### v2

# Privacy in Speech Technology

Abstract
Speech technology for communication, accessing information, and services has rapidly improved in quality. It is convenient and appealing because speech is the primary mode of communication for humans. Such technology, however, also presents proven threats to privacy. Speech is a tool for communication and it will thus inherently contain private information. Importantly, it however also contains a wealth of side information, such as information related to health, emotions, affiliations, and relationships, all of which are private. Exposing such private information can lead to serious threats such as price gouging, harassment, extortion, and stalking. This paper is a tutorial on privacy issues related to speech technology, modeling their threats, approaches for protecting users’ privacy, measuring the performance of privacy-protecting methods, perception of privacy as well as societal and legal consequences. In addition to a tutorial overview, it also presents lines for further development where improvements are most urgently needed.
Index Terms:
speech technology, privacy and security, machine learning
I Introduction
Speech is a mode of communication and thus inherently contains a wealth of information (see table I). Communicating information already known by the receiver is pointless, and efficient communication will thus mainly contain information that is not widely known or which is private. In addition, speech contains also a wide range of side information like the state of health and emotions, as well as physical, psychological, and social identity, most of which is private information. Finally, speaking is a dynamic interaction between two or more speakers. Dialogues thus also contain information about the relationship between participants such as their level of familiarity, affiliation, intimacy, relative hierarchy, shared interests, and history. From an information content perspective, we can thus expect that privacy is a multifaceted issue in all areas of speech technology.
TABLE I: Private information potentially identifiable from speech signals, and the extent to which they are sustained over time and can be willfully controlled, summarized by and reflecting the personal judgment of the author.
Category Examples Permanence Control
Biological Body characteristics Sustained No
State of health Variable No
Level of intoxication Variable Yes
Psychological Emotions Variable Partly
Intelligence Sustained No
Education, skill Sustained Yes
Gender identity Sustained No
Message Text, emphasis, style,
expression
Variable Yes
Mannerisms, context Variable Partly
Language
choice & skills
Partly Partly
Affiliation Ethnic, national,
cultural, religious,
political, etc.
Sustained Partly
Relationship
character
Hierarchy, familiarity,
attraction, intimacy
Sustained Partly
Has met with person Variable Yes
Physical
environment
Background sounds,
distance to sensor,
transmission distance,
reverberation, location
Variable Yes
Hardware
used
Sensor type & manufacturer Variable Yes
TABLE II: Practical examples of threats related to private and sensitive information that is conveyed by people’s voices.
Threat Example
Price gouging Signs of depression or other health problems in users’ voices could be misused to trigger an increase in their insurance premiums. Signs of users’ emotions could be exploited to offer them products at higher prices.
Tracking, stalking Voice re-identification could link users across platforms, i.e., from work-related social media to online support groups and dating apps, making it possible to follow them anywhere.
Extortion, public
humiliation
Private health problems, and romantic affairs could be detected in the voice and used for blackmail or made public against a user’s wishes.
Algorithmic stereotyping Recommender systems based on voice can become biased with respect to age, identity, religion, or ethnicity, in ways that are nearly impossible to monitor.
Harassment, inappropriate advances Users in chat rooms or virtual reality could be automatically singled out by gender or opinions, making them a target for unwanted attention and harassment.
Fear of
monitoring
The subjective feeling of being continuously monitored may cause psychological damage. It may also stifle political expression damaging democratic societies.
Studying and improving privacy in speech technology is important because breaches in privacy can have serious consequences. Table II lists examples of threats and exploits. The list is incomplete, and we can expect new threats to be discovered. The generic solution template is, however, typically always the same: Minimize transmission and storage of as well as access to sensitive information that is irrelevant to the service that the users want. It is then ”merely” a question of how such minimization is achieved, what information is relevant, and how to determine what the user wants (provide information and enable control of services and threats).
With respect to threats, media attention is often focused on breaches which have large economic consequences [Lynskey2019GuardianPrivacyScandal, hern2019, brewster2021]. While such breaches are important by themselves, the media attention introduces unfortunate biases in two ways. First, a breach is a worst-case event, whereas threats that have not yet led to a breach event can already have a large impact. For example, users may choose not to use systems that threaten their privacy; known weaknesses and even a lax attitude of the service provider toward privacy can, therefore, have an effect on the adoption, retention, and sales of products and services. Importantly, users can avoid systems that are perceived threatening, even when there is no actual threat. To maintain users’ trust, it is therefore important to both uphold the users’ actual privacy but also design systems to actively and frequently communicate the level of privacy, including known threats, their current status, and measures taken to protect against them.
Second, while a single breach in a large service can have large economic, psychological, societal, and legal consequences, small crimes are so common that their joint effect is comparable or larger in size [levy2020intimate]. With speech technology, such ”small” crimes include stalking, extortion, harassment, humiliation, and inappropriate advances (see table II). While the economic damage of a single such incident can be small, when combined, their joint psychological and societal effect is potentially large and their prevalence makes them a considerable threat also economically.
This paper presents a tutorial overview of privacy in speech technology that covers a wide range of threats, methodologies, and algorithms. Analysis of threats, however, demonstrates that while attack surfaces have great variety, the threat models are similar (see section III). Protections against those threats have five distinct categories of approaches, isolating target information by removing private information, limiting access to private the message through cryptography, reducing centralized threats through non-centralized processing, limiting access to reproduced audio, and improving overall system performance (see section IV). To quantify the extent of protections, we further need methods for evaluation of privacy and utility (see section V). While objective evaluation quantifies the actual level of privacy, users’ impressions of privacy do not necessarily follow the objective level. For the best user experience, therefore, we need to quantify and understand users’ perceptions, experiences, and preferences related to privacy, as well as design systems accordingly (see section VI). The range of applications and content types where privacy-preserving processing methods are needed is vast, and section VII presents a brief overview thereof. Finally, threats to and breaches in privacy have considerable consequences on both societal, economic, and individual levels, which has motivated governments to increasingly regulate the use of technology (see section VIII).
To the author’s knowledge, this is the widest tutorial overview of privacy in speech technology. Recent works in the field, however, include a technical overview [nautsch2019preserving], as well as a popular-science review [backstrom2020privacy] of this area, and several doctoral theses have their respective summaries, e.g. [Nelus2022, das2021thesis, zarazaga2022thesis, pathak2012privacy, williams2022learning, portelo2015privacy]. Privacy has been extensively discussed in other areas of science, such as the neighboring field of natural language processing [mahendran2021privacy], in statistical theory of privacy [cummings2023challenges], in social sciences [petronio2002boundaries] and psychology [Derlega1977PrivacyAS]. There is even an excellent and thorough meta-study of all areas of science that discuss privacy [van2018handbook].
Bob Acoustic pathway User
Alice
Acoustic pathway Local device Network Remote service
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Eve Local device Secondary service
XX5765432X1
Figure 1: Threats to the privacy of a user Alice when speaking with another person Bob or a local device, where information is leaked or shared through the acoustic pathways, the local device, the network or the cloud service, to another person Eve, device or service, contrary to the preferences of Alice (red arrows). Each threat is marked with the number of the corresponding part of section III-D. Dotted red arrows marked with “X” indicate threats outside the current scope and dashed black arrows indicate the potential retransmission of leaked data beyond the primary attack surface.
II Goals and Definitions
This paper focuses on the two primary uses of speech technology: telecommunication and human-computer interfaces. Concerning telecommunication approaches to ethical, legal, and technological questions related to privacy in telecommunication over landlines are well-established, and open discussions are primarily related to how the existing regulation and oversight should be extended to cover also mobile telecommunication and voice-over-Internet protocols (VoIP) [penders2004privacy, wong2009privacy, arapinis2017analysis]. The primary focus of this paper is thus human-computer interfaces where a computer processes speech signals (see fig. 1), but threats to communication applications are also considered where applicable. Focus is here, however, limited to the acoustic speech signal, also known as the voice since natural language processing is a distinct and largely independent field that warrants treating it separately (e.g. [mahendran2021privacy]).
Privacy is closely related to challenges in security and it is often difficult to distinguish between them. Here, we strike a balance by considering privacy scenarios where an agent has legitimate access to some private speech data of the user but uses it for purposes contradicting with, or gains access beyond, the users’ expectations or preferences. For example, a voice interface can be used to control home automation, but if the service provider shares that voice data with advertisers against the users’ preferences, then it violates privacy (see fig. 1). Consequently, security concerns such as identity spoofing, deep fakes, and attacks on devices or networks are excluded here, as these threats are more related to security than privacy.
A private message is the content of a communication that is required to fulfill the purpose of said communication. Side-information is any and all information in a communication other than the private message. See fig. 2.
We use the terms attacker and unintended recipient loosely to describe any entity or agent that gains access to or uses private information in a way that conflicts with the user’s preferences. An malicious attacker is intentionally seeking to gain private information, a targeted attacker attempts to gain such access to a specific person, an opportunistic attacker to any person, a motivated attacker is seeking ways to attack, and a curious attacker seizes emerging opportunities.
Attributes of speech are properties that can be distinguished from a speech signal and describe the speaker(s) or their state, the speech content, style or environment, the recording configuration, or any other related features. Table I lists examples of speech attributes with private information. Attribute or property inference refers to the identification of such private attributes of speakers of known identity or otherwise. Re-identification refers to identifying individuals from apparently anonymized information [ISObiometric, ISOIEC224745_2022].
Enrollment data is reference data, typically labeled with, for example, speakers’ identities or other attribute(s), to which observations are compared with the purpose of identifying said identity or attribute(s). It is used in the context of trusted tasks, where the agent is authorized to access this information. In threat tasks, where an attacker has gained access to similar reference data, the corresponding data is known as found data. Conversely, speech samples that are compared to a database of samples are here referred to as observations, and in literature, they are sometimes known as trial samples, recognition samples, or public data. [tomashenko2022voiceprivacy]
When a speaker makes a speech command to a computer or service, this implies that the speaker is giving a task to the computer and this is known as the trusted task. In contrast, we use threat tasks to model the activity of attackers trying to extract information or perform tasks in a nefarious way.
The attack surface is the specific point in the system where private information leaves the authorized path to make its way to the attacker.
Voice conversion refers to any speech processing method that replaces the speaker’s identity or other attributes in a speech signal with another. Anonymization of a speech signal refers to voice conversion where the purpose is to conceal private information. Pseudonymization further refers to reversible anonymization, where the original signal can be recovered if the concealed attributes are known. [^1] These definitions of anonymization and pseudonymization are outcomes of a community-wide discussion about personally identifying information (PII) at the Lorentz Center workshop “Speech as PII: Ethics & Usability in vis-à-vis with Security & Privacy” in 2021. The anonymized speech signal then corresponds to a speech signal of an artificial person, who is known as a pseudospeaker.
III Threat and attack models
To better understand the decisive aspects of threats, it is useful to model threats and their components. The most important components of such models include the exposed data sources (section III-A), tactics used by, and data available to the attacker (section III-B), the flow of data on an abstract level, also known as the attack model (section III-C), as well as the specific point in the system where private information leaves the intended path, known as the attack surface (section III-D).
Sender Channel RecipientRefer to caption
LegitimateRefer to caption Undesired
!! Private message Side-information
Figure 2: The high-level threat model of speech interaction, where a private message is sent through a channel to the legitimate recipient, but consequential side-information is bundled to that message. It is a threat to privacy when an undesired recipient gains access to that private message or side information (marked by red arrows and exclamation marks).
III-A Data sources
Speech data can be exposed in two ways (see fig. 2): First, when a private message is transmitted, it is a threat to privacy when contrary to the preferences of the user, it is used in an unexpected way or transmitted to a third party. Second, since speech contains a wide range of private information (see table I) which is bundled in the acoustic signal in complicated ways, it is challenging to extract only the desired message. Typically, speech information thus always contains consequential side information bundled into the private message. It is a threat to privacy when such private side-information is transmitted to an undesired recipient alongside the private message. Per definition, we here assume that the legitimate recipient can be trusted with the side information and that any undesirable use of it is labeled as an undesirable service.
The difference between the two forms of information makes an important difference in approaches to mitigation. Transmitting the private message to the recipient is the desired outcome of the interaction and we can thus not complete the desired task without sending the message. In some cases, though, we can reduce the amount of information content in the private message, and more generally, we should always secure the transmission for example through cryptography (see sections IV-D, IV-B and V-A2). Some cryptographic methods, such as homomorphic encryption, even keep the message secret from the receiver but allow the processing of a desired task. However, when side-information is exposed, as additional methods we can also use signal processing to better remove, replace, or distort such side-information (see section IV-A).
III-B Attack types
There are three main types of attacks; motivated, opportunistic, and incidental. In motivated attacks, the attacker attempts to extract information about a specific person. First, the attacker can try to extract private information from speech samples from that person. For example, an attacker could extract private health information from a public speech of a famous person. Second, an attacker can browse through a database of anonymous speakers in an attempt to identify the target person. This is known as re-identification, as the attacker thus recovers the identity of a speaker from a database where the identities were not available. When the target speaker is re-identified, the attacker can then extract private information from the re-identified sample. Moreover, the fact that the voice of the target person is in the database can be private or sensitive information in itself. For example, if a database contains speech samples from the members of a particular association, then the re-identification of a target person reveals that person’s affiliation with that association.
Opportunistic attacks are usually known as birthday attacks, referring to the birthday paradox [menezes2018handbook], where the attacker uses a known weakness on a large pool of potential victims. It is like spam email; if the attacker sends a large enough number of such emails, then randomly someone triggers the trap by mistake. In this form of attack, anyone could be the victim. Where motivated attacks typically target high-value victims such as famous persons where both protections and rewards are high, birthday attacks typically target ordinary people where the likelihood of success is larger but rewards are smaller.
The third category of “attacks” is incidental, where an agent gains excessive access to information unintentionally or by mistake. For example, if a speech device misinterprets an utterance, it can trigger functionalities that reveal private information to unauthorized parties. Such incidents are thus neither intentional nor nefarious, but in the interest of consistency, here they are still called attacks.
Attacks can further be classified according to the amount of information available for the attacker, such as information about the speakers and about the models used. Access to speech samples from the target user (known as enrollment data) are particularly important for the attacker. For example, in the anonymization task of the VoicePrivacy 2022 challenge, the objective was to replace (pseudonymize) speaker identity but retain all other speech characteristics such as linguistic information. The anonymized sentences are known as trial utterances. The attack scenarios were then classified as [tomashenko2022voiceprivacy]:
  1. 1.
    Unprotected: no anonymization is performed; attackers have access to original trial and enrollment data.
  2. 2.
    Ignorant attacker: Trial data is anonymized, but attackers are unaware of it, hence they use original data for enrollment.
  3. 3.
    Lazy-informed: Both trial and enrollment data are available to attackers, but anonymized with different pseudo-speaker.
  4. 4.
    Semi-informed: As in lazy-informed, attackers can also train their model with the same anonymization system but different pseudo-speakers.
With increasing information, obviously, the attacker’s task becomes easier and the accuracy of the attacker’s models improves. Further, such attack scenarios can then be devised according to the specific use case. For example, if the task is to anonymize emotional or health status, we can define different attack scenarios depending on the extent to which the attacker has information about categories of emotions and health status used in anonymization.
III-C Attack model
In the design of privacy-protecting methods, attack models simulate and quantify an attacker’s performance. In other words, safeguards designers can put themselves in the attacker’s place to better understand weaknesses and quantify risks.
Figure 3 illustrates an attack model for measuring the extent of privacy [maouche20_interspeech]. We assume that there is some private data, which is anonymized such that it can be used in a trusted task. The attacker has access to the anonymized data, which is here called “public data”. Observe that the public data is not necessarily openly available for anyone to see, but it only indicates that the data is sufficiently freely available that the attacker has access to it. The attacker has also access to some other data about speakers, found from some other source (found data), which helps in extracting private information. Again, the term “found data” is a loose term and denotes any data other than found data that is available to the attacker.
Private data Anonymization Public data Downstream application ! Private
information
Found data Attacker
Figure 3: An attack model for the evaluation of privacy-preserving anonymization, where private data is anonymized to remove private information, and the anonymized data is shared publicly. An attacker uses any available (found) data and anonymized public data to infer private information contrary to the users’ preferences. Anonymized data flow is indicated by dashed lines and the attack by red lines and an exclamation mark.
III-D Attack Surfaces
We categorize scenarios according to the attack surface from where information is extracted (see fig. 1): a channel between remote (cloud) services (section III-D1), a user interface to the local (edge) device (section III-D2) or to the cloud service (section III-D3), from the local network (section III-D4), the acoustic pathway (section III-D5), or through the shared user interface of the local device (section III-D7). We consider only cases where a piece of technology is receiving or transmitting information and exclude human-to-human communication without devices. We also assume that devices and network connections are secured such that only authorized services can communicate with them. The threats we focus on are thus illustrated with numbered, solid red lines in fig. 1.
III-D1 Cloud Leak
Suppose a user accesses a remote server, known as a cloud service through a local device, known as an edge device (see fig. 4). The cloud service is thus assumed to have legitimate access to the private messages of the user. The cloud server, however, can then use that private message or bundled side information for some other purpose than that requested by the user, extract more information than anticipated, combine it with other information, or share information with a third party. For example, the cloud server of a voice assistant could inappropriately share information with an advertiser. Observe that though figs. 4 and 1 show the secondary remote service as an impersonal cloud server, this analysis includes harms done by the people and organizations that control those servers. For the current analysis, it, however, does not make a difference who is the actor in perpetuating harm or is responsible for the harm caused to the user, but the pertinent difference is the first point where private information departs from the intended flow of information (cf. section III-D3).
User Alice Local device Remote service Secondary service Refer to caption Refer to caption Refer to caption Refer to caption !
Figure 4: Threat scenario “Cloud Leak”, where a user Alice accesses a (primary) remote service using a local device, but the information is shared to a secondary service contrary to preferences (red arrow and exclamation mark).
User Alice Local device Remote service Refer to caption Refer to caption Refer to caption !
Figure 5: Threat scenario “False Activation”, where a user Alice accesses a local device, but the information is shared to a cloud service contrary to preferences (red arrow and exclamation mark).
III-D2 False Activation
Devices with speech interfaces can hear all conversations in the same acoustic space as the device and, therefore, need mechanisms to determine which utterances are intended for the speech interface (see fig. 5). A popular approach is to use a specific utterance, known as a wake word, to start all interactions with the interface [wang2021wake]. The wake word is then like a rudimentary password, which prevents the interface from activating when speech is not directed to the device. Unfortunately, designing wake word detectors is non-trivial, and they will occasionally make mistakes. They might sometimes miss a wake word when it is spoken (false negative) or mistake some other unrelated sound as a wake word (false positive). While false negatives are annoying for the user when the service does not activate, false positives potentially present serious threats to privacy. In some famous cases, speech interfaces have activated from sounds on the television to buy unwanted items at the users’ behest, and users’ private conversations have been leaked to third parties [liptak2017alexa, moscaritolo2018alexa].
User Local device AliceRefer to caption Refer to caption EveRefer to caption Refer to caption Remote service Cloud storage Refer to caption Refer to caption !!
Figure 6: Threat scenario “Cloud Access”, where user Alice uses a remote service through a local device, where it is potentially stored in the cloud and shared with user Eve, contrary to Alice’s preferences (red arrow and exclamation mark).
III-D3 Cloud Access
Figure 6 illustrates the threat where a user Alice accesses a cloud service through an edge device, and where some private information of Alice is stored. A second user, Eve, can then access the same cloud service through another device and potentially gain access to the stored private information, contrary to Alice’s preferences. This attack surface is similar to the Cloud Leak scenario, with the main difference being the recipient, which is here a person, whereas, in the Cloud Leak scenario, it is an automated agent.
An example of this threat is medical records, which can be collected from patients to a centralized database. A researcher with authorization to access the database can then potentially extract private information beyond the expected. [Teixeira2018]
User Devices in same network Remote service Refer to caption Refer to captionRefer to captionRefer to caption Refer to caption Refer to caption Refer to caption ! Outside the room Room
Figure 7: Threat scenario “WASN Authentication”, where a user uses a service through a wireless acoustic sensor network (WASN) inside a room, but another device, outside the room in the same network, joins the distributed sensor network and shares information contrary to preferences (red arrow and exclamation mark).
III-D4 WASN Authentication
The audio quality of speech pickup (in terms of, for example, signal-to-noise ratio or perceptual quality), as well as the usability of voice interfaces, can be improved by using all available connected devices with microphones in the same room or acoustic space [das2021enhancement, zarazaga2020acoustic, sigg2020provable]. Such collaboration can be realized with wireless acoustic sensor networks (WASNs), where several independent devices simultaneously pick up speech, and those channels are combined to obtain a high-quality signal (see fig. 7[Nelus2022, zarazaga2022thesis]. A sensitive question is, however, authorization: Which devices are allowed to share information with each other?
One approach is to assume that devices that are in the same room or acoustic space can hear the same signal [zarazaga2022thesis, koppelmann22_interspeech]. Presence in the same space is thus already an implicit authorization to participate in a joint signal pickup. To protect privacy, it is then necessary to determine which devices reside in the same acoustic space. Devices in a different room can belong to the same company or family and be connected to the same network, but they are still outside the sphere of the current discussion. Without proper authorization mechanisms, devices outside the room could then gain access to private speech inside the room.
User Local device Remote service AliceRefer to caption Refer to caption Refer to caption EveRefer to caption Refer to caption Refer to caption !Room
Figure 8: Threat scenario “Speech Interface Leak”, where a user, Alice, accesses an edge device (potentially including also a remote service), but another local person, Eve, or device in the same room or acoustic space overhears the speech interaction contrary to preferences (red arrow and exclamation mark).
III-D5 Speech Interface Leak
Figure 8 illustrates a scenario where a device or person overhears a discussion between a user, Alice, and a local device. The defining property is that the leak is in the acoustic pathway, necessitating that the eavesdropper is physically present in the same acoustic space as the private communication. The difference between these leaks and False Activation (sections III-D2 and 5) is that here, the eavesdropper is collecting data inappropriately, whereas in False Activation, an incorrect functionality is activated.
In the case where Eve overhears Alice’s speech, we assume that people have a learned awareness of which other people are present in the same room and adjust their speech accordingly. Then either Alice does not mind that Eve overhears her speech (inconsequential information), Alice can change her speech style to a whisper, or change content such that Eve does not hear anything private (reduced information transmission), or she can go to a different room to continue the interaction in private (modified acoustic channel). Whether users are similarly aware of devices present in the room, through anthropomorphization of devices or otherwise, and adjust their speech accordingly is an open question.
Local dialogue Local device Remote service Alice Bob Refer to caption Refer to caption Refer to caption Refer to caption !!
Figure 9: Threat scenario “Discussion Leak”, where a dialogue between two users, Alice and Bob, is picked up by a local device and potentially shared to a remote service contrary to preferences (red arrow and exclamation mark).
III-D6 Discussion Leak
Figure 9 illustrates a discussion between two persons overheard by a local device and, like False Activation, triggers unintended functionalities in the local device. The main difference to a Speech Interface Leak (sections III-D5 and 8) is that the type of speech is different. Interaction with a speech interface at the current level of technology is, roughly speaking, limited to speech commands, question-answer pairs, and one-sided monologues, whereas spoken human-to-human interaction has potentially a much more dynamic character. This is potentially beneficial in the sense that the speaking style of human-to-human interaction could be recognized, preventing unintended triggering of functionalities.
UserAliceRefer to caption EveRefer to caption Local device Storage Refer to caption Refer to caption !
Figure 10: Threat scenario “Shared Device”, where user Alice uses an edge device, where information is stored and shared with another user, Eve, contrary to Alice’s preferences (red arrow and exclamation mark).
III-D7 Shared Device
Many practical scenarios include multiple users sharing one or several speech-operated devices. For example, a family can share a smart speaker or television, an office meeting room can have smart devices and customer service points can have a phone shared across duty officers. Each of these devices can collect private information about the user(s) over time and can potentially share it with other users (see fig. 10). Notably, this scenario highlights that leak happens over a distance in time, whereas threats occurring over a spatial distance often receive most of the attention.
This threat model clearly demonstrates that devices and services used by multiple users need to employ access control and authorization management if they store any private information. It is also obviously related to many security threats – unauthorized access to devices should be prevented – but those are outside the scope of this work.
IV Protections
The main avenues for protecting privacy in speech technology are: 1. Information isolation, where private information is filtered or otherwise removed from a transmitted (or stored) channel of speech information (section IV-A), 2. Secure processing, where access to private information is prevented with methods of cryptography (section IV-B), 3. Privacy-preserving architectures, where private information is not sent to the cloud, but processed either only locally (on the edge), or in a distributed network of (nearby) devices (section IV-A3), 4. Acoustic interventions, where reproduction of audio is changed such that private information is not revealed to eavesdroppers (section IV-D), and 5. Improving performance, where the intended functionality of a system is improved such that processing errors leak less private information (section IV-E).
IV-A Information isolation
By isolation of information, we refer to the design of a channel of information where the target information is isolated from other (undesired) private information. Such isolation can be enforced through two main approaches; 1. we can reduce the capacity of the channel such that it can represent only the target information and has no room for other private information, which is known as an information bottleneck [tishby2000information] (see section IV-A1), or, 2. we can use adversarial training of the model, where the model is trained such that an attacker cannot extract a specific type of private information from the channel (see section IV-A2). Such isolation can be implemented for any category of information. More generally, a system can include multiple channels of information, each designed to carry a distinct category of information, in an approach known as disentanglement. This allows applications to cherry-pick the private information that should be transmitted case by case (see section IV-A3).
InferenceTraining User EncoderBottleneckDecoder Loss Refer to caption OutputLabelLocal device Remote service !
Figure 11: Training of a privacy-preserving speech analysis method with the information bottleneck principle. The point of attack is indicated by the red arrow and exclamation mark.
InferenceTraining User EncoderBottleneckDecoder Loss Refer to caption Output--!$+$
Figure 12: Training of a privacy-preserving an autoencoder structure as an example of the information bottleneck principle. The point of attack is indicated by the red arrow and exclamation mark.
IV-A1 Information Bottleneck
The purpose of an information bottleneck is to use an encoder to distill the information needed to perform a downstream (trusted) task into a bottleneck, such that all other (private) information is removed.222 A concept related to information bottlenecks is the privacy funnel [makhdoumi2014information]. Where the original definition of the information bottleneck [tishby2000information] aims to maximize the utility of a channel with a fixed rate, the privacy funnel aims to minimize the information rate with a fixed utility. The information bottleneck is, however, much more widely known in the community and often informally used as an umbrella term to cover both cases. A decoder can then be used to decipher the information in the bottleneck into a format usable for the downstream task. The overall approach is generic and can be applied for example in cases where 1. the decoder resynthesizes the signal with some attributes anonymized [wang2019vector] or 2. information extraction scenarios such as classification [tishby2000information], where the decoder can often be trivial.
To distill only the private message and discard any side-information, speech information is passed through a bottleneck so tight, that only the legitimate message at full accuracy can pass through [Tishby2015, makhdoumi2014information] (see fig. 11). This approach thus provides protection against an attacker who has access to the output of the bottleneck. The challenges are to design a training methodology and model structure with a sufficiently tight bottleneck so that the legitimate message is accurately retained while private side information is discarded.
A frequently used approach to implementing a bottleneck is an autoencoder structure, consisting of an encoder, bottleneck, and a decoder, where the objective is to reconstruct the input signal from the bottleneck output [malekzadeh2018replacement, perero2022x, lu2013speech] (see fig. 12). However, to be privacy-preserving, the bottleneck should be sufficiently tight that the original speech signal cannot be perfectly reconstructed to resemble the original in their waveform with respect to some distance function. Designing a loss function is then challenging when the bottleneck is so tight that the decoded signal is very far from the original. One solution akin to the machine learning method known as representation learning is to feed the output of the autoencoder again to the encoder and compare the bottleneck features [neumann2019improving, chorowski2019unsupervised].
Information content at the bottleneck can be reduced by either reducing the signal rank (viz. dimensionality reduction, e.g., [perero2022x]) or by quantizing and coding the signal at a low bitrate (e.g., [wang2019vector, williams2021revisiting]). The trade-off between the accuracy of trusted task and bottleneck entropy then determines the extent of privacy (see section V). Observe that dimensionality reduction is, in general, not alone sufficient because if the channel’s floating- (or fixed)-point representation has $K$ dimensions each with $B$ bits, then a sufficiently advanced encoder can pack $KB$ bits of information to that channel. Since typical representations have $B=32$ or $B=64$, which can encode up to $2^{B}$ unique speaker identities, we see that the bottleneck is not particularly tight. As a general rule, therefore, bottlenecks should always be quantized to the lowest bitrate sufficient to represent the target information.
A central advantage of the information bottleneck approach is that it gives a provable level of privacy. If the bitrate of the bottleneck is $B_{b}$ and that of the target information is $B_{t}$, then the privacy leak is at most $B_{leak}\leq B_{t}-B_{b}$. This applies to single vectors of target information passed through a bottleneck with full accuracy. Streaming data through the bottleneck, however, over time, leaks a flow of private information that increases in size without bound [kairouz2015composition]. This approach also addresses only the average disclosure of private information but does not give guarantees of worst-case disclosure [nautsch20_interspeech, vali2024porcupine].
Information bottlenecks, or more accurately, privacy funnels, are theoretically well-suited for privacy-preserving applications. They have also been used in other areas, such as privacy-aware transmission of images [sun2024privacyaware] and face recognition [razeghi2024deep]. However, the above-mentioned applications in speech either do not specifically apply it for privacy or use information bottlenecks optimized for privacy with additional components.
InferenceTraining User Encoder Trusted Task Trusted Loss Refer to caption Output Threat Task Threat Loss !$+$--
Figure 13: Training of a privacy-preserving speech analysis method with an adversarial approach. The trusted task is authorized to extract some information, while the threat task (drawn in red) is extracting some other private information. During training, the trusted task competes with the adversarial threat task, such that in the encoder block, all private information is removed. During inference, the threat task can be ignored.
IV-A2 Adversarial Approach
An alternative to reducing the size of the information bottleneck is to use an adversarial model during training, such that side-information in the bottleneck is minimized [noe21_interspeech, srivastava19_interspeech, nelus2018itg]. Figure 13 illustrates the model structure, where the trusted and threat tasks correspond respectively to utility and privacy, which in turn respectively correspond to the extraction of the private message and side-information. The threat task is independently optimized to extract private side information from the bottleneck output. While keeping the threat task fixed, the encoder and the trusted tasks are jointly optimized to maximize the accuracy [^3] Accuracy refers here to some similarity measure between the message and a corrupt version of the message. of the private message and to minimize the accuracy of private side information.
Adversarial training thus combines maximizing and minimizing the accuracy of private side information. This forces the encoder to minimize private side information passed through the bottleneck since the threat task tries to maximize the extraction of private side information. Similar effects can be achieved by a variety of other formulations, for example, by Siamese training, where the analysis model is specifically optimized to minimize and maximize feature distances between speakers within and across target classes [nelus19b_interspeech]. This approach should cluster extracted information into groups corresponding to the trusted task while minimizing private information.
An advantage of an adversarial configuration is that the designer can specifically choose which type of threat and information category is removed from the data. The system can also be optimized end-to-end, such that all components are jointly optimized for the best performance. This can lead to an efficient model in terms of the trade-off between computational cost and output quality.
The principal issue with adversarial training is that it does not give any formal assurances of privacy (cf. section V-A). The best it can do is to demonstrate the extent to which the chosen adversarial model was unable to extract private side information from the chosen category of private information. We can thus expect that some other attacker with a better model could still extract private side-information and that other categories of private information are potentially present in the bottleneck output.
User EncoderBottleneckDecoderRefer to caption OutputAnonymization!
Figure 14: Privacy-preserving speech processing through disentanglement, where speech is decomposed into independent streams of information, each representing a distinct category of information and where the level of anonymization can be individually chosen for each category. The point of attack is indicated by the red arrow and exclamation mark.
IV-A3 Disentanglement
With disentanglement, we refer to methods that split a speech signal into channels corresponding to distinct categories of information. For example, selecting from table I we can choose to decompose speech into linguistic information, the speaker’s (voice) identity, and emotions [zhou2021vaw]. Such disentanglement can be achieved by either or both through adversarial training or by generalizing information bottlenecks to encompass multiple bottleneck channels, where each bottleneck represents a disentangled representation (see fig. 14). Note that provable privacy requires that such channels are statistically independent of each other. That is, each channel represents a distinct and statistically independent attribute of the signal such that the extent of anonymization can be cherry-picked per channel as per use-case e.g. [aloufi2020privacy, stoidis21_interspeech, zhou2021vaw, williams2022learning].
An important challenge with disentanglement is to define the constraints with which information is funneled to the corresponding channel and the methods by which they are made independent of each other. For example, a model can be trained to match specific channels with labels in the dataset (e.g. [stoidis21_interspeech, zhou2021vaw]). Alternatively, using representation learning, we can constrain channels based on objective criteria such as time-scope, i.e. the length of time over which data is integrated (e.g. [Niekerk2020, neumann2019improving, lotfidereshgi2022cognitive]). Specifically, attributes such as the speaker’s identity are constant over an utterance, while phonemes, intensity and pitch change over a time scale of tens of milliseconds. By modeling the signal in terms of such time scales, we can disentangle attributes that appear on the corresponding time scales.
User Alice Local device Remote service Refer to caption Refer to caption Refer to caption Encrypted!
Figure 15: Protecting privacy by encrypting communication and processing. The gray area indicates the domain that is encrypted and red, double arrows marked with exclamation marks indicate the protected stream of data.
IV-B Secure processing
If the transmitted information is encrypted, then it cannot be used for malicious purposes without access to the key. Surprisingly, however, it is possible to process encrypted information when using secure multi-party computation (SMPC), such that the encrypted result of the computation can be returned and opened [yao1986generate, portelo2015logsum, portelo2014privacy, Teixeira2018, cramer2015mpc] (see fig. 15). The SMPC umbrella contains several categories of methods such as homomorphic encryption, oblivious transfer, and garbled circuits. The general principle is that in an interaction between two agents, Alice and Bob, Alice can send encrypted data for processing or a request to Bob, such that Bob cannot interpret the input but can perform the requested task. Furthermore, when Alice receives Bob’s response, she cannot deduce the operations that Bob has performed, such that the desired task is completed with mutual privacy preserved. Another category of approaches is distance-preserving hashing techniques, where data is obfuscated through randomization such that the desired communication is preserved while other private information is kept secret [portelo2015privacy, jimenez2015secure].
Such secure techniques can be applied, for example, in the extraction of spectral features of speech or speech recognition in the cloud, without plain text access to the speech signal [thaine19_interspeech, Zhang2019]. It can also be applied in privacy-preserving speaker recognition, authentication or verification [teixeira22_interspeech, pathak11_interspeech, pathak2012ppsa, portelo2014bprivacy], speech search [portelo2015bprivacy], or for secure sound classification [shashanka2006secure]. Finally, we can compare the similarity of audio signals without revealing them [rane2013privacy, perez2020, zarazaga2022thesis].
A central problem with such cryptographic protocols is that the operations that can be performed in the encrypted domain are limited, typically to only additions and multiplications [smaragdis2007framework, pathak2012ppsv, pathak2011privacy, pathak2013privacy]. In contrast, typical speech-processing algorithms also require other operations, such as non-linearities, conditionals, and recursions [chialva2018conditionals], making the algorithms impossible to implement. While many operations have standard workarounds [teixeira2019privacy], such as approximation with additions and multiplications, such approximations are often computationally prohibitively complex, see e.g. [thaine19_interspeech]. In addition, the communication overhead, in terms of the amount of data to send in comparison to the size of the secret message, is very large. For a broader overview of cryptographic methods in speech, see [nautsch2019preserving].
User Local device Remote service Secondary service (a)Refer to caption Refer to caption Refer to caption Refer to caption (b)Refer to caption Refer to caption Reduced information ! Local processing
Figure 16: Two alternative protections against the “Cloud Leak” scenario, where the transmission of private information to a remote service is either (a) reduced (by some means and to some extent) or (b) prevented.
Users Edge devices Remote service Refer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to caption Refer to caption Model updates !!
Figure 17: Federated learning as an example of distributed learning, where devices train models locally and transmit only model updates to the cloud. Red dashed arrows marked with exclamation marks indicate the reduced flow of information.
IV-C Privacy-preserving architectures
Perhaps the most obvious solution to protecting privacy, for example in the Cloud Leak–scenario, is to limit information flow (to some extent) from the local edge device to the cloud server (see fig. 16). Such limitation or reduction in information flow can be implemented by removing, replacing, or distorting private content as in fig. 16(a) (cf. section IV-A). At the extreme, the edge device can be entirely disconnected from the cloud or network, and process information only locally, as in fig. 16(b) [shi2016edge]. In doing so, we then need to assume that the local device has sufficient capacity and software to complete the requested tasks, that it is not compromised, and that the restriction of information flow is sufficient to ensure that private information is not leaked. Observe, for protecting privacy in scenarios 1 to 4 in fig. 1, local processing is the best protection as nothing is transmitted to a cloud service.
A majority of advanced speech processing today, however, uses machine learning, which has to be trained using large databases of speech. The best quality data correspond closely to scenarios where the services are used. Recording users’ interactions with their devices is then attractive for training improved models since it corresponds exactly to the use case. This presents a considerable threat to privacy because such unrestricted recording could capture a wide range of private information, including all interactions with the device but potentially also any and all speech in the vicinity of the device.
Distributed learning is an approach to training models without the need for centralized data collection [liu2021machine]. Models are trained on local edge devices and only model updates are shared between the nodes and/or with a central server (see fig. 17). The model updates are, however, calculated from the raw signal and thus describe the original signal in some convoluted way. Though the raw data thus never leaves the edge device in distributed learning, this corresponds to a privacy protection approach where information flow is restricted (cf. section IV-A).
Federated learning is one of the flavors of distributed learning, where a central, cloud server collects, merges, and redistributes model updates. It has been used for example in speaker and emotion recognition, language modeling, as well as for unsupervised estimation of microphone clusters in sensor networks [feng22b_interspeech, granqvist20_interspeech, zewoudie2021federated, nelus2021, shoemate2022sottovoce, ro21_interspeech, nguyen2023federated, feng2021attribute].
Overall, distributed learning is a promising approach, even if it has several challenges. First, constructing, training, and testing systems architectures is much more complicated than regular machine learning. Second, model updates are model-specific and cannot easily be reused if the model structure is updated. The learning accumulated during training is then effectively lost every time the model is updated, jeopardizing the fair comparison of competing approaches. Third, even if distributed learning does improve privacy, it is not a guarantee for privacy since model updates can also contain private information [Orekondy2018gradientleaks, liu2021machine]. To be specific, except for federated learning, distributed learning protects efficiently against threats 1 to 3 in fig. 1, but it may potentially leak private information to other devices, and federated learning may leak information to the cloud.
IV-D Acoustic interventions
UserLoudspeaker Local device AliceRefer to caption Refer to caption Refer to caption EveRefer to caption Refer to caption Microphone!
Figure 18: Protection against a reproduction leak by authorization tracking of users, where the device keeps track of people present in the room such that private information is not shared with unauthorized listeners (red dashed line and exclamation mark).
UserLoudspeakers Local device AliceRefer to caption Refer to caption Refer to caption EveRefer to caption Refer to caption !
Figure 19: Protection against a reproduction leak by using sound zones, where constructive interference between loudspeakers is used to retain intelligibility for user Alice, and destructive interference distorts it for user Eve (red dashed lines and exclamation mark).
When a user interacts with a speech interface, then the spoken answer of the speech interface can contain private information. This private information can be overheard by other users in the same acoustic space, and that presents a threat to privacy (see figs. 19 and III-D5). In theory, it would be possible to identify and track users in the same acoustic space and communicate private information only when it does not pose a threat to privacy (see fig. 18). However, tracking itself also poses a threat to privacy, as it is yet another form of private information extracted from the users. This places great trust in the local device’s ability to track and identify the authorization levels of the people present in the acoustic space. Such systems have, however, not yet been widely published.
Figure 19 presents another solution, which uses acoustic interventions to create physical spaces where the private information is, respectively, intelligible or inaudible [gardner1998_3daudio, pulkki2018parametric]. This acoustic effect can be achieved either as a sound zone through constructive and destructive interference between loudspeakers (see e.g. [donley2016, nielsen2018, WALLACE2022101285]) or as an audio spotlight through non-linear interference effects between ultrasound sources [yoneyama1983audio]. By choosing the spatial location where speech is intelligible and assuming we know where the target user is located, we can thus limit access to private information only to the target speaker.
Note that sound zones with destructive interference do receive a partial observation of the private message. Still, it is (hopefully) low-energy and distorted to the extent that it is unintelligible. The central challenges of this approach are to make the constructive sound zone large enough that it allows for small head movements and to make the destructive interference uniform everywhere else, such that there are no isolated points with constructive interference outside the desired sound zone. A benefit of this approach is, however, that it requires tracking of only the target listener, which is, while difficult, still much simpler than tracking all the people in the room.
IV-E Improving Performance
In many use cases, like False Activation, the actual culprit is the inadequate performance of the service. For example, if a wake word detector is incorrectly triggered (false positive), the system will start to listen to a conversation when it was not supposed to, thus breaching privacy. Since the inadequate wake word detector is thus causing a privacy breach, the best solution is to improve the wake word detector. The treatment thus addresses the cause rather than the symptoms thereof. Preserving privacy thus becomes an issue of systems design and implies using the privacy-by-design approach [Kalloniatis2009, van2011privacy].
This approach has, however, two principal challenges. First, improving performance often requires an increase in computational power and other resource consumption. This is financially costly and has an environmental penalty [patterson2021carbon]. In particular, balancing the computational load is easier on a cloud server, and thus, existing resources can be more efficiently used. Second, speech interfaces will always have occasional errors, even with improved performance. That is, the design of privacy-preserving speech technology must include multiple layers of protection, especially when requested operations have large consequences [hong2004privacy, backstrom2021edpbcomments]. Say, when the wake word detector is activated, then before any actions which would potentially breach privacy, the system could require that the speaker is in the same room, or the speaker identity is verified, or an extra confirmation step “Are you sure?” or similar. The intrusiveness of such additional protections should then reflect the severity of the potential breach such that the protections are not perceived as overly obtrusive and decrease the utility of the service.
Improving performance can also mean that the system requires additional functionalities. For example, in the case a second local device overhears Alice’s interaction with Bob (Discussion Leak) or the primary device (Speech Interface Leak), the required privacy protection is that the secondary device either 1. is aware that it is not the intended recipient of speech (speech analysis), or 2. notifies the user of its presence (user-interface design).
User Alice Local device Noise Remote service Refer to caption Refer to caption $+$ Refer to caption !
Figure 20: Protecting privacy by adding noise to private data, following the idea of differential privacy.
TABLE III: Dimensions of performance in privacy-preserving speech processing.
Performance
Objective Subjective
Privacy The consequences and
likelihood of attacks
Users’ perception and
experience of privacy
Utility Performance in the
downstream, trusted task
User experience (UX) and
quality of experience (QoE)
V Evaluating the Performance of Privacy-Preserving Methods
To evaluate the performance of any privacy-preserving methods, we need performance measures for both utility and privacy, corresponding respectively to the trusted and threat tasks (see fig. 2). With utility, we refer here to the usefulness of the service in its intended task. While the objective performance in both utility and privacy are prerequisites for any good speech application, to the user, it is central that their experience with the service or application is good. Correspondingly, performance can always be measured on two dimensions, privacy and utility, as well as objective and subjective (see table III). Performance with respect to objective and subjective privacy and utility, are presented in the following subsections (sections V-A, V-B, V-C and V-D), but the subjective experience of privacy is a sufficiently large and independent topic that it deserves its own section (section VI). We further need to define how the objective measures are applied in systems design to ensure privacy (see section V-E).
The evaluation protocol of the VoicePrivacy challenge is a state-of-the-art, practical example of evaluation of both privacy and utility [tomashenko2022voiceprivacy].
V-A Objective measures of privacy
Evaluation of objective privacy can be approached through two alternative approaches, through empirical methods or using theoretical concepts. The two categories emanate from different traditions and have their own benefits and areas of application, though the methodologies have some overlaps. More complete integration of the two categories would be highly beneficial, potentially providing the benefits of both categories, but at the time of writing, methods for such unification are not available. More comprehensive treatments of objective privacy metrics are provided in [nautsch2019preserving, noe2022towards].
V-A1 Empirical Privacy
Evaluation of privacy in, for example, the VoicePrivacy challenge [tomashenko2022voiceprivacy], is based on direct measurement of identifiability using signal processing and machine learning methods. The evaluation objective in such empirical privacy is to determine the extent to which an attacker can re-identify an anonymized speaker or anonymized attribute(s) of a speaker (see fig. 3).
Measures commonly used to analyze privacy in speaker recognition, -verification, and biometrics include [natgunanathan2016protection]:
  1. 1.
    The false rejection rate (FJR) (or false negative rate) is the proportion of times a sample from a speaker with a true attribute is not accepted.
  2. 2.
    The false acceptance rate (FAR) (or false positive rate) is the proportion of times a sample from a speaker without an attribute is accepted,
  3. 3.
    The equal error rate (EER) is the proportion of misclassified samples when the threshold of acceptance is adjusted to give equal FJR and FAR [maouche20_interspeech, ISObiometric]. An increase in EER means that the attacker has made more errors, and privacy has improved.
  4. 4.
    The application-independent log-likelihood-ratio cost function $C^{\min}_{llr}$ generalizes the EER by considering optimal thresholds over all possible prior probabilities and all possible error cost functions [maouche20_interspeech, brummer2006application].
  5. 5.
    The expected privacy disclosure from the Privacy ZEBRA framework [nautsch20_interspeech] quantifies the empirical cross-entropy (the leak magnitude) integrated over all possible prior likelihoods.
  6. 6.
    The worst-case disclosure also from the Privacy ZEBRA framework [nautsch20_interspeech] is the largest disclosure observed in a dataset.
Analysis of anonymized speech signals is straightforward since the above measures can be applied with conventional speech analysis methods to identify the anonymized attributes. Existing methodologies for evaluating speech analysis algorithms can thus be reused such that the evaluation is based on a body of prior research. On the other hand, empirical evaluation of any other form of signal, such as embeddings of machine learning representations or the compressed data of a telecommunications channel, requires analysis methods developed specifically for that data format. While the benefit is that such analysis directly evaluates the signal with any potentially confounding or obfuscation due to additional processing, the overhead of purpose-built analysis can be great.
A central downside of empirical privacy evaluation is that it is based on the current state-of-the-art, which can be surpassed by future technology. In other words, weaknesses in anonymization that are not (yet) known to the evaluator can compromise privacy. Moreover, it is widely known that the performance of machine learning models can be improved by making the model larger or using more data to train such models. A speech analysis model can thus be (likely) improved, not only by future developments but with this currently known methodology that only requires additional effort. Empirical evaluation can thus never give proofs, but just characterizations of the level of privacy.
As mentioned above, empirical evaluation of privacy using EER has been applied, for example, in the VoicePrivacy challenge [tomashenko2022voiceprivacy]. Both EER and $C^{\min}_{llr}$ are widely used in any speaker recognition [KINNUNEN201012]. The worst-case disclosure in a quantized bottleneck in speaker recognition has been studied in [vali2024porcupine].
V-A2 Theoretical Privacy
Theoretical evaluation of privacy aims to give assurances or proofs of the level of privacy based on notions of information theory. The idea is to quantify the new information about a speaker gained through an observed speech signal in terms of entropy and units of nits or bits.
In other words, suppose an attacker has prior assumptions about a speaker, quantified as a probability distribution $p(X)$ of the possible states $X$ of the speaker. The information content of that model is the negative logarithm $I(X)=-\log p(X)$ and its entropy the expectation thereof $H(X)=E\left[I(X)\right]$. The attacker then gains access to or makes an observation $Y$. This new information modifies the attacker’s beliefs about the speaker, expressed as the conditional model of the speaker $X$ given the information $Y$, or $X|Y$. Correspondingly, the posterior, conditional entropy $H(X|Y)$ quantifies the amount of information the attacker has about the speaker after (posterior) to the observation.
The leak of private information $\epsilon$ is then defined as the change in entropy
$\epsilon=H\left(X\right)-H\left(X|Y\right).$ (1)
The leak is thus equal to the mutual information of $X$ and $Y$.
A related concept is the log-likelihood ratio (LLR), defined as the logarithm of the ratio of likelihoods $p(Y|H_{k})$ of the observation $Y$ conditioned with the hypotheses that either the sample is $H_{1}$, or is not $H_{0}$, the target speaker [nist2021SRE]
$LLR(s)=\log\left(\frac{p\left(Y|H_{0}\right)}{p\left(Y|H_{1}\right)}\right)=I(Y|H_{0})-I(Y|H_{1}).$ (2)
A large $LLR$ indicates that $H_{0}$ is more likely than $H_{1}$. Comparing to eq. 1, we can conclude that where eq. 1 is the expectation of information content, eq. 2 quantifies information content explicitly.
Another important and related formulation is differential privacy, which quantifies how the amount of information about a speaker’s affiliation to a particular group changes due to an observation [dwork2006differential, dwork2008differential, nautsch2019preserving]. In particular, suppose an algorithm $M()$ is applied on a set $D$ to give an answer $M(D)=S$. The leakage of information is then new information gained through a slightly different dataset $D^{\prime}$ as
$\log\left|\frac{p\left(M(D)=S\right)}{p\left(M(D^{\prime})=S\right)}\right|\leq\epsilon.$ (3)
Leaks can also be formulated as a function of the groups’ sizes where the speaker is known to belong before and after an observation. Suppose we know that a speaker $x$ belongs to group $x\in A=\{a_{1},\dotsc,\,a_{M}\}$ with $M$ members. The speaker can be any one of the group members with equal likelihood $p\left(x=a_{k}\right)=\frac{1}{M}$. The identity within this group, which is unknown to the attacker, then has an information content of $\log M$ nits. If an attacker somehow manages to identify that the speaker belongs to a subset $B\subset A$ of size $N$, then the remaining information content unknown to the attacker is $\log N$ nits. The reduction in uncertainty is thus the leak magnitude
$\epsilon=\log M-\log N=\log\frac{M}{N}.$ (4)
A leak is thus a function of relative group size before and after an observation.
It is arguably more informative for users to state the absolute group size within which they remain anonymous. If, through the processing of an observation, an attacker can reduce the group size within which the user remains anonymous to $k$, then we say that the process is $k$-anonymous [sweeney2002k].
A central issue with theoretical metrics is that they are based on mathematical definitions, whereas speech is a natural signal that does not submit to exact formulas [nautsch20_interspeech]. To use these formulas, we must, therefore, first choose models to describe the distributions of observations and then estimate their parameters. Both steps lead to inaccuracies that dilute the evidentiary power of the proofs of privacy.
Theoretical methodology for privacy has been applied in, among others, differentially private speech recognition [shoemate2022sottovoce], emotion recognition [feng22b_interspeech], speaker anonymization [shamsabadi2023diff], and federated learning in acoustic sensor network [el2022differential]. As theoretical proofs of privacy have obvious advantages, the adoption of this methodology will likely increase.
V-B Subjective evaluation of privacy
In speaker recognition, computers are now on par or better than humans  [hansen2015speaker] and are not fooled by voice mimicry (i.e., voice actors) even when humans struggle to do so [vestman2020methods, vestman2020voice]. We can thus expect that computers are better at evaluating privacy in well-defined tasks. However, there are at least two caveats. First, in identifying private information, human listeners can potentially use properties or features of speech signals other than those computers use. Subjective listeners can thus potentially discover new weaknesses that computers miss. Second, in scenarios where humans are the potential attackers extracting information (as in threats 3, 5, and 7 in fig. 1), the true measure of privacy is quantified by subjective listening. Thus, it is often prudent to complement objective privacy metrics with subjective tests.
In the subjective evaluation of privacy, the listening task should be defined to match the threat. For example, if the threat is re-identification, subjects should be tasked with identifying subjects from anonymized speech signals. Similarly, if the threat is speakers’ attribute inference, the subjects should evaluate the extent to which they can identify such attributes. For a practical implementation of such subjective evaluation of privacy, see [tomashenko2022voiceprivacy].
V-C Objective performance in downstream tasks
The utility of a privacy-performing system can be quantified by measuring the performance of the downstream, trusted task. By the downstream task, we here refer to the intended functionality of the speech service. The downstream task is, in the functional flow diagram, generally implemented after privacy-protection elements, which is why it is known as a downstream task. In some cases, we might not trust the provider of the downstream service and need to apply privacy protections specifically because we have no trust. The concept of a “trusted task” is, in such cases, misleading, and the concept of a “downstream task” can be preferred.
Metrics used for quantifying utility depend on the application. In applications where a computer analyzes the speech signal, we quantify utility by measuring the accuracy of analysis. For example, in automatic speech recognition, where the computer extracts the text content of a speech signal, a commonly used measure is word error rate (WER), which is the proportion of correctly identified words [rohit2024asr]. By measuring the word error rate with and without privacy-preserving processing, we can determine how much privacy-protections impede utility.
V-D Subjective performance in downstream tasks
If a human user is to listen to the output signal, then that user is the ultimate measure of performance [zacharov2018sensory]. For example, if we implement speech enhancement to reduce the perceived noise in a signal, then a user can subjectively evaluate how close the enhanced signal is to the noiseless original [vincent2018audio]. Such perceptual evaluation is best implemented using subjective listening test standards such as P.800 or MUSHRA [P800, MUSHRA]. Similarly, if privacy protections degrade a speech signal, we can readily quantify such degradations by listening tests.
Subjective listening tests are, however, expensive, time-consuming, and noisy. Individual subjects, especially those without training (known as naïve subjects), are typically unable to repeat a sequence of tests such that results would match. Subjects are also slow since they need to listen to samples one at a time, at real-time speeds, require time to answer between each sample, a multitude of subjects need to do the same test to gain statistical confidence in results, and tests need oversight by trained personnel. Automated tests, known as objective measures of perceptual quality, are an alternative to such subjective tests. They estimate the outcome of subjective listening tests by computational means. For example, a classic objective measure known as PESQ estimates the outcome of a P.800 test [rix2001perceptual]. Later improvements include POLQA and ViSQOL [beerends2013perceptual, chinen2020visqol]. Such objective measures can be used to automate quality testing at a lower price and faster speed so that repeated estimates with the same input give the same result. Observe that researchers should carefully review the specifications of such algorithms to determine whether they apply to the problem at hand to ensure reliable results.
V-E Privacy-preserving systems
Even if individual modules of a system can assure privacy, their interaction in a system can expose privacy. Therefore, the privacy-preservation performance of processes involving speech or other biometric signals must be evaluated on a systems level. Such evaluation uses a concept of biometric reference, which is a data record that stores an attribute of the user, such as the speakers’ identity. New samples are compared to such biometric references to find matching attributes. The three key concepts in such evaluation are [ISOIEC224745_2022, nautsch2019preserving]
  • Unlinkability – comparing two or more biometric references will not reveal whether they are from the same or different persons.
  • Irreversibility – information about the biometric sample cannot be derived from the biometric reference.
  • Renewability – multiple, unlikable biometric references can be created from the same biometric samples and used to identify the person without revealing information about the biometric samples.
Together, unlinkability and irreversibility thus minimize the amount of information leaked if biometric references are revealed. At the same time, renewability ensures that leaked references can be revoked and new ones created to replace them.
VI Perception and Psychology of Privacy
People tend to have a deep sense of ownership of some things, both material and immaterial. In particular, people have a feeling of ownership toward information about themselves [cichy2014extending, dawkins2017psychological]. Such feelings are, to some extent, detached from the material consequences of breaches in privacy. For example, even if publishing an audio recording revealing a secret intimate relationship would not have direct economic consequences, it can cause psychological damage and harm personal relationships. The effect of threats to privacy then necessarily becomes a question of psychology and social sciences. To qualitatively or quantitatively measure such effects, we also need subjective user tests.
Concurrently, people have well-established social rules regarding human-to-human privacy [petronio2002boundaries]. Moreover, users also tend to anthropomorphize technology, that is, treat devices and services as if they were human [nass1993anthropomorphism, cornelius2021acceptance], such that they will likely assume that devices apply human-like social rules to privacy. Observe that such social rules are related to human-like behavior and performance. That does not reveal whether they are applicable to the super-human performance that computers can possess, such as the ability to integrate information over massive databases and to retain accurate records of events long past. How machines should then behave, especially with respect to super-human capabilities, is then not only a question of user-interface design but of moral psychology [thompson2021machine]: We need discussions on a societal level of how automated services should behave and how they are allowed to behave (see also section VIII).
VI-A Perception and Experience of Privacy
Irrespective of whether a system actually preserves privacy or not, some users can perceive the system as threatening, and others can be oblivious to the threats it poses [pepino2020distrust]. Clearly, users do not like using services they perceive as threatening. Understanding how people perceive privacy is interesting in its own right, but such understanding is essential in the design of effective user interfaces.
Perception and experience of privacy can be approached in two alternative ways. We can make user studies where people interact with either machines or each other [leschanowsky2024review]. The distinction is important in the sense that human-to-human interaction relies on social rules that are well-established and developed over a long time. We can thus expect them to be stable over time, whereas human-computer interaction is continuously evolving as people learn more. Moreover, by using studies of human-to-human interaction, we can learn effects related to human-like performance but can probably not rely on observing effects related to super-human performance. Studies of human-to-human interaction are, however, always a proxy if the actual target is to design human-to-computer interfaces [leschanowsky2024review].
Studies of human-computer interaction thus characterize the desired phenomenon directly. The compromise is, however, that people’s understanding of privacy with devices might not reflect the true level of privacy, and conclusions made based on the users’ opinions might then not reflect their true preferences and those preferences are moving targets. As people gain more experience with and as they learn more about technology, their attitudes, and perceptions of it change [leschanowsky2024review].
Despite these shortcomings, both types of experiments are essential for improving our understanding of privacy and for improving technology. For example, human-to-human studies have revealed that people experience privacy differently in different acoustic environments; a noisy cafeteria can be better at masking sounds and defending against eavesdropping than a reverberant hallway [zarazaga19_interspeech, das2021thesis, zarazaga2022thesis, backstrom2021intuitive, Leschanowsky2020] and multiple-bed patient rooms in hospitals have privacy-concerns [reinten2017speech]. Similarly, human-machine studies have revealed that when a chatbot actively communicates choices related to privacy it improves users’ experience of privacy [bruggemeier2022perceptions], voice interfaces with unknown features cause fear in users, and reduces retention of services [yeasmin2020privacy, Lau2018, lutz2021privacy] and breaches in privacy are highly detrimental to the trust in services and reduce users’ willingness to use the services, but such trust can be rebuilt [Kurz2021].
VI-B User Interface Design
For usable privacy, it is important that information about privacy is readily provided, changes in the extent of privacy are promptly notified, and that users have control of the level of privacy [bruggemeier2022perceptions, so2021voice, lewis1998designing, Lau2018, Kurz2021, dhiya2023privacy, iachello2007end, sivaraman2018smart]. As a parallel to speech interfaces, visual interfaces can use lights or icons for monitoring and tactile interfaces for controlling the level of privacy. While sound can be used to monitor system status with sonification [hildebrandt2016continuous, iber2021auditory], this is not in widespread use within signaling of privacy. It is a compromise between filling the acoustic space with information and the tendency of the human auditory system to block out monotonous sounds [pulkki2015communication]. To be effective, the sound should be perceivable when the user consciously wants to check the privacy level and changes in privacy level should evoke correspondingly and appropriately large changes in the soundscape that users consciously register such changes. Furthermore, user interfaces should give correct information about and allow controlling the privacy level [Lau2018, backstrom2020privacy].
In any case, it is imperative that services are designed such that they reflect the true extent of privacy. Observe that it is likely possible to design systems that communicate an advanced level of privacy even when the system does not respect user privacy. In fact, service providers may have short-term incentives to follow such approaches to design as long as they improve overall user satisfaction. However, such approaches to design are known as dark patterns or deceptive design patterns, and they are considered unethical [willis2020deception]. Through deception, such dark patterns lull users into believing they are safe when, in fact, they are abused. Therefore, good design of privacy should actively communicate and enable control of the true extent of privacy. Such design practice is not only ethical but also rewards service providers in the long term by improving retention of those services [Kurz2021].
VI-C Privacy and Trust
The level of privacy provided by a speech service is intimately linked to the level of trust users experience in that service [leschanowsky2024review]. In particular, if users experience that a service is detrimental to their privacy or, more generally, to their well-being, they will obviously not trust it. We thus discuss trust mainly from a psychological perspective, unlike in security research and cryptography, where trust has a specific and unambiguous mathematical definition [balsa2022cryptography].
While the meaning of trust may seem intuitively clear, it is however difficult to define exactly. A classic definition is [deutch1973resolution]
Trust: confidence that [one] finds what is desired rather than feared.
This definition highlights the psychological nature of trust [simpson2007psychological]. From the current engineering perspective, this translates to the statistical confidence that the outcome of a process is useful. Here, usefulness refers to what the user, subjectively perceives as useful. Implicitly, this furthermore requires that we predict the outcome of that process in the form of a statistical distribution of possible outcomes and that the expectation over the possible outcomes is that the process is useful. Another way of expressing this is to say that trust requires dimensions of competence, integrity, and benevolence [Chen2003, Xie2009].
Observe the semantic difference between a system that is “benevolent” and one that is “useful”. The casual use of the concept of trust seems to require that the actor is indeed benevolent. However, it is possible to imagine a system with malicious objectives concerning Alice but produces side effects that consistently benefit Bob. Bob thus could potentially trust the system to act beneficially even if the system is not benevolent towards him. It is furthermore questionable whether we can assign an “intent” to a computational system due to its close association with psychology. The concept of benevolence is thus ambiguous in the context of speech technology.
It is useful and ethical to design speech services that are trustworthy. This means that systems should be such that it would be logical for users to trust them [hardin1996trustworthiness]. Speech technology should thus be designed to have useful outcomes with statistically significant confidence, especially with respect to privacy.
A practical approach for reducing the likelihood of negative outcomes is to build multiple layers of checks and verifications, similar to multi-factor authentication [ometov2018multi]. In fig. 4, this can be implemented by applying layers of privacy protection in each successive connection and device as well as in their internal modules and components, as described in section IV-E. Activation of an undesired functionality of the system or leaking private information would thus require that every subsequent module fails simultaneously, which is unlikely in most scenarios.
Conversely, users likely experience most trust in devices and services that are “close” to them and over which they exert control. Thus, it seems obvious that users will tend to trust more those devices they have in their possession or homes, rather than remote cloud services.
VII Content Categories and Applications
To give a complete picture of the implications of privacy in speech technology, this section provides a brief discussion both on categories of private information, complementing table I as well as the application of such information. Observe that this is not a complete list (nor is table I) of the content nor of application categories. The purpose is merely to provide a characterization of work done and challenges in the research field, with accompanying references.
First, note that while all presented categories are (potentially) private information, all sustained information (in the sense that it does not change over a “long” time) can potentially be used to identify a speaker. For example, while the current emotional state cannot alone identify a speaker, the tendency to display emotional states can aid in identification. However, where we want to verify that a speaker is who they claim to be, we can use only information that cannot be willfully changed. That is, speech style is a particular example of a property that a good voice actor can freely choose, making it “easy” to change also for fraudulent purposes. Using variable attributes, or their distributions, to identify or verify an identity is thus likely to be complicated.
Recognizing the speaker’s identity then becomes the natural starting point for studies in privacy. We can attempt to recognize who is speaking (speaker recognition), verify whether a speaker has the claimed identity (speaker verification), clustering audio to segments with a single speaker (speaker diarization) and we can develop methods for deceiving identity (spoofing), e.g. [chouchane21_interspeech, granqvist20_interspeech, mawalim20_interspeech, mawalim20_interspeech, nautsch19b_interspeech, nelus19_interspeech, nelus19b_interspeech, noe21_interspeech, perero2022x, zewoudie2021federated, zewoudie2021use, teixeira22_interspeech, treiber2019privacy, nelus19_interspeech, nelus19b_interspeech, BAI202165, han2020voice, KINNUNEN201012, hansen2015speaker, parthasarathi2013wordless, pathak11_interspeech, parthasarathi11_interspeech, pathak2012ppsa, jimenez2017two, portelo2014bprivacy]. Similarly, by voice conversion we can anonymize a speaker identity by replacing it with a random identity (anonymization) or a specific one (pseudonymization), e.g. [noe20_interspeech, pierre22_interspeech, prajapati21_interspeech, shamsabadi2023diff, srivastava20_interspeech, Srivastava2022taslp, mawalim20_interspeech, maouche20_interspeech, Fang2019]. Speaker characterization is the natural complement of speaker idenfication [nautsch2019preserving, nguyen2010automatic]. Such methods related to speaker identity can be applied for example to recognize the user of a device, verify a customer at a bank, or as a voice avatar in online gaming. Similarly, anonymization and pseudonymization can be used to hide the identity from the public media and gaming.
Speech recognition, as in speech-to-text, is probably the largest sub-area of speech research. With respect to privacy, it contains two obvious challenges for privacy. We can try to limit side-information, such as eliminating all non-text information from the data stream, e.g. [shoemate2022sottovoce, nguyen2023federated, aloufi21_interspeech, garg20b_interspeech, shah21_interspeech, srivastava19_interspeech, tomashenko2022privacy], or we can use natural language processing to anonymize the text content, e.g. [liu22_interspeech, maouche2022enhancing, adelani20_interspeech, novotney21_interspeech, mahendran2021privacy, habernal2021differential].
Privacy is even more important in always-on applications like wake-word detection, i.e. when the interface is triggered by a specific keyword like “Computer” in “Computer, lights-off”. This application is more sensitive exactly because it is always on as users cannot choose when data is processed [koppelmann21_interspeech, koppelmann22_interspeech, stoidis22_interspeech, yang21_interspeech]. The always-on characteristic is also prominent in assisted-living applications, which can ”monitor people’s daily exercises, consumption of calories and sleeping patterns, and to provide coaching interventions to foster positive behavior” [haider19_interspeech]. Such ambient voice interfaces are often implemented through acoustic sensor networks which pose their own challenges [koppelmann22_interspeech, oconnor20_interspeech, malekzadeh2018replacement, zewoudie2021use, zarazaga2020acoustic, nelus2021].
Speech enhancement refers to removing background noises and distortions from the desired speech signal (private message) [benesty2008springer]. This task can have an impact on privacy in two ways. First, the recording environment can reveal private information about the speaker, and second, background noises, like a competing talker, can contain private information [zarazaga21_interspeech, rech2022multi]. Attenuating background noises and competing speakers, as well as removing reverberation, can thus improve privacy. Again, while using a sensor network to capture speech can improve utility, it introduces novel threats as well [nelus2021, zarazaga2022thesis, hendriks2013ppdistributed].
Speech is typically considered to be biometric information in its entirety, including all information categories in table I as such categories can be used to identify a person [adler2006towards, nautsch20_interspeech, nautsch2019preserving, ISObiometric, edpb2021guidelines]. Some, however, limit the term biometric to refer only to physical and behavioral characteristics. Such properties include health [natarajan2021prioris], emotions [feng22b_interspeech] and gender identity [nelus2018itg, stoidis21_interspeech, stoidis22_interspeech, nelus19_interspeech], each warranting their own treatment.
VIII Legal and Societal Landscape
[…]
While we have not (yet) seen breaches related to speech technology with consequences of comparable magnitude, these parallels highlight the potential effect of breaches. Scandals directly related to speech technology include eavesdropping on private persons by employees and contractors of service providers [Lynskey2019GuardianPrivacyScandal, hern2019, day2019].
Section VI-B also makes the point that service and technology providers have clear short-term incentives that conflict with users’ preferences for privacy. We can easily find examples where users do not have real choices in protecting their privacy. For example, suppose all friends and family of a user use a particular platform for social and voice interaction even if its privacy configuration is inadequate. The individual user faces then the choice of either disconnecting from their social network or compromising privacy. This applies to all users with a sufficiently large portion of their communities participating in that platform. This is a variation of the prisoner’s dilemma, where no single user has sufficient incentive to transition to a solution that would be best for everyone.
On the individual level, the examples of potential exploits in table II cause clear damage to individuals, including psychological harm in particular (e.g. [natarajan2021prioris]). Additional dangers include stalkers [lu2019abusers, hulsen2022stalkers, avast2021stalkerware]. While such individual damage is “small” on the societal level, their prevalence makes their joint impact significant [levy2020intimate].
These examples demonstrate the inherent need for society-level regulation of speech technology with respect to privacy. Governments have already responded to this need, with the European Union spearheading the process with the General Data Protection Regulation (GDPR) [eu2016gdpr, nautsch19c_interspeech], with the State of California following soon thereafter with the California Consumer Privacy Act (CCPA) of 2018 [california2018ccpa]. While these laws cover only a small percentage of the global population, as cloud services typically operate globally, they need the capability to follow local laws. In many cases, it can be easiest to apply the strictest laws on all users, such that the most strict laws benefit the privacy of all users. Service providers have, therefore, widely adopted the requirements of GDPR and CCPA, and that has likely had a large impact also on users outside the scope of these regulations.
With respect to regulation, an important consequence of the objective measures of privacy in section V is that our tools and measurements will give as an output only statistical characterizations of the accuracy of estimated attributes, but they can never give absolute confidence. This is in stark contrast with the concept of unique identifiability used in legal documents, such as the General Data Protection Regulation (GDPR) by the European Union [eu2016gdpr], which does not explicitly leave room for statistical uncertainty. This is reflected for example in the Guidelines for virtual voice assistants by the European Data Protection Board, which states that: [edpb2021guidelines, page 13, §31]
… voice data is inherently biometric personal data. As a result, when such data is processed for the purpose of uniquely identifying a natural person … the processing must have a valid legal basis …
This leaves the interpretation open. It is possible to argue that it is never possible to obtain absolute confidence in speaker identification such that the GDPR is never triggered. It is also possible to argue that all voice data contains personal information that can be used to uniquely identify a person, such that all processing must have a valid legal basis. Both interpretations lead to absurdity, which suggests that the truth must lie somewhere in the middle. In fact, the GDPR in practice requires (see [edpb2021guidelines, page 4]) that the design process of voice assistants includes a data protection impact assessment, where the risks and consequences are evaluated such that the designer can take appropriate precautions to preserve privacy. Authors of the GDPR are thus clearly aware that it is impossible to give absolute guarantees of privacy but that the impact assessment (i.e., objective measures of privacy) must necessarily be based on statistical measures, even if such measures have not been defined.
However, even in the absence of legal requirements, the current author would encourage all researchers, developers, data analysts, project managers, and policymakers to perform impact assessments in data projects, data management, and data policies, anytime when there is or could be the slightest concern that the dataset imposes risks to privacy. A practical tool to aid impact assessment is the “Data Ethics Decision Aid (DEDA)” that formalizes and enumerates all the questions that a representative group of stakeholders should review, discuss, and consider [franzke2021data].
While governments are in the process of regulating privacy, corporations and non-governmental organizations have also realized that proper privacy is an opportunity. For example, the Open Voice Network seeks to develop and standardize open technical standards and ethical guidelines for voice assistance [ovn2023] and the MyData Global seeks to help people and organizations to benefit from personal data in a human-centric way [mydata2023, poikola2020mydata]. Within the research community, the author of this paper has been involved in establishing a special interest group within the International Speech Communication Association (ISCA) devoted to “Security and Privacy in Speech Communication” [spsc2023]. It is as far as we know, the world-wide largest community focused on this topic.
IX Discussion and Conclusions
The quality and use of speech interfaces have increased rapidly in recent years. As with any new technology, the rapid progress has also revealed dangers, particularly the threats to privacy it demonstrably poses. Unprotected users are exposed to threats like stalking, algorithmic stereotyping, harassment, and price gouging. Researchers, service providers, and governments thus have the impetus to protect the users, not only because it is ethical but also because it makes for better products and long-term business.
This paper is a tutorial on privacy for speech technology. Its most notable novelty is an exhaustive categorization of threats (see figs. 1 and III). Protections against those threats are further categorized according to whether they relate to the private message or side information. The pertinent difference is that transmitting a private message is the whole purpose of communication, and there is not much we can do to protect it other than encryption. With side-information, that is, all the other information that is bundled into a speech, like health status and gender identity, we have a much larger arsenal of protections. The primary approach is, however, to remove as much of the side information as possible as early as possible. As the private message is all that communication we need, all side-information should be removed to the extent possible. Such removal rapidly, however, demonstrates that paralinguistic information like speech style is often very useful in conveying the intended message. It is thus not always clear what constitutes a legitimate private message.
The first conclusion from this paper is that the range of possible threats to privacy is vast. Each agent – be it human or device – participating in an interaction as well as the acoustic pathways and network connection through which they are connected, is a potential attack surface. Any actor that can interact with the other agent or listen to the connection is a potential eavesdropper. Since we define privacy as a scenario where an agent is authorized with some access but over-exceeds that authorization (intentionally or inadvertently), we cannot just cut connections but need more refined designs and methodologies. We thus need to dynamically adjust access according to need. Conversely, systems must actively monitor the privacy status to determine appropriate actions.
Second, we find that privacy and ethics are largely overlapping challenges. Our ethical values govern our preferences for privacy. Most potential breaches of ethics in speech technology are based on breaching privacy. That means that we need a society-wide ethical discussion about what is allowed with respect to user privacy. Such discussions are needed to prevent a Cambridge Analytica-style scandal for speech technology [heikkila2023cambridge_ai].
A third implication of this paper is that, while research in this field has started to accelerate only recently, there is already a substantial body of research available. The research is not, however, mature but in a phase of rapid development, and there are important sub-areas that have not yet seen much work. This makes it a fruitful area for research as we can expect important understanding to be discovered in the coming years.
Particular research questions where the author sees an urgent need for and expects to see new results include:
Consent
While management of acquiring informed consent has established traditions and best practices for most interface types [friedman2000informed], speech, audio, and ambient systems are notably unique. Namely, acoustic information is a time-varying stream. Reading out a pages-long consent form before an interaction can start is clearly much too obtrusive and unnecessarily detailed. Privacy requirements also vary over time. Consent should thus be acquired on an actual need basis. In addition to being more usable, it would also make choices better connected to the actual needs since consent is acquired only once it is actually needed.
Metrics for Streaming
The available theoretical metrics reflect privacy with respect to a finite dataset, whereas speech is an open-ended stream of data. The consequence is that, in theory, we can resolve any private attribute or identity, provided that it has a unique probability distribution and we have a sufficiently long observation. We would thus need methodologies for characterizing the effect of the length of observation on privacy.
Metrics for Out-of-category Information
The metrics discussed in section V are all related to specific categories of private information and in particular, we can provide protection only to identified threats. For example, we can measure the threat to privacy related to health information, but that does not say anything about the threat related to information about ethnic background. We thus need methods for evaluating privacy jointly with respect to all categories of private information except the private message.
Future-proof Metrics
Empirical metrics are generally based on a model of the signal or the attacker. The metrics are thus subject to change when those models are improved in the future, and it will likely expose new threats. Though it is likely difficult, it would be extremely useful if we could characterize, for example, privacy threats as a function of computational complexity, similarly as cryptography provides assurances over security based on the difficulty of solving certain difficult computational problems. Where cryptographic security is often seen as a threshold of computational complexity, however, for privacy in speech applications, it would be beneficial to also assess the leak magnitude (e.g. in bits) as a function of computational complexity. Over time, improvements in efficiency reduce algorithmic complexity, and as hardware with better capabilities becomes available, this would allow characterizing how protections will survive the test of time.
Attack and attacker characterization
The biometrics community has a long tradition of characterizing attacks as, for example, opportunistic and motivated attacks, as well as in terms of the potential rewards of such attacks [roberts2007biometric]. Characterization of speech-specific attacks has received much less attention, though figs. 1 and III in this paper are attempts in that direction. Further work is needed to gain privacy-protections on a systems level.
Multi-user interaction
Privacy research is categorically focused on personal and user-centric privacy. However, speech is, by definition, communication between multiple agents, and when exposed, threatens all participants simultaneously. This is not an issue from a legal point of view, because privacy protections apply to all individual users equally. However, from an authorization and consent management perspective, this is an underappreciated issue. If user A records a discussion with user B, then both clearly have some level of ownership and privacy requirements on that recording. Another case is smart technology with multiple users, like smart TVs; even if one user has consented to data collection, that does not mean that others would agree. We do not yet have any widely accepted standard approaches for handling privacy, ownership, and consent in such multi-user scenarios.
Disentangelement
If we could disentangle all categories of speech information as in fig. 14, it would be easy to anonymize each category to an appropriate degree. This approach thus seemingly solves all our problems. The issue is that we do not yet have sufficiently sophisticated methods for such disentanglement. The difficulty in developing disentanglement algorithms is that information categories in table I are vaguely and heuristically defined and there is significant overlap between them. We cannot even demonstrate that this would be a complete list of information categories. Without exact definitions of those categories, we have no hope of developing methods for them. An alternative approach is to use representation learning methods to create unsupervised clustering of information categories, thus circumventing the need to explicitly define categories. The compromise is that we cannot guarantee that the learned representations correspond to heuristically meaningful categories. Still, since disentanglement is the ideal solution, that should continue to be a central focus of research.
Perception, Experience, and Design of Privacy
Most of the speech-specific research on privacy has focused on privacy-preserving processing and systems structures. This is useful because it is the mandatory prerequisite for privacy-preserving technology. However, as discussed in section VI-B, users’ experience of services is to some extent independent of the objective level of privacy. We need many more user studies on how, for example, voice characteristics and word choices influence trust, how the privacy level can be monitored during interactions and how changes are notified, how the environment and content of interaction influence user experiences, etc. By improving the user experience with respect to privacy, we are likely to improve user satisfaction and retention of the overall service, while also improving the service objectively.
In conclusion, threats and breaches of privacy have significant negative consequences on individual, societal, ethical, and economic levels. While further improvements in smart technology are expected to improve the utility of the technology, it likely also introduce new threats. The protection of privacy in speech technology has thus been important already for a long time and the importance is increasing. Fortunately, research in the area has picked up speed and this tutorial presents the most important concepts, approaches, and methodology. It is, however, likely that fundamental results and new technologies will be introduced in the near future. This is thus an exciting time for researchers in the area.
Acknowledgments
A majority of the content in this paper has risen directly from discussions with a long list of colleagues, who were far too many to include as co-authors. These colleagues include, but are not limited to, in alphabetical order, Chris Brzuska, Mads Græsbøll Christensen, Sneha Das, Nicholas Evans, Johannes Fischer, Francisco Florez, Florin Ghido, Meeri Haataja, Ivan Habernal, Aki Härmä, Dorothea Kolossa, Michael Laakasuo, Martha Larson, Anna Leschanowsky, Rainer Martin, Joachim Meyer, Sebastian Möller, Andreas Nautsch, Birgit Popp, Nitin Sawhney, Ingo Siegert, Stephan Sigg, Rech Silas, Isabel Trancoso, Sanna Toropainen, Emmanuel Vincent, Jennifer Williams, and Pablo Pérez Zarazaga. I will remain forever grateful.
[…]
[Uncaptioned image] Tom Bäckström D.Sc. (tech.) is an associate professor at Aalto University, Finland (2019-). He obtained his Master’s and Doctoral degrees at Helsinki University of Technology (the predecessor of Aalto) in 2001 and 2004, respectively. During his time at International Audio Laboratories Erlangen, Germany (2008-2016), he made contributions to several international speech and audio coding standards such as MPEG USAC and 3GPP EVS, and became a professor (W2) at Friedrich-Alexander University Erlangen-Nürnberg (FAU) (2012-2016). Before his current position, he was a professor practice at Aalto University (2016-2019). He was the initiator and chair of ISCA SIG “Security and Privacy in Speech Communication” (2019-2022). His current research interests include privacy, coding, enhancement, and transmission of speech as well as machine learning.

### v3

@@
 Speech is a mode of communication and thus inherently contains a wealth of information (see table I).
-Communicating information already known by the receiver is pointless, and efficient communication will thus mainly contain information that is not widely known or which is private.
-In addition, speech contains also a wide range of side information like the state of health and emotions, as well as physical, psychological, and social identity, most of which is private information.
-Finally, speaking is a dynamic interaction between two or more speakers.
-Dialogues thus also contain information about the relationship between participants such as their level of familiarity, affiliation, intimacy, relative hierarchy, shared interests, and history.
+In addition to the desired information of the communication, speech thus also contains a wide range of side information like the state of health and emotions, as well as physical, psychological, and social identity, most of which is private information.
+Moreover, speaking is a dynamic interaction between two or more speakers.
+Dialogues thus also contain information about the relationship between participants, such as their level of familiarity, affiliation, intimacy, relative hierarchy, shared interests, and history.
 From an information content perspective, we can thus expect that privacy is a multifaceted issue in all areas of speech technology.
-TABLE I: Private information potentially identifiable from speech signals, and the extent to which they are sustained over time and can be willfully controlled, summarized by and reflecting the personal judgment of the author.
+TABLE I: Private information potentially identifiable from speech signals, and the extent to which they are sustained over time and can be willfully controlled [1, 2, 3, 4].
 Category Examples Permanence Control
-Biological Body characteristics Sustained No
-State of health Variable No
-Level of intoxication Variable Yes
-Psychological Emotions Variable Partly
-Intelligence Sustained No
-Education, skill Sustained Yes
-Gender identity Sustained No
+Biological Body characteristics Long-term No
+State of health Short-term No
+Level of intoxication Short-term Yes
+Psychological Emotions Short-term Partly
+Health Mixed Partly
+Intelligence Long-term No
+Education, skill Long-term Yes
+Gender identity Long-term No
 Message Text, emphasis, style,
 expression
-Variable Yes
-Mannerisms, context Variable Partly
-Language
-choice & skills
-Partly Partly
+Short-term Yes
+Mannerisms, context Short-term Partly
+Language choice Short-term Yes
+Language skills Long-term Partly
 Affiliation Ethnic, national,
@@
 political, etc.
-Sustained Partly
+Long-term Partly
 Relationship
@@
 attraction, intimacy
-Sustained Partly
-Has met with person Variable Yes
+Mixed Partly
 Physical
@@
 reverberation, location
-Variable Yes
-Hardware
-used
-Sensor type & manufacturer Variable Yes
+Short-term Yes
+Hardware &
+software used
+Sensor and software
+type & manufacturer
+Short-term Yes
 TABLE II: Practical examples of threats related to private and sensitive information that is conveyed by people’s voices.
@@
 Private health problems, and romantic affairs could be detected in the voice and used for blackmail or made public against a user’s wishes.
-Algorithmic stereotyping Recommender systems based on voice can become biased with respect to age, identity, religion, or ethnicity, in ways that are nearly impossible to monitor.
+Algorithmic stereotyping Recommender systems based on voice can become biased with respect to age, identity, religion, or ethnicity in ways that are nearly impossible to monitor.
 Harassment, inappropriate advances Users in chat rooms or virtual reality could be automatically singled out by gender or opinions, making them a target for unwanted attention and harassment.
 Fear of
-monitoring
-The subjective feeling of being continuously monitored may cause psychological damage. It may also stifle political expression damaging democratic societies.
+monitoring, loss of
+dignity
+The subjective feeling of being continuously monitored may cause psychological damage and stifle political expression damaging democratic societies.
+Users may feel they lose control of personal aspects of their lives.
 Studying and improving privacy in speech technology is important because breaches in privacy can have serious consequences.
@@
 It is then ”merely” a question of how such minimization is achieved, what information is relevant, and how to determine what the user wants (provide information and enable control of services and threats).
-With respect to threats, media attention is often focused on breaches which have large economic consequences [Lynskey2019GuardianPrivacyScandal, hern2019, brewster2021].
+With respect to threats, media attention is often focused on breaches which have large economic consequences [5, 6, 7].
 While such breaches are important by themselves, the media attention introduces unfortunate biases in two ways.
@@
 To maintain users’ trust, it is therefore important to both uphold the users’ actual privacy but also design systems to actively and frequently communicate the level of privacy, including known threats, their current status, and measures taken to protect against them.
-Second, while a single breach in a large service can have large economic, psychological, societal, and legal consequences, small crimes are so common that their joint effect is comparable or larger in size [levy2020intimate].
+Second, while a single breach in a large service can have large economic, psychological, societal, and legal consequences, small crimes are so common that their joint effect is comparable or larger in size [8].
 With speech technology, such ”small” crimes include stalking, extortion, harassment, humiliation, and inappropriate advances (see table II).
@@
 This paper presents a tutorial overview of privacy in speech technology that covers a wide range of threats, methodologies, and algorithms.
-Analysis of threats, however, demonstrates that while attack surfaces have great variety, the threat models are similar (see section III).
-Protections against those threats have five distinct categories of approaches, isolating target information by removing private information, limiting access to private the message through cryptography, reducing centralized threats through non-centralized processing, limiting access to reproduced audio, and improving overall system performance (see section IV).
-To quantify the extent of protections, we further need methods for evaluation of privacy and utility (see section V).
+The range of applications and content types where privacy-preserving processing methods are needed is vast, and section II presents a brief overview thereof, and section III defines the vocabulary.
+Analysis of threats demonstrates that while attack surfaces have great variety, the threat models are similar (see section IV).
+Protections against those threats have distinct categories of approaches, isolating target information by removing private information, limiting access to private the message through cryptography, reducing centralized threats through non-centralized processing, limiting access to reproduced audio, improving overall system performance, and design principles from theoretical frameworks (see section V).
+To quantify the extent of protections, we further need methods for evaluation of privacy and utility (see section VI).
 While objective evaluation quantifies the actual level of privacy, users’ impressions of privacy do not necessarily follow the objective level.
-For the best user experience, therefore, we need to quantify and understand users’ perceptions, experiences, and preferences related to privacy, as well as design systems accordingly (see section VI).
-The range of applications and content types where privacy-preserving processing methods are needed is vast, and section VII presents a brief overview thereof.
+For the best user experience, therefore, we need to quantify and understand users’ perceptions, experiences, and preferences related to privacy, as well as design systems accordingly (see section VII).
 Finally, threats to and breaches in privacy have considerable consequences on both societal, economic, and individual levels, which has motivated governments to increasingly regulate the use of technology (see section VIII).
+This paper focuses primarily on speech technology in human-computer interfaces (see fig.
+1), but will also address telecommunication where appropriate.
+Focus is here, however, limited to the acoustic speech signal, also known as the voice since natural language processing is a distinct and largely independent field that warrants treating it separately (e.g.
+[9]).
+Concerning telecommunication, approaches to ethical, legal, and technological questions related to privacy in telecommunication over landlines are well-established, and open discussions are primarily related to how the existing regulation and oversight should be extended to cover also mobile telecommunication and voice-over-Internet protocols (VoIP) [10, 11, 12].
 To the author’s knowledge, this is the widest tutorial overview of privacy in speech technology.
-Recent works in the field, however, include a technical overview [nautsch2019preserving], as well as a popular-science review [backstrom2020privacy] of this area, and several doctoral theses have their respective summaries, e.g.
-[Nelus2022, das2021thesis, zarazaga2022thesis, pathak2012privacy, williams2022learning, portelo2015privacy].
-Privacy has been extensively discussed in other areas of science, such as the neighboring field of natural language processing [mahendran2021privacy], in statistical theory of privacy [cummings2023challenges], in social sciences [petronio2002boundaries] and psychology [Derlega1977PrivacyAS].
-There is even an excellent and thorough meta-study of all areas of science that discuss privacy [van2018handbook].
+Recent works in the field, however, include a technical overview [13], a survey on privacy issues with voice assistants [14], as well as a popular-science review [15] of this area, and several doctoral theses have their respective summaries, e.g.
+[16, 17, 18, 19, 20, 21, 22, 23].
+Privacy has been extensively discussed in other areas of science, such as the neighboring field of natural language processing [9], in statistical theory of privacy [24], in social sciences [25] and psychology [26].
+There is even an excellent and thorough meta-study of all areas of science that discuss privacy [27].
+Central contributions of this work include 1.
+structuring and collecting information, research results, and vocabulary into a unified format, 2.
+proposing a framework for analysis of attack vectors (see fig.
+1), and 3.
+an introduction to the psychological aspects of privacy that is not often discussed in engineering-oriented publications (see section VII).
+II Content categories and applications
+To give a complete picture of the implications of privacy in speech technology, this section provides a brief discussion both on categories of private information, complementing table I as well as the application of such information.
+Observe that this is not a complete list (nor is table I) of the content nor of application categories.
+The purpose is merely to provide a characterization of work done and challenges in the research field, with accompanying references.
+First, note that while all presented categories are (potentially) private information, all sustained information (in the sense that it does not change over a “long” time) can potentially be used to identify a speaker.
+For example, while the current emotional state cannot alone identify a speaker, the tendency to display emotional states can aid in identification.
+However, where we want to verify that a speaker is who they claim to be, we can use only information that cannot be willfully changed.
+That is, speech style is a particular example of a property that a good voice actor can freely choose, making it “easy” to change also for fraudulent purposes.
+Using variable attributes, or their distributions, to identify or verify an identity is thus likely to be complicated.
+Recognizing the speaker’s identity then becomes the natural starting point for studies in privacy.
+We can attempt to recognize who is speaking (speaker recognition), verify whether a speaker has the claimed identity (speaker verification), clustering audio to segments with a single speaker (speaker diarization) and we can develop methods for deceiving identity (spoofing), e.g.
+[28, 37, 44, 44, 45, 46, 47, 48, 49, 29, 30, 31, 32, 46, 47, 33, 34, 35, 36, 38, 39, 40, 41, 42, 43].
+Similarly, by voice conversion we can anonymize a speaker identity by replacing it with a random identity (anonymization) or a specific one (pseudonymization), e.g.
+[50, 51, 52, 53, 54, 55, 44, 56, 57].
+Speaker characterization is the natural complement of speaker idenfication [13, 58].
+Such methods related to speaker identity can be applied for example to recognize the user of a device, verify a customer at a bank, or as a voice avatar in online gaming.
+Similarly, anonymization and pseudonymization can be used to hide the identity from the public media and gaming.
+Speech recognition, as in speech-to-text, is probably the largest sub-area of speech research.
+With respect to privacy, it contains two obvious challenges for privacy.
+We can try to limit side-information, such as eliminating all non-text information from the data stream, e.g.
+[59, 60, 61, 62, 63, 64, 65], or we can use natural language processing to anonymize the text content, e.g.
+[66, 67, 68, 69, 9, 70].
+Privacy is even more important in always-on applications like wake-word detection, i.e.
+when the interface is triggered by a specific keyword like “Computer” in “Computer, lights-off”.
+This application is more sensitive exactly because it is always on as users cannot choose when data is processed [71, 72, 73, 74].
+The always-on characteristic is also prominent in assisted-living applications, which can “monitor people’s daily exercises, consumption of calories and sleeping patterns, and to provide coaching interventions to foster positive behavior” [75, 76].
+Such ambient voice interfaces are often implemented through acoustic sensor networks which pose their own challenges [72, 77, 78, 30, 79, 80, 81].
+Speech enhancement refers to removing background noises and distortions from the desired speech signal (target information) [82].
+This task can have an impact on privacy in two primary ways.
+First, the recording environment can reveal private information about the speaker, and second, background noises, like a competing talker, can contain private information [83, 84].
+Attenuating or removing background noises and competing speakers, as well as removing reverberation, can thus improve privacy.
+In addition, while using a sensor network to capture speech can improve utility, it introduces novel threats as well [80, 18, 85].
+Speech is typically considered to be biometric information in its entirety, including all information categories in table I as such categories can be used to identify a person [86, 87, 13, 88, 89].
+Some, however, limit the term biometric to refer only to physical and behavioral characteristics.
+Such properties include health [90], emotions [91] and gender identity [92, 93, 73, 46], each warranting their own treatment.
 Bob Acoustic pathway User
@@
 Figure 1: Threats to the privacy of a user Alice when speaking with another person Bob or a local device, where information is leaked or shared through the acoustic pathways, the local device, the network or the cloud service, to another person Eve, device or service, contrary to the preferences of Alice (red arrows).
-Each threat is marked with the number of the corresponding part of section III-D.
+Each threat is marked with the number of the corresponding part of section IV-D.
 Dotted red arrows marked with “X” indicate threats outside the current scope and dashed black arrows indicate the potential retransmission of leaked data beyond the primary attack surface.
-II Goals and Definitions
-This paper focuses on the two primary uses of speech technology: telecommunication and human-computer interfaces.
-Concerning telecommunication approaches to ethical, legal, and technological questions related to privacy in telecommunication over landlines are well-established, and open discussions are primarily related to how the existing regulation and oversight should be extended to cover also mobile telecommunication and voice-over-Internet protocols (VoIP) [penders2004privacy, wong2009privacy, arapinis2017analysis].
-The primary focus of this paper is thus human-computer interfaces where a computer processes speech signals (see fig.
-1), but threats to communication applications are also considered where applicable.
-Focus is here, however, limited to the acoustic speech signal, also known as the voice since natural language processing is a distinct and largely independent field that warrants treating it separately (e.g.
-[mahendran2021privacy]).
-Privacy is closely related to challenges in security and it is often difficult to distinguish between them.
-Here, we strike a balance by considering privacy scenarios where an agent has legitimate access to some private speech data of the user but uses it for purposes contradicting with, or gains access beyond, the users’ expectations or preferences.
+III Definitions
+III-A Privacy and security
+Privacy is closely related to challenges in security, and it is often difficult to distinguish between them.
+Here, information security refers to enforcing rightful access to information in scenarios where the access rights are binary (granted or not granted).
+In contrast, information privacy refers to the application of rules to information access and use, in scenarios where access permissions are partial and rules context-dependent (adapted from [94]).
+The rules may distinguish between the ability to access and de facto accessing the data, different ways and purposes of using the data, and sharing the data with third parties.
+The rules may further be explicit (e.g., verbalized or documented) or implicit (as based on, e.g., social norms and tacit expectations) and can include requirements on monitoring, reporting, and control.
 For example, a voice interface can be used to control home automation, but if the service provider shares that voice data with advertisers against the users’ preferences, then it violates privacy (see fig.
@@
 Consequently, security concerns such as identity spoofing, deep fakes, and attacks on devices or networks are excluded here, as these threats are more related to security than privacy.
-A private message is the content of a communication that is required to fulfill the purpose of said communication.
-Side-information is any and all information in a communication other than the private message.
+III-B Information categories
+Target information is the content of a communication that is required to fulfill the purpose of said communication.
+Side-information (also known as auxiliary information, external knowledge, or background knowledge [95]) is any and all information in a communication other than the target information.
 See fig.
 2.
-We use the terms attacker and unintended recipient loosely to describe any entity or agent that gains access to or uses private information in a way that conflicts with the user’s preferences.
-An malicious attacker is intentionally seeking to gain private information, a targeted attacker attempts to gain such access to a specific person, an opportunistic attacker to any person, a motivated attacker is seeking ways to attack, and a curious attacker seizes emerging opportunities.
-Attributes of speech are properties that can be distinguished from a speech signal and describe the speaker(s) or their state, the speech content, style or environment, the recording configuration, or any other related features.
+Attributes (or characteristics [4]) of speech are properties that can be distinguished from a speech signal and describe the speaker(s) or their state, the speech content, style or environment, the recording configuration, or any other related features.
 Table I lists examples of speech attributes with private information.
-Attribute or property inference refers to the identification of such private attributes of speakers of known identity or otherwise.
-Re-identification refers to identifying individuals from apparently anonymized information [ISObiometric, ISOIEC224745_2022].
-Enrollment data is reference data, typically labeled with, for example, speakers’ identities or other attribute(s), to which observations are compared with the purpose of identifying said identity or attribute(s).
-It is used in the context of trusted tasks, where the agent is authorized to access this information.
-In threat tasks, where an attacker has gained access to similar reference data, the corresponding data is known as found data.
-Conversely, speech samples that are compared to a database of samples are here referred to as observations, and in literature, they are sometimes known as trial samples, recognition samples, or public data.
-[tomashenko2022voiceprivacy]
-When a speaker makes a speech command to a computer or service, this implies that the speaker is giving a task to the computer and this is known as the trusted task.
-In contrast, we use threat tasks to model the activity of attackers trying to extract information or perform tasks in a nefarious way.
-The attack surface is the specific point in the system where private information leaves the authorized path to make its way to the attacker.
-Voice conversion refers to any speech processing method that replaces the speaker’s identity or other attributes in a speech signal with another.
+Observe that while the target information is often the lexical information contained in speech, in many applications, the target information can be some other attribute.
+For example, in medical analysis, the target information is the health state of the speaker.
+In that case, the lexical information may be private side-information.
+Sender Channel RecipientRefer to caption
+LegitimateRefer to caption Undesired
+!! Target information Side-information
+Figure 2: The high-level threat model of speech interaction, where a target information is sent through a channel to the legitimate recipient, but consequential side-information is bundled to that message.
+It is a threat to privacy when an undesired recipient gains access to that target information or side information (marked by red arrows and exclamation marks).
+III-C Tasks and processing objectives
+When a speaker makes a speech command to a computer or service, this implies that the speaker is giving a task to the computer, and this is known as the trusted task.
+In contrast, we use threat tasks to model the activity of attackers trying to extract information or perform other nefarious tasks.
+[47, 92].
+Voice conversion refers to any speech processing method that replaces the speaker’s identity or other attributes in a speech signal with those of another.
 Anonymization of a speech signal refers to voice conversion where the purpose is to conceal private information.
@@
 The anonymized speech signal then corresponds to a speech signal of an artificial person, who is known as a pseudospeaker.
-III Threat and attack models
+Attribute or property inference refers to the identification of private attributes of speakers of known identity or otherwise [96].
+Re-identification refers to identifying individuals from apparently anonymized information [88, 97].
+In anonymization tasks, enrollment data is reference data, typically labeled with, for example, speakers’ identities or other attribute(s), to which observations are compared with the purpose of identifying said identity or attribute(s).
+It is used in the context of trusted tasks, where the agent is authorized to access this information.
+In threat tasks, where an attacker has gained access to similar reference data, the corresponding data is known as found data.
+Conversely, speech samples that are compared to a database of samples are here referred to as observations, and in literature, they are sometimes known as trial samples, recognition samples, or public data.
+[98]
+III-D Attack and attacker properties
+We use the terms attacker and unintended recipient loosely to describe any entity or agent that gains access to or uses private information in a way that conflicts with the user’s preferences.
+A malicious attacker is intentionally seeking to gain private information, a targeted attacker attempts to gain such access to a specific person, an opportunistic attacker to any person, a motivated attacker is seeking ways to attack, and a honest-but-curious attacker seizes emerging opportunities [99, 100].
+The attack surface or vector is the specific point in the system where private information leaves the authorized path to make its way to the attacker [101].
+IV Threat and attack models
 To better understand the decisive aspects of threats, it is useful to model threats and their components.
-The most important components of such models include the exposed data sources (section III-A), tactics used by, and data available to the attacker (section III-B), the flow of data on an abstract level, also known as the attack model (section III-C), as well as the specific point in the system where private information leaves the intended path, known as the attack surface (section III-D).
-Sender Channel RecipientRefer to caption
-LegitimateRefer to caption Undesired
-!! Private message Side-information
-Figure 2: The high-level threat model of speech interaction, where a private message is sent through a channel to the legitimate recipient, but consequential side-information is bundled to that message.
-It is a threat to privacy when an undesired recipient gains access to that private message or side information (marked by red arrows and exclamation marks).
-III-A Data sources
+The most important components of such models include the exposed data sources (section IV-A), tactics used by, and data available to the attacker (section IV-B), the flow of data on an abstract level, also known as the attack model (section IV-C), as well as the specific point in the system where private information leaves the intended path, known as the attack surface (section IV-D).
+IV-A Vulnerable data sources
 Speech data can be exposed in two ways (see fig.
-2): First, when a private message is transmitted, it is a threat to privacy when contrary to the preferences of the user, it is used in an unexpected way or transmitted to a third party.
+2): First, when a target information is transmitted, it is a threat to privacy when contrary to the preferences of the user, it is used in an unexpected way or transmitted to a third party.
 Second, since speech contains a wide range of private information (see table I) which is bundled in the acoustic signal in complicated ways, it is challenging to extract only the desired message.
-Typically, speech information thus always contains consequential side information bundled into the private message.
-It is a threat to privacy when such private side-information is transmitted to an undesired recipient alongside the private message.
+Typically, speech always contains consequential side information bundled into the target information.
+It is a threat to privacy when such private side-information is transmitted to an undesired recipient alongside the target information.
 Per definition, we here assume that the legitimate recipient can be trusted with the side information and that any undesirable use of it is labeled as an undesirable service.
 The difference between the two forms of information makes an important difference in approaches to mitigation.
-Transmitting the private message to the recipient is the desired outcome of the interaction and we can thus not complete the desired task without sending the message.
-In some cases, though, we can reduce the amount of information content in the private message, and more generally, we should always secure the transmission for example through cryptography (see sections IV-D, IV-B and V-A2).
+Transmitting the target information to the recipient is the desired outcome of the interaction and we can thus not complete the desired task without sending the message.
+In some cases, though, we can reduce the amount of information content in the target information, and more generally, we should always secure the transmission for example through cryptography (see sections V-A, V-B and V-F).
 Some cryptographic methods, such as homomorphic encryption, even keep the message secret from the receiver but allow the processing of a desired task.
-However, when side-information is exposed, as additional methods we can also use signal processing to better remove, replace, or distort such side-information (see section IV-A).
-III-B Attack types
+However, when side-information is exposed, as additional methods we can also use signal processing to better remove, replace, or distort such side-information (see section V-A).
+IV-B Attack types
 There are three main types of attacks; motivated, opportunistic, and incidental.
@@
 For example, if a database contains speech samples from the members of a particular association, then the re-identification of a target person reveals that person’s affiliation with that association.
-Opportunistic attacks are usually known as birthday attacks, referring to the birthday paradox [menezes2018handbook], where the attacker uses a known weakness on a large pool of potential victims.
-It is like spam email; if the attacker sends a large enough number of such emails, then randomly someone triggers the trap by mistake.
+Opportunistic attacks are usually known as birthday attacks, referring to the birthday paradox [102], where the attacker uses a known weakness on a large pool of potential victims.
+It is like spam email; if the attacker sends a large enough number of such emails, then randomly, someone triggers the trap by mistake.
 In this form of attack, anyone could be the victim.
-Where motivated attacks typically target high-value victims such as famous persons where both protections and rewards are high, birthday attacks typically target ordinary people where the likelihood of success is larger but rewards are smaller.
+For example, a speech service could analyze private speaker attributes (for, e.g., health problems) or affiliations (with, e.g., minority groups), for purposes of targeted advertisement or surveillance.
+Where motivated attacks typically target high-value victims such as famous persons, where both protections and rewards are high, birthday attacks typically target ordinary people, where the likelihood of success is larger but rewards are smaller.
 The third category of “attacks” is incidental, where an agent gains excessive access to information unintentionally or by mistake.
@@
 Such incidents are thus neither intentional nor nefarious, but in the interest of consistency, here they are still called attacks.
-Attacks can further be classified according to the amount of information available for the attacker, such as information about the speakers and about the models used.
+IV-C Attack model
+In the design of privacy-protecting methods, attack models simulate and quantify an attacker’s performance.
+In other words, by putting themselves in the attacker’s place, safeguards designers can better understand pertinent aspects of weaknesses and quantify risks.
+Figure 3 illustrates an attack model for measuring the extent of privacy [56].
+We assume that there is some private data, which is anonymized such that it can be used in a trusted task.
+The attacker has access to the anonymized data, which is here called “public data”.
+Observe that the public data is not necessarily openly available for anyone to see, but it only indicates that the data is sufficiently freely available that the attacker has access to it.
+The attacker has also access to some other data about speakers, found from some other source (found data), which helps in extracting private information.
+Again, the term “found data” is a loose term and denotes any data other than found data that is available to the attacker.
+To complement an Attack(er) model, it can also be useful to define a Protector model for the use-case of the system.
+Jointly, the two can be documented using the Scenario-of-Use scheme, as illustrated in table III [103].
+The accurate definition of the protector model and use-cases is essential, since they define both the target information needed as well as how performance should be evaluated [104].
+TABLE III: The Scenario-of-Use scheme [103].
+Attacker Model (Attack on Privacy) Protector Model (Protection of Privacy)
+Objective The information about the target speakers the attacker seeks to obtain. Objective Defense objective: The information in the spoken audio that must be protected.
+Utility objective: The information in the spoken audio that must be retained.
+Opportunity The attacker’s access to the target, including the availability of the target speakers’ audio.
+Also, access to (and any knowledge about) the protection.
+Opportunity The possibilities available for protection (i.e., counter-measures).
+Also, access to (and any knowledge about) the attack.
+Additional
+resources
+The knowledge, data, and compute available to the attacker to carry out the attack (beyond the access to the target specified under Opportunity). Additional
+resources
+The knowledge, data, and compute available to the protector to defend against attack (beyond the access to the attack specified under Opportunity).
+Attacks can be classified according to the amount of information available for the attacker, such as information about the speakers and about the models used.
 Access to speech samples from the target user (known as enrollment data) are particularly important for the attacker.
@@
 The anonymized sentences are known as trial utterances.
-The attack scenarios were then classified as [tomashenko2022voiceprivacy]:
+The attack scenarios were then classified as [98]:
   1. 1.
@@
 For example, if the task is to anonymize emotional or health status, we can define different attack scenarios depending on the extent to which the attacker has information about categories of emotions and health status used in anonymization.
-III-C Attack model
-In the design of privacy-protecting methods, attack models simulate and quantify an attacker’s performance.
-In other words, safeguards designers can put themselves in the attacker’s place to better understand weaknesses and quantify risks.
-Figure 3 illustrates an attack model for measuring the extent of privacy [maouche20_interspeech].
-We assume that there is some private data, which is anonymized such that it can be used in a trusted task.
-The attacker has access to the anonymized data, which is here called “public data”.
-Observe that the public data is not necessarily openly available for anyone to see, but it only indicates that the data is sufficiently freely available that the attacker has access to it.
-The attacker has also access to some other data about speakers, found from some other source (found data), which helps in extracting private information.
-Again, the term “found data” is a loose term and denotes any data other than found data that is available to the attacker.
 Private data Anonymization Public data Downstream application ! Private
@@
 Anonymized data flow is indicated by dashed lines and the attack by red lines and an exclamation mark.
-III-D Attack Surfaces
+IV-D Attack Surfaces
 We categorize scenarios according to the attack surface from where information is extracted (see fig.
-1): a channel between remote (cloud) services (section III-D1), a user interface to the local (edge) device (section III-D2) or to the cloud service (section III-D3), from the local network (section III-D4), the acoustic pathway (section III-D5), or through the shared user interface of the local device (section III-D7).
+1): a channel between remote (cloud) services (section IV-D1), a user interface to the local (edge) device (section IV-D2) or to the cloud service (section IV-D3), from the local network (section IV-D4), the acoustic pathway (section IV-D5), or through the shared user interface of the local device (section IV-D7).
 We consider only cases where a piece of technology is receiving or transmitting information and exclude human-to-human communication without devices.
@@
 1.
-III-D1 Cloud Leak
-Suppose a user accesses a remote server, known as a cloud service through a local device, known as an edge device (see fig.
+IV-D1 Cloud leak
+Suppose a user accesses a remote server, known as a cloud service, through a local device, known as an edge device (see fig.
 4).
-The cloud service is thus assumed to have legitimate access to the private messages of the user.
-The cloud server, however, can then use that private message or bundled side information for some other purpose than that requested by the user, extract more information than anticipated, combine it with other information, or share information with a third party.
+The cloud service is thus assumed to have legitimate access to the target information of the user.
+The cloud server, however, can then use that target information or bundled side information for some other purpose than that requested by the user, extract more information than anticipated, combine it with other information, or share information with a third party.
 For example, the cloud server of a voice assistant could inappropriately share information with an advertiser.
+Another example is data collection for training of deep neural networks.
+A majority of advanced speech processing today uses machine learning, which has to be trained using large databases of speech.
+The best quality data correspond closely to scenarios where the services are used.
+Recording users’ interactions with their devices is then attractive for training improved models since it corresponds exactly to the use case.
+This presents a considerable threat to privacy because such unrestricted recording could capture a wide range of private information, including all interactions with the device but potentially also any and all speech in the vicinity of the device.
 Observe that though figs.
@@
 For the current analysis, it, however, does not make a difference who is the actor in perpetuating harm or is responsible for the harm caused to the user, but the pertinent difference is the first point where private information departs from the intended flow of information (cf.
-section III-D3).
+section IV-D3).
 User Alice Local device Remote service Secondary service Refer to caption Refer to caption Refer to caption Refer to caption !
@@
 Figure 5: Threat scenario “False activation”, where a user Alice accesses a local device, but the information is shared to a cloud service contrary to preferences (red arrow and exclamation mark).
-III-D2 False Activation
+IV-D2 False activation
 Devices with speech interfaces can hear all conversations in the same acoustic space as the device and, therefore, need mechanisms to determine which utterances are intended for the speech interface (see fig.
 5).
-A popular approach is to use a specific utterance, known as a wake word, to start all interactions with the interface [wang2021wake].
+A popular approach is to use a specific utterance, known as a wake word, to start all interactions with the interface [105].
 The wake word is then like a rudimentary password, which prevents the interface from activating when speech is not directed to the device.
@@
 While false negatives are annoying for the user when the service does not activate, false positives potentially present serious threats to privacy.
-In some famous cases, speech interfaces have activated from sounds on the television to buy unwanted items at the users’ behest, and users’ private conversations have been leaked to third parties [liptak2017alexa, moscaritolo2018alexa].
+In some famous cases, speech interfaces have activated from sounds on the television to buy unwanted items at the users’ behest, and users’ private conversations have been leaked to third parties [106, 107].
 User Local device AliceRefer to caption Refer to caption EveRefer to caption Refer to caption Remote service Cloud storage Refer to caption Refer to caption !!
 Figure 6: Threat scenario “Cloud access”, where user Alice uses a remote service through a local device, where it is potentially stored in the cloud and shared with user Eve, contrary to Alice’s preferences (red arrow and exclamation mark).
-III-D3 Cloud Access
+IV-D3 Cloud access
 Figure 6 illustrates the threat where a user Alice accesses a cloud service through an edge device, and where some private information of Alice is stored.
@@
 An example of this threat is medical records, which can be collected from patients to a centralized database.
-A researcher with authorization to access the database can then potentially extract private information beyond the expected.
-[Teixeira2018]
+A researcher with authorization to access the database can then potentially extract private information beyond the expected [108].
 User Devices in same network Remote service Refer to caption Refer to captionRefer to captionRefer to caption Refer to caption Refer to caption Refer to caption ! Outside the room Room
 Figure 7: Threat scenario “WASN authentication”, where a user uses a service through a wireless acoustic sensor network (WASN) inside a room, but another device, outside the room in the same network, joins the distributed sensor network and shares information contrary to preferences (red arrow and exclamation mark).
-III-D4 WASN Authentication
-The audio quality of speech pickup (in terms of, for example, signal-to-noise ratio or perceptual quality), as well as the usability of voice interfaces, can be improved by using all available connected devices with microphones in the same room or acoustic space [das2021enhancement, zarazaga2020acoustic, sigg2020provable].
+IV-D4 WASN authentication
+The audio quality of speech pickup (in terms of, for example, signal-to-noise ratio or perceptual quality), as well as the usability of voice interfaces, can be improved by using all available connected devices with microphones in the same room or acoustic space [109, 79, 110].
 Such collaboration can be realized with wireless acoustic sensor networks (WASNs), where several independent devices simultaneously pick up speech, and those channels are combined to obtain a high-quality signal (see fig.
-7[Nelus2022, zarazaga2022thesis].
+7[16, 18].
 A sensitive question is, however, authorization: Which devices are allowed to share information with each other?
-One approach is to assume that devices that are in the same room or acoustic space can hear the same signal [zarazaga2022thesis, koppelmann22_interspeech].
+One approach is to assume that devices that are in the same room or acoustic space can hear the same signal [18, 72].
 Presence in the same space is thus already an implicit authorization to participate in a joint signal pickup.
@@
 Figure 8: Threat scenario “Speech interface leak”, where a user, Alice, accesses an edge device (potentially including also a remote service), but another local person, Eve, or device in the same room or acoustic space overhears the speech interaction contrary to preferences (red arrow and exclamation mark).
-III-D5 Speech Interface Leak
+IV-D5 Speech interface leak
 Figure 8 illustrates a scenario where a device or person overhears a discussion between a user, Alice, and a local device.
 The defining property is that the leak is in the acoustic pathway, necessitating that the eavesdropper is physically present in the same acoustic space as the private communication.
-The difference between these leaks and False Activation (sections III-D2 and 5) is that here, the eavesdropper is collecting data inappropriately, whereas in False Activation, an incorrect functionality is activated.
+The difference between these leaks and False Activation (sections IV-D2 and 5) is that here, the eavesdropper is collecting data inappropriately, whereas in False Activation, an incorrect functionality is activated.
 In the case where Eve overhears Alice’s speech, we assume that people have a learned awareness of which other people are present in the same room and adjust their speech accordingly.
@@
 Figure 9: Threat scenario “Discussion leak”, where a dialogue between two users, Alice and Bob, is picked up by a local device and potentially shared to a remote service contrary to preferences (red arrow and exclamation mark).
-III-D6 Discussion Leak
+IV-D6 Discussion leak
 Figure 9 illustrates a discussion between two persons overheard by a local device and, like False activation, triggers unintended functionalities in the local device.
-The main difference to a Speech Interface Leak (sections III-D5 and 8) is that the type of speech is different.
+The main difference to a Speech interface leak (sections IV-D5 and 8) is that the type of speech is different.
 Interaction with a speech interface at the current level of technology is, roughly speaking, limited to speech commands, question-answer pairs, and one-sided monologues, whereas spoken human-to-human interaction has potentially a much more dynamic character.
@@
 Figure 10: Threat scenario “Shared device”, where user Alice uses an edge device, where information is stored and shared with another user, Eve, contrary to Alice’s preferences (red arrow and exclamation mark).
-III-D7 Shared Device
+IV-D7 Shared device
 Many practical scenarios include multiple users sharing one or several speech-operated devices.
@@
 It is also obviously related to many security threats – unauthorized access to devices should be prevented – but those are outside the scope of this work.
-IV Protections
+V Protections
 The main avenues for protecting privacy in speech technology are: 1.
-Information isolation, where private information is filtered or otherwise removed from a transmitted (or stored) channel of speech information (section IV-A), 2.
-Secure processing, where access to private information is prevented with methods of cryptography (section IV-B), 3.
-Privacy-preserving architectures, where private information is not sent to the cloud, but processed either only locally (on the edge), or in a distributed network of (nearby) devices (section IV-A3), 4.
-Acoustic interventions, where reproduction of audio is changed such that private information is not revealed to eavesdroppers (section IV-D), and 5.
-Improving performance, where the intended functionality of a system is improved such that processing errors leak less private information (section IV-E).
-IV-A Information isolation
+Information isolation, where private information is filtered or otherwise removed from a transmitted (or stored) channel of speech information (section V-A), 2.
+Secure processing, where access to private information is prevented with methods of cryptography (section V-B), 3.
+Privacy-preserving architectures, where private information is not sent to the cloud, but processed either only locally (on the edge), or in a distributed network of (nearby) devices (section V-A3), 4.
+Acoustic interventions, where reproduction of audio is changed such that private information is not revealed to eavesdroppers (section V-D), 5.
+Improving utility, where the intended functionality of a system is improved such that processing errors leak less private information (section V-E) and 6.
+Design of provable protectionsby applying theoretical frameworks such as differential privacy (section V-F).
+V-A Information isolation
 By isolation of information, we refer to the design of a channel of information where the target information is isolated from other (undesired) private information.
-Such isolation can be enforced through two main approaches; 1.
-we can reduce the capacity of the channel such that it can represent only the target information and has no room for other private information, which is known as an information bottleneck [tishby2000information] (see section IV-A1), or, 2.
-we can use adversarial training of the model, where the model is trained such that an attacker cannot extract a specific type of private information from the channel (see section IV-A2).
-Such isolation can be implemented for any category of information.
-More generally, a system can include multiple channels of information, each designed to carry a distinct category of information, in an approach known as disentanglement.
-This allows applications to cherry-pick the private information that should be transmitted case by case (see section IV-A3).
-InferenceTraining User EncoderBottleneckDecoder Loss Refer to caption OutputLabelLocal device Remote service !
-Figure 11: Training of a privacy-preserving speech analysis method with the information bottleneck principle. The point of attack is indicated by the red arrow and exclamation mark.
-InferenceTraining User EncoderBottleneckDecoder Loss Refer to caption Output--!$+$
-Figure 12: Training of a privacy-preserving an autoencoder structure as an example of the information bottleneck principle. The point of attack is indicated by the red arrow and exclamation mark.
-IV-A1 Information Bottleneck
-The purpose of an information bottleneck is to use an encoder to distill the information needed to perform a downstream (trusted) task into a bottleneck, such that all other (private) information is removed.222 A concept related to information bottlenecks is the privacy funnel [makhdoumi2014information].
-Where the original definition of the information bottleneck [tishby2000information] aims to maximize the utility of a channel with a fixed rate, the privacy funnel aims to minimize the information rate with a fixed utility.
+This applies at least and in particular to the attack vectors Cloud leak (section IV-D1), False activation (section IV-D2) and Cloud access (section IV-D3), where a reduction in flow of private information would reduce threats to privacy.
+Central concepts in information isolation include 1.
+information bottlenecks(or privacy funnels) , where information is compressed to a low-capacity channel such that only target information is passed through (see section V-A1) and where the neural network implementation usually takes the structure of an autoencoder [111], 2.
+adversarial training, where an information channel (such as an information bottleneck) is trained to contain target information to fulfill the trusted task and simultaneously trained such that a competing, attacking adversarial model fails to extract private information (threat task), (see section V-A2) and 3.
+disentanglement, where different attributes of speech are isolated to separate channels, as a generalization of information bottlenecks (see section V-A3).
+Note that these concepts are overlapping and should be seen as characterizations of algorithms and methodology, rather than alternative approaches.
+InferenceTraining User EncoderBottleneckDecoder Loss Refer to caption OutputLabelLocal device Remote service !!
+Figure 11: Training of a privacy-preserving speech analysis method with the information bottleneck principle. Potential points of attack are indicated by red, dashed arrows and exclamation marks.
+InferenceTraining User EncoderBottleneckDecoder Loss Refer to caption Output-!++!
+Figure 12: Training of a privacy-preserving an autoencoder structure as an example of the information bottleneck principle.
+Potential points of attack are indicated by red, dashed arrows and exclamation marks.
+V-A1 Information bottleneck
+The purpose of an information bottleneck is to use an encoder to distill the information needed to perform a downstream (trusted) task into a bottleneck, such that all other (private) information is removed.
+A decoder can then be used to decipher the information in the bottleneck into a format usable for the downstream task.
+The overall approach is generic and can be applied, for example, in cases where 1.
+the decoder resynthesizes the signal with some attributes anonymized [112] or 2.
+information extraction scenarios such as classification [113], where the decoder can often be trivial.
+A concept related to information bottlenecks is the privacy funnel [114].
+Where the original definition of the information bottleneck [113] aims to maximize the utility of a channel with a fixed rate, the privacy funnel aims to minimize the information rate with a fixed utility.
 The information bottleneck is, however, much more widely known in the community and often informally used as an umbrella term to cover both cases.
-A decoder can then be used to decipher the information in the bottleneck into a format usable for the downstream task.
-The overall approach is generic and can be applied for example in cases where 1.
-the decoder resynthesizes the signal with some attributes anonymized [wang2019vector] or 2.
-information extraction scenarios such as classification [tishby2000information], where the decoder can often be trivial.
-To distill only the private message and discard any side-information, speech information is passed through a bottleneck so tight, that only the legitimate message at full accuracy can pass through [Tishby2015, makhdoumi2014information] (see fig.
+To distill only the target information and discard any side-information, speech information can be passed through a bottleneck so tight, that only the legitimate message at full accuracy can pass through [115, 114] (see fig.
 11).
@@
 The challenges are to design a training methodology and model structure with a sufficiently tight bottleneck so that the legitimate message is accurately retained while private side information is discarded.
-A frequently used approach to implementing a bottleneck is an autoencoder structure, consisting of an encoder, bottleneck, and a decoder, where the objective is to reconstruct the input signal from the bottleneck output [malekzadeh2018replacement, perero2022x, lu2013speech] (see fig.
+Observe that a generic implementation of an information bottleneck does not guarantee privacy as side-information can sneak through the bottleneck alongside the target information [116].
+Instead, to be privacy preserving, a channel must be designed or optimized for privacy, warranting the term privacy funnel.
+A frequently used approach to implementing a bottleneck is an autoencoder structure [111], consisting of an encoder, bottleneck, and a decoder, where the objective is to reconstruct the input signal from the bottleneck output [78, 49, 117] (see fig.
 12).
-However, to be privacy-preserving, the bottleneck should be sufficiently tight that the original speech signal cannot be perfectly reconstructed to resemble the original in their waveform with respect to some distance function.
+To be privacy-preserving, the bottleneck should be sufficiently tight that the original speech signal cannot be perfectly reconstructed to resemble the original in its waveform with respect to some distance function.
 Designing a loss function is then challenging when the bottleneck is so tight that the decoded signal is very far from the original.
-One solution akin to the machine learning method known as representation learning is to feed the output of the autoencoder again to the encoder and compare the bottleneck features [neumann2019improving, chorowski2019unsupervised].
+One solution akin to the machine learning method known as representation learning is to feed the output of the autoencoder again to the encoder and compare the bottleneck features [118, 119].
 Information content at the bottleneck can be reduced by either reducing the signal rank (viz.
-dimensionality reduction, e.g., [perero2022x]) or by quantizing and coding the signal at a low bitrate (e.g., [wang2019vector, williams2021revisiting]).
-The trade-off between the accuracy of trusted task and bottleneck entropy then determines the extent of privacy (see section V).
+dimensionality reduction, e.g., [49]) or by quantizing and coding the signal at a low bitrate (e.g., [112, 120]).
+The trade-off between the accuracy of trusted task and bottleneck entropy then determines the extent of privacy (see section VI).
 Observe that dimensionality reduction is, in general, not alone sufficient because if the channel’s floating- (or fixed)-point representation has $K$ dimensions each with $B$ bits, then a sufficiently advanced encoder can pack $KB$ bits of information to that channel.
@@
 As a general rule, therefore, bottlenecks should always be quantized to the lowest bitrate sufficient to represent the target information.
-A central advantage of the information bottleneck approach is that it gives a provable level of privacy.
+A central advantage of the information bottleneck approach is that it can give a provable level of privacy.
 If the bitrate of the bottleneck is $B_{b}$ and that of the target information is $B_{t}$, then the privacy leak is at most $B_{leak}\leq B_{t}-B_{b}$.
 This applies to single vectors of target information passed through a bottleneck with full accuracy.
-Streaming data through the bottleneck, however, over time, leaks a flow of private information that increases in size without bound [kairouz2015composition].
-This approach also addresses only the average disclosure of private information but does not give guarantees of worst-case disclosure [nautsch20_interspeech, vali2024porcupine].
-Information bottlenecks, or more accurately, privacy funnels, are theoretically well-suited for privacy-preserving applications.
-They have also been used in other areas, such as privacy-aware transmission of images [sun2024privacyaware] and face recognition [razeghi2024deep].
+Streaming data through the bottleneck, however, over time, leaks a flow of private information that increases in size without bound [121, 122].
+This approach also addresses only the average disclosure of private information but does not give guarantees of worst-case disclosure [87, 123].
+Information bottlenecks, or more accurately, privacy funnels, are thus theoretically well-suited for privacy-preserving applications.
+They have also been used in other areas, such as privacy-aware transmission of images [124] and face recognition [125].
 However, the above-mentioned applications in speech either do not specifically apply it for privacy or use information bottlenecks optimized for privacy with additional components.
-InferenceTraining User Encoder Trusted Task Trusted Loss Refer to caption Output Threat Task Threat Loss !$+$--
+InferenceTraining User Encoder Trusted Task Trusted Loss Refer to caption Output Threat Task Threat Loss !++-
 Figure 13: Training of a privacy-preserving speech analysis method with an adversarial approach.
@@
 During inference, the threat task can be ignored.
-IV-A2 Adversarial Approach
-An alternative to reducing the size of the information bottleneck is to use an adversarial model during training, such that side-information in the bottleneck is minimized [noe21_interspeech, srivastava19_interspeech, nelus2018itg].
-Figure 13 illustrates the model structure, where the trusted and threat tasks correspond respectively to utility and privacy, which in turn respectively correspond to the extraction of the private message and side-information.
+V-A2 Adversarial training
+Privacy-preserving neural networks can be trained using adversarial methods, such that side-information in the bottleneck is minimized [48, 64, 92].
+Figure 13 illustrates the model structure, where the trusted and threat tasks correspond respectively to utility and privacy, which in turn respectively correspond to the extraction of the target information and side-information.
 The threat task is independently optimized to extract private side information from the bottleneck output.
 While keeping the threat task fixed, the encoder and the trusted tasks are jointly optimized to maximize the accuracy [^2] Accuracy refers here to some similarity measure between the message and a corrupt version of the message.
-of the private message and to minimize the accuracy of private side information.
+of the target information and to minimize the accuracy of private side information.
 Adversarial training thus combines maximizing and minimizing the accuracy of private side information.
 This forces the encoder to minimize private side information passed through the bottleneck since the threat task tries to maximize the extraction of private side information.
-Similar effects can be achieved by a variety of other formulations, for example, by Siamese training, where the analysis model is specifically optimized to minimize and maximize feature distances between speakers within and across target classes [nelus19b_interspeech].
+Similar effects can be achieved by a variety of other formulations, for example, by Siamese training, where the analysis model is specifically optimized to minimize and maximize feature distances between speakers within and across target classes [47].
 This approach should cluster extracted information into groups corresponding to the trusted task while minimizing private information.
@@
 This can lead to an efficient model in terms of the trade-off between computational cost and output quality.
-The principal issue with adversarial training is that it does not give any formal assurances of privacy (cf.
-section V-A).
-The best it can do is to demonstrate the extent to which the chosen adversarial model was unable to extract private side information from the chosen category of private information.
+The principal issue with adversarial training is that theoretical frameworks for certifying performance (e.g.
+[126]) are not yet widely accepted and there are thus no formal assurances of privacy (cf.
+section VI-A).
+The best it can do is to demonstrate the extent to which the chosen adversarial model was unable to extract private side information from the selected category of private information.
 We can thus expect that some other attacker with a better model could still extract private side-information and that other categories of private information are potentially present in the bottleneck output.
-User EncoderBottleneckDecoderRefer to caption OutputAnonymization!
+User EncoderBottleneckDecoderRefer to caption OutputAnonymization!!
 Figure 14: Privacy-preserving speech processing through disentanglement, where speech is decomposed into independent streams of information, each representing a distinct category of information and where the level of anonymization can be individually chosen for each category.
-The point of attack is indicated by the red arrow and exclamation mark.
-IV-A3 Disentanglement
+Potential points of attack are indicated by red, dashed arrows and exclamation marks.
+V-A3 Disentanglement
 With disentanglement, we refer to methods that split a speech signal into channels corresponding to distinct categories of information.
-For example, selecting from table I we can choose to decompose speech into linguistic information, the speaker’s (voice) identity, and emotions [zhou2021vaw].
+For example, selecting from table I we can choose to decompose speech into linguistic information, the speaker’s (voice) identity, and emotions [127].
 Such disentanglement can be achieved by either or both through adversarial training or by generalizing information bottlenecks to encompass multiple bottleneck channels, where each bottleneck represents a disentangled representation (see fig.
@@
 That is, each channel represents a distinct and statistically independent attribute of the signal such that the extent of anonymization can be cherry-picked per channel as per use-case e.g.
-[aloufi2020privacy, stoidis21_interspeech, zhou2021vaw, williams2022learning].
+[128, 93, 127, 20].
+The privacy-preserving performance of a disentangling system can be readily formulated, provided that channels $k$ are uncorrelated.
+If the bitrates for private information of respective channels are $B_{k}$, then using the composition theorem, the joint disclosure is $\sum_{k}B_{k}$ [121].
+Per-channel disclosures are thus additive such that the design and performance analysis of the privacy-preserving systems is straightforward.
 An important challenge with disentanglement is to define the constraints with which information is funneled to the corresponding channel and the methods by which they are made independent of each other.
 For example, a model can be trained to match specific channels with labels in the dataset (e.g.
-[stoidis21_interspeech, zhou2021vaw]).
+[93, 127]).
 Alternatively, using representation learning, we can constrain channels based on objective criteria such as time-scope, i.e.
 the length of time over which data is integrated (e.g.
-[Niekerk2020, neumann2019improving, lotfidereshgi2022cognitive]).
+[129, 118, 130]).
 Specifically, attributes such as the speaker’s identity are constant over an utterance, while phonemes, intensity and pitch change over a time scale of tens of milliseconds.
@@
 The gray area indicates the domain that is encrypted and red, double arrows marked with exclamation marks indicate the protected stream of data.
-IV-B Secure processing
+V-B Secure processing
 If the transmitted information is encrypted, then it cannot be used for malicious purposes without access to the key.
-Surprisingly, however, it is possible to process encrypted information when using secure multi-party computation (SMPC), such that the encrypted result of the computation can be returned and opened [yao1986generate, portelo2015logsum, portelo2014privacy, Teixeira2018, cramer2015mpc] (see fig.
+Surprisingly, however, it is possible to process encrypted information when using secure multi-party computation (SMPC), such that the encrypted result of the computation can be returned and opened [131, 132, 133, 108, 134] (see fig.
 15).
-The SMPC umbrella contains several categories of methods such as homomorphic encryption, oblivious transfer, and garbled circuits.
+This applies largely to the same attack vectors as Information isolation, including Cloud leak (section IV-D1), False activation (section IV-D2) and Cloud access (section IV-D3), but also WASN authentication (section IV-D4) and Shared device (section IV-D7).
+The SMPC umbrella contains several categories of methods such as homomorphic encryption, oblivious transfer, secret sharing, and garbled circuits.
 The general principle is that in an interaction between two agents, Alice and Bob, Alice can send encrypted data for processing or a request to Bob, such that Bob cannot interpret the input but can perform the requested task.
 Furthermore, when Alice receives Bob’s response, she cannot deduce the operations that Bob has performed, such that the desired task is completed with mutual privacy preserved.
-Another category of approaches is distance-preserving hashing techniques, where data is obfuscated through randomization such that the desired communication is preserved while other private information is kept secret [portelo2015privacy, jimenez2015secure].
-Such secure techniques can be applied, for example, in the extraction of spectral features of speech or speech recognition in the cloud, without plain text access to the speech signal [thaine19_interspeech, Zhang2019].
-It can also be applied in privacy-preserving speaker recognition, authentication or verification [teixeira22_interspeech, pathak11_interspeech, pathak2012ppsa, portelo2014bprivacy], speech search [portelo2015bprivacy], or for secure sound classification [shashanka2006secure].
-Finally, we can compare the similarity of audio signals without revealing them [rane2013privacy, perez2020, zarazaga2022thesis].
-A central problem with such cryptographic protocols is that the operations that can be performed in the encrypted domain are limited, typically to only additions and multiplications [smaragdis2007framework, pathak2012ppsv, pathak2011privacy, pathak2013privacy].
-In contrast, typical speech-processing algorithms also require other operations, such as non-linearities, conditionals, and recursions [chialva2018conditionals], making the algorithms impossible to implement.
-While many operations have standard workarounds [teixeira2019privacy], such as approximation with additions and multiplications, such approximations are often computationally prohibitively complex, see e.g.
-[thaine19_interspeech].
+Another category of approaches is distance-preserving hashing techniques, where data is mapped into a space where distances between points are approximately maintained, allowing some operations like clustering, without revealing the actual data [21, 135].
+Such secure techniques can be applied, for example, in the extraction of spectral features of speech or speech recognition in the cloud, without plain text access to the speech signal [136, 137].
+It can also be applied in privacy-preserving speaker recognition, authentication or verification [31, 39, 41, 43], speech search [138], or for secure sound classification [139].
+Finally, we can compare the similarity of audio signals without revealing them [140, 141, 18].
+A central problem with such cryptographic protocols is that the floating- or fixed-point operations that can be performed in the encrypted domain are limited, typically to only additions and multiplications [142, 143, 144, 145].
+The limitation does not apply to the boolean operations offered by fast fully homomorphic encryption over the torus scheme, garbled circuits, and boolean secret sharing [146].
+In contrast, typical speech-processing algorithms also require other floating- or fixed-point operations, such as non-linearities, conditionals, and recursions [147], making the algorithms impossible to implement.
+While many operations have standard workarounds [148], such as approximation with additions and multiplications, such approximations are often computationally prohibitively complex, see e.g.
+[136].
 In addition, the communication overhead, in terms of the amount of data to send in comparison to the size of the secret message, is very large.
-For a broader overview of cryptographic methods in speech, see [nautsch2019preserving].
+For a broader overview of cryptographic methods in speech, see [13].
 User Local device Remote service Secondary service (a)Refer to caption Refer to caption Refer to caption Refer to caption (b)Refer to caption Refer to caption Reduced information ! Local processing
@@
 Figure 17: Federated learning as an example of distributed learning, where devices train models locally and transmit only model updates to the cloud.
-Red dashed arrows marked with exclamation marks indicate the reduced flow of information.
-IV-C Privacy-preserving architectures
+The red dashed arrows marked with exclamation marks indicate the reduced flow of information.
+V-C Privacy-preserving architectures
 Perhaps the most obvious solution to protecting privacy, for example in the Cloud leak–scenario, is to limit information flow (to some extent) from the local edge device to the cloud server (see fig.
@@
 16(a) (cf.
-section IV-A).
+section V-A).
 At the extreme, the edge device can be entirely disconnected from the cloud or network, and process information only locally, as in fig.
-16(b) [shi2016edge].
+16(b) [149].
 In doing so, we then need to assume that the local device has sufficient capacity and software to complete the requested tasks, that it is not compromised, and that the restriction of information flow is sufficient to ensure that private information is not leaked.
@@
 1, local processing is the best protection as nothing is transmitted to a cloud service.
-A majority of advanced speech processing today, however, uses machine learning, which has to be trained using large databases of speech.
-The best quality data correspond closely to scenarios where the services are used.
-Recording users’ interactions with their devices is then attractive for training improved models since it corresponds exactly to the use case.
-This presents a considerable threat to privacy because such unrestricted recording could capture a wide range of private information, including all interactions with the device but potentially also any and all speech in the vicinity of the device.
-Distributed learning is an approach to training models without the need for centralized data collection [liu2021machine].
+Distributed learning is an approach to training models without the need for centralized data collection [150].
 Models are trained on local edge devices and only model updates are shared between the nodes and/or with a central server (see fig.
@@
 Though the raw data thus never leaves the edge device in distributed learning, this corresponds to a privacy protection approach where information flow is restricted (cf.
-section IV-A).
+section V-A).
 Federated learning is one of the flavors of distributed learning, where a central, cloud server collects, merges, and redistributes model updates.
-It has been used for example in speaker and emotion recognition, language modeling, as well as for unsupervised estimation of microphone clusters in sensor networks [feng22b_interspeech, granqvist20_interspeech, zewoudie2021federated, nelus2021, shoemate2022sottovoce, ro21_interspeech, nguyen2023federated, feng2021attribute].
+It has been used for example in speaker and emotion recognition, language modeling, as well as for unsupervised estimation of microphone clusters in sensor networks [91, 37, 29, 80, 59, 151, 60, 152].
 Overall, distributed learning is a promising approach, even if it has several challenges.
@@
 The learning accumulated during training is then effectively lost every time the model is updated, jeopardizing the fair comparison of competing approaches.
-Third, even if distributed learning does improve privacy, it is not a guarantee for privacy since model updates can also contain private information [Orekondy2018gradientleaks, liu2021machine].
+Third, even if distributed learning does improve privacy, it is not a guarantee for privacy since model updates can also contain private information [153, 150].
 To be specific, except for federated learning, distributed learning protects efficiently against threats 1 to 3 in fig.
 1, but it may potentially leak private information to other devices, and federated learning may leak information to the cloud.
-IV-D Acoustic interventions
+V-D Acoustic interventions
 UserLoudspeaker Local device AliceRefer to caption Refer to caption Refer to caption EveRefer to caption Refer to caption Microphone!
@@
 This private information can be overheard by other users in the same acoustic space, and that presents a threat to privacy (see figs.
-19 and III-D5).
+19 and IV-D5), as in attack vectors False activation (section IV-D2), WASN authentication (section IV-D4), Speech interface leak (section IV-D5), Discussion leak (section IV-D6), and Shared device (section IV-D7).
 In theory, it would be possible to identify and track users in the same acoustic space and communicate private information only when it does not pose a threat to privacy (see fig.
@@
 Such systems have, however, not yet been widely published.
-Figure 19 presents another solution, which uses acoustic interventions to create physical spaces where the private information is, respectively, intelligible or inaudible [gardner1998_3daudio, pulkki2018parametric].
+Figure 19 presents another solution, which uses acoustic interventions to create physical spaces where the private information is, respectively, intelligible or inaudible [154, 155].
 This acoustic effect can be achieved either as a sound zone through constructive and destructive interference between loudspeakers (see e.g.
-[donley2016, nielsen2018, WALLACE2022101285]) or as an audio spotlight through non-linear interference effects between ultrasound sources [yoneyama1983audio].
+[156, 157, 158]) or as an audio spotlight through non-linear interference effects between ultrasound sources [159].
 By choosing the spatial location where speech is intelligible and assuming we know where the target user is located, we can thus limit access to private information only to the target speaker.
-Note that sound zones with destructive interference do receive a partial observation of the private message.
+Note that sound zones with destructive interference do receive a partial observation of the target information.
 Still, it is (hopefully) low-energy and distorted to the extent that it is unintelligible.
@@
 A benefit of this approach is, however, that it requires tracking of only the target listener, which is, while difficult, still much simpler than tracking all the people in the room.
-IV-E Improving Performance
-In many use cases, like False Activation, the actual culprit is the inadequate performance of the service.
+V-E Improving utility
+In many use cases, like False activation (section IV-D2), Speech interface leak (section IV-D4), and Discussion leak (section IV-D6), the actual culprit is the inadequate performance of the service.
 For example, if a wake word detector is incorrectly triggered (false positive), the system will start to listen to a conversation when it was not supposed to, thus breaching privacy.
@@
 The treatment thus addresses the cause rather than the symptoms thereof.
-Preserving privacy thus becomes an issue of systems design and implies using the privacy-by-design approach [Kalloniatis2009, van2011privacy].
+Preserving privacy thus becomes an issue of systems design and implies using the privacy-by-design approach [160, 161].
 This approach has, however, two principal challenges.
-First, improving performance often requires an increase in computational power and other resource consumption.
-This is financially costly and has an environmental penalty [patterson2021carbon].
+First, improving utility often requires an increase in computational power and other resource consumption.
+This is financially costly and has an environmental penalty [162].
 In particular, balancing the computational load is easier on a cloud server, and thus, existing resources can be more efficiently used.
-Second, speech interfaces will always have occasional errors, even with improved performance.
-That is, the design of privacy-preserving speech technology must include multiple layers of protection, especially when requested operations have large consequences [hong2004privacy, backstrom2021edpbcomments].
+Second, speech interfaces will always have occasional errors, even with improved utility.
+That is, the design of privacy-preserving speech technology must include multiple layers of protection, especially when requested operations have large consequences [163, 164].
 Say, when the wake word detector is activated, then before any actions which would potentially breach privacy, the system could require that the speaker is in the same room, or the speaker identity is verified, or an extra confirmation step “Are you sure?” or similar.
@@
 notifies the user of its presence (user-interface design).
-User Alice Local device Noise Remote service Refer to caption Refer to caption $+$ Refer to caption !
+User Alice Local device Noise Remote service Refer to caption Refer to caption ++ Refer to caption !
 Figure 20: Protecting privacy by adding noise to private data, following the idea of differential privacy.
-TABLE III: Dimensions of performance in privacy-preserving speech processing.
+V-F Privacy-by-design and provable protections
+Privacy-by-design is a philosophy or approach to design, where privacy is taken into account from the start of the design process and not as an afterthought [165].
+This increases the chances that the system does not violate privacy unnecessarily, since the privacy-preserving performance is considered at every step of the design process.
+The privacy-preserving performance of a system can be designed to satisfy design requirements provably, using theoretical frameworks, and differential privacy in particular [166, 167, 13].
+Privacy is generally enforced by reducing the accuracy of private information sufficiently, for example, through the addition of noise or by quantization (see fig.
+20 and cf.
+section V-A).
+To determine how much accuracy must be reduced, differential privacy quantifies how the amount of information available about a speaker’s affiliation to a particular group changes due to an observation.
+Specifically, suppose an algorithm $M()$ is applied on a set $D$ to give an answer $M(D)=S$.
+The limit $\epsilon$ for leakage of information (in nits) is then the amount of new information gained through a slightly different dataset $D^{\prime}$ as
+$\log\left|\frac{p\left(M(D)=S\right)}{p\left(M(D^{\prime})=S\right)}\right|\leq\epsilon.$ (1)
+Leaks can also be formulated as a function of the groups’ sizes, where the speaker is known to belong before and after an observation.
+Suppose we know that a speaker $x$ belongs to group $x\in A=\{a_{1},\dotsc,\,a_{M}\}$ with $M$ members.
+Without prior knowledge, the speaker can be any one of the group members with equal likelihood $p\left(x=a_{k}\right)=\frac{1}{M}$.
+The identity within this group, which is unknown to the attacker, then has an information content of $\log M$ nits.
+If an attacker somehow manages to identify that the speaker belongs to a subset $B\subset A$ of size $N$, then the remaining information content unknown to the attacker is $\log N$ nits.
+The reduction in uncertainty is thus the leak magnitude
+$\epsilon=\log M-\log N=\log\frac{M}{N}.$ (2)
+A leak is thus a function of the relative group size before and after an observation.
+It is arguably more informative for users to state the absolute group size within which they remain anonymous.
+If, through the processing of an observation, an attacker can reduce the group size within which the user remains anonymous to $k$, then we say that the process is $k$-anonymous [168].
+Even if individual modules of a system can assure privacy, their interaction in a system can expose privacy.
+Therefore, the privacy-preservation performance of processes involving speech or other biometric signals must be evaluated on a systems level.
+Such evaluation uses a concept of biometric reference, which is a data record that stores an attribute of the user, such as the speakers’ identity.
+New samples are compared to such biometric references to find matching attributes.
+The three key concepts in such evaluation are [97, 13]
+  • Unlinkability – comparing two or more biometric references will not reveal whether they are from the same or different persons.
+  • Irreversibility – information about the biometric sample cannot be derived from the biometric reference.
+  • Renewability – multiple, unlinkable biometric references can be created from the same biometric samples and used to identify the person without revealing information about the biometric samples.
+Together, unlinkability and irreversibility thus minimize the amount of information leaked if biometric references are revealed.
+At the same time, renewability ensures that leaked references can be revoked and new ones created to replace them.
+Such concepts are central to the legal definitions of privacy and first proposals for their mathematical characterization have emerged [169, 170].
+These theoretical design tools and concepts can be used in the design of, say, information bottlenecks (section V-A), to quantify the maximal leak of private information.
+However, care must be taken to ensure that prior assumptions, such as the equal likelihood of all labels, are met (or enforced) in practical scenarios (cf.
+[123]).
+Theoretical methodology for privacy has been applied in, among others, differentially private speech recognition [59], emotion recognition [91], speaker anonymization [53], and federated learning in acoustic sensor network [171].
+As theoretical proofs of privacy have obvious advantages, the adoption of this methodology will likely increase.
+TABLE IV: Dimensions of performance in privacy-preserving speech processing.
 Performance
@@
 quality of experience (QoE)
-V Evaluating the Performance of Privacy-Preserving Methods
+VI Evaluating the performance of privacy-preserving methods
 To evaluate the performance of any privacy-preserving methods, we need performance measures for both utility and privacy, corresponding respectively to the trusted and threat tasks (see fig.
@@
 While the objective performance in both utility and privacy are prerequisites for any good speech application, to the user, it is central that their experience with the service or application is good.
-Correspondingly, performance can always be measured on two dimensions, privacy and utility, as well as objective and subjective (see table III).
-Performance with respect to objective and subjective privacy and utility, are presented in the following subsections (sections V-A, V-B, V-C and V-D), but the subjective experience of privacy is a sufficiently large and independent topic that it deserves its own section (section VI).
-We further need to define how the objective measures are applied in systems design to ensure privacy (see section V-E).
-The evaluation protocol of the VoicePrivacy challenge is a state-of-the-art, practical example of evaluation of both privacy and utility [tomashenko2022voiceprivacy].
-V-A Objective measures of privacy
-Evaluation of objective privacy can be approached through two alternative approaches, through empirical methods or using theoretical concepts.
-The two categories emanate from different traditions and have their own benefits and areas of application, though the methodologies have some overlaps.
-More complete integration of the two categories would be highly beneficial, potentially providing the benefits of both categories, but at the time of writing, methods for such unification are not available.
-More comprehensive treatments of objective privacy metrics are provided in [nautsch2019preserving, noe2022towards].
-V-A1 Empirical Privacy
-Evaluation of privacy in, for example, the VoicePrivacy challenge [tomashenko2022voiceprivacy], is based on direct measurement of identifiability using signal processing and machine learning methods.
+Correspondingly, performance can always be measured on two dimensions, privacy and utility, as well as objective and subjective (see table IV).
+Performance with respect to objective and subjective privacy and utility, are presented in the following subsections (sections VI-A, VI-B, VI-C and VI-D), but the subjective experience of privacy is a sufficiently large and independent topic that it deserves its own section (section VII).
+We further need to define how the objective measures are applied in systems design to ensure privacy (see section V-F).
+The evaluation protocol of the VoicePrivacy challenge is a state-of-the-art, practical example of evaluation of both privacy and utility [98].
+It is based on direct measurement of identifiability using signal processing and machine learning methods.
 The evaluation objective in such empirical privacy is to determine the extent to which an attacker can re-identify an anonymized speaker or anonymized attribute(s) of a speaker (see fig.
 3).
-Measures commonly used to analyze privacy in speaker recognition, -verification, and biometrics include [natgunanathan2016protection]:
+VI-A Objective measures of privacy
+An information-theoretic approach to the objective measurement is to quantify the new information about a speaker gained through an observed speech signal in terms of entropy and units of nits or bits.
+In other words, suppose an attacker has prior assumptions about a speaker, quantified as a probability distribution $p(X)$ of the possible states $X$ of the speaker.
+The information content of that model is the negative logarithm $I(X)=-\log p(X)$ and its entropy is the expectation thereof $H(X)=E\left[I(X)\right]$.
+The attacker then gains access to or makes an observation $Y$.
+This new information modifies the attacker’s beliefs about the speaker, expressed as the conditional model of the speaker $X$ given the information $Y$, or $X|Y$.
+Correspondingly, the posterior, conditional entropy $H(X|Y)$ quantifies the amount of information the attacker has about the speaker after (posterior to) the observation.
+The leak of private information $\epsilon$ is then defined as the change in entropy
+$\epsilon=H\left(X\right)-H\left(X|Y\right).$ (3)
+The leak is thus equal to the mutual information of $X$ and $Y$.
+A central issue with these metrics is that they are based on mathematical models of probability distributions, whereas speech is a natural signal that does not submit to exact formulas [87].
+To use these formulas, we must, therefore, first choose models to describe the distributions of observations and then estimate their parameters.
+Both steps lead to inaccuracies that dilute the evidentiary power of the proofs of privacy.
+Measures commonly used to analyze privacy in speaker recognition, -verification, and biometrics include [172]:
   1. 1.
@@
   3. 3.
-    The equal error rate (EER) is the proportion of misclassified samples when the threshold of acceptance is adjusted to give equal FJR and FAR [maouche20_interspeech, ISObiometric].
+    The equal error rate (EER) is the proportion of misclassified samples when the threshold of acceptance is adjusted to give equal FJR and FAR [56, 88].
 An increase in EER means that the attacker has made more errors, and privacy has improved.
   4. 4.
-    The application-independent log-likelihood-ratio cost function $C^{\min}_{llr}$ generalizes the EER by considering optimal thresholds over all possible prior probabilities and all possible error cost functions [maouche20_interspeech, brummer2006application].
-  5. 5.
-    The expected privacy disclosure from the Privacy ZEBRA framework [nautsch20_interspeech] quantifies the empirical cross-entropy (the leak magnitude) integrated over all possible prior likelihoods.
-  6. 6.
-    The worst-case disclosure also from the Privacy ZEBRA framework [nautsch20_interspeech] is the largest disclosure observed in a dataset.
-Analysis of anonymized speech signals is straightforward since the above measures can be applied with conventional speech analysis methods to identify the anonymized attributes.
-Existing methodologies for evaluating speech analysis algorithms can thus be reused such that the evaluation is based on a body of prior research.
-On the other hand, empirical evaluation of any other form of signal, such as embeddings of machine learning representations or the compressed data of a telecommunications channel, requires analysis methods developed specifically for that data format.
-While the benefit is that such analysis directly evaluates the signal with any potentially confounding or obfuscation due to additional processing, the overhead of purpose-built analysis can be great.
-A central downside of empirical privacy evaluation is that it is based on the current state-of-the-art, which can be surpassed by future technology.
-In other words, weaknesses in anonymization that are not (yet) known to the evaluator can compromise privacy.
-Moreover, it is widely known that the performance of machine learning models can be improved by making the model larger or using more data to train such models.
-A speech analysis model can thus be (likely) improved, not only by future developments but with this currently known methodology that only requires additional effort.
-Empirical evaluation can thus never give proofs, but just characterizations of the level of privacy.
-As mentioned above, empirical evaluation of privacy using EER has been applied, for example, in the VoicePrivacy challenge [tomashenko2022voiceprivacy].
-Both EER and $C^{\min}_{llr}$ are widely used in any speaker recognition [KINNUNEN201012].
-The worst-case disclosure in a quantized bottleneck in speaker recognition has been studied in [vali2024porcupine].
-V-A2 Theoretical Privacy
-Theoretical evaluation of privacy aims to give assurances or proofs of the level of privacy based on notions of information theory.
-The idea is to quantify the new information about a speaker gained through an observed speech signal in terms of entropy and units of nits or bits.
-In other words, suppose an attacker has prior assumptions about a speaker, quantified as a probability distribution $p(X)$ of the possible states $X$ of the speaker.
-The information content of that model is the negative logarithm $I(X)=-\log p(X)$ and its entropy the expectation thereof $H(X)=E\left[I(X)\right]$.
-The attacker then gains access to or makes an observation $Y$.
-This new information modifies the attacker’s beliefs about the speaker, expressed as the conditional model of the speaker $X$ given the information $Y$, or $X|Y$.
-Correspondingly, the posterior, conditional entropy $H(X|Y)$ quantifies the amount of information the attacker has about the speaker after (posterior) to the observation.
-The leak of private information $\epsilon$ is then defined as the change in entropy
-$\epsilon=H\left(X\right)-H\left(X|Y\right).$ (1)
-The leak is thus equal to the mutual information of $X$ and $Y$.
-A related concept is the log-likelihood ratio (LLR), defined as the logarithm of the ratio of likelihoods $p(Y|H_{k})$ of the observation $Y$ conditioned with the hypotheses that either the sample is $H_{1}$, or is not $H_{0}$, the target speaker [nist2021SRE]
+    The log-likelihood ratio (LLR), defined as the logarithm of the ratio of likelihoods $p(Y|H_{k})$ of the observation $Y$ conditioned with the hypotheses that either the sample is $H_{1}$, or is not $H_{0}$, the target speaker [173]
     $LLR(s)=\log\left(\frac{p\left(Y|H_{0}\right)}{p\left(Y|H_{1}\right)}\right)=I(Y|H_{0})-I(Y|H_{1}).$ (4)
@@
 Comparing to eq.
-1, we can conclude that where eq.
+3, we can conclude that where eq.
 3 is the expectation of information content, eq.
 4 quantifies information content explicitly.
-Another important and related formulation is differential privacy, which quantifies how the amount of information about a speaker’s affiliation to a particular group changes due to an observation [dwork2006differential, dwork2008differential, nautsch2019preserving].
-In particular, suppose an algorithm $M()$ is applied on a set $D$ to give an answer $M(D)=S$.
-The leakage of information is then new information gained through a slightly different dataset $D^{\prime}$ as
-$\log\left|\frac{p\left(M(D)=S\right)}{p\left(M(D^{\prime})=S\right)}\right|\leq\epsilon.$ (3)
-Leaks can also be formulated as a function of the groups’ sizes where the speaker is known to belong before and after an observation.
-Suppose we know that a speaker $x$ belongs to group $x\in A=\{a_{1},\dotsc,\,a_{M}\}$ with $M$ members.
-The speaker can be any one of the group members with equal likelihood $p\left(x=a_{k}\right)=\frac{1}{M}$.
-The identity within this group, which is unknown to the attacker, then has an information content of $\log M$ nits.
-If an attacker somehow manages to identify that the speaker belongs to a subset $B\subset A$ of size $N$, then the remaining information content unknown to the attacker is $\log N$ nits.
-The reduction in uncertainty is thus the leak magnitude
-$\epsilon=\log M-\log N=\log\frac{M}{N}.$ (4)
-A leak is thus a function of relative group size before and after an observation.
-It is arguably more informative for users to state the absolute group size within which they remain anonymous.
-If, through the processing of an observation, an attacker can reduce the group size within which the user remains anonymous to $k$, then we say that the process is $k$-anonymous [sweeney2002k].
-A central issue with theoretical metrics is that they are based on mathematical definitions, whereas speech is a natural signal that does not submit to exact formulas [nautsch20_interspeech].
-To use these formulas, we must, therefore, first choose models to describe the distributions of observations and then estimate their parameters.
-Both steps lead to inaccuracies that dilute the evidentiary power of the proofs of privacy.
-Theoretical methodology for privacy has been applied in, among others, differentially private speech recognition [shoemate2022sottovoce], emotion recognition [feng22b_interspeech], speaker anonymization [shamsabadi2023diff], and federated learning in acoustic sensor network [el2022differential].
-As theoretical proofs of privacy have obvious advantages, the adoption of this methodology will likely increase.
-V-B Subjective evaluation of privacy
-In speaker recognition, computers are now on par or better than humans  [hansen2015speaker] and are not fooled by voice mimicry (i.e., voice actors) even when humans struggle to do so [vestman2020methods, vestman2020voice].
+  5. 5.
+    The application-independent log-likelihood-ratio cost function $C^{\min}_{llr}$ generalizes the EER by considering optimal thresholds over all possible prior probabilities and all possible error cost functions [56, 174].
+  6. 6.
+    The expected privacy disclosure from the Privacy ZEBRA framework [87] quantifies the empirical cross-entropy (the leak magnitude) integrated over all possible prior likelihoods.
+  7. 7.
+    The worst-case disclosure also from the Privacy ZEBRA framework [87] is the largest disclosure observed in a dataset.
+Private information can be potentially extracted (and privacy-preserving performance evaluated) from either 1.
+the waveform of a speech signal or 2.
+from any intermediate representation of speech or its attributes within a speech processing system, such as an information bottleneck.
+In the analysis of waveforms, we can use generic speech analysis tools to determine the magnitude of disclosure.
+In contrast, intermediate representations depend on the configuration and design of the speech processing system.
+Analysis and evaluation tools thus have to be purpose-built and/or optimized for every different speech processing system, which is usually a time and labour-intensive effort.
+The benefit is, however, that when analyzing the intermediate representations, such as an information bottleneck, we analyze a compact representation of the potential leak, without the potentially confounding effects introduced by later processing steps, such as the decoder of an autoencoder.
+Analyzing waveforms can thus be easier, while analysis of intermediate representations is more effort but potentially more informative.
+A central downside of empirical privacy evaluation, such as the analysis of the accuracy of a speaker recognition model, is that it is based on a current state-of-the-art model, which can be surpassed by future technology.
+In other words, weaknesses in anonymization that are not (yet) known to the evaluator can compromise privacy.
+Moreover, it is widely known that the performance of machine learning models can be improved by making the model larger or using more data to train such models.
+A speech analysis model can thus be (likely) improved, not only by future developments but also with this currently known methodology that only requires additional effort.
+Empirical evaluation can thus never give proofs, but just characterizations of the level of privacy.
+In contrast, provable privacy can be provided by theoretical frameworks such as differential privacy (cf.
+section V-F).
+As mentioned above, empirical evaluation of privacy using EER has been applied, for example, in the VoicePrivacy challenge [98].
+Both EER and $C^{\min}_{llr}$ are widely used in any speaker recognition [35].
+The worst-case disclosure in a quantized bottleneck in speaker recognition has been studied in [123].
+More comprehensive treatments of objective privacy metrics are provided in [13, 175].
+VI-B Subjective evaluation of privacy
+In speaker recognition, computers are now on par or better than humans [36] and are not fooled by voice mimicry (i.e., voice actors) even when humans struggle to do so [176, 177].
 We can thus expect that computers are better at evaluating privacy in well-defined tasks.
@@
 First, in identifying private information, human listeners can potentially use properties or features of speech signals other than those computers use.
-Subjective listeners can thus potentially discover new weaknesses that computers miss.
+Subjective listeners can thus potentially discover new weaknesses that computers miss (though the opposite is also true).
 Second, in scenarios where humans are the potential attackers extracting information (as in threats 3, 5, and 7 in fig.
@@
 Similarly, if the threat is speakers’ attribute inference, the subjects should evaluate the extent to which they can identify such attributes.
-For a practical implementation of such subjective evaluation of privacy, see [tomashenko2022voiceprivacy].
-V-C Objective performance in downstream tasks
+As an example, in the VoicePrivacy 2020 challenge, subjective speaker verifiability was measured by asking subjects to rate, on a scale of 1 to 10110, the similarity between an anonymized target and a reference speaker, where the reference was either the same or a different speaker [98].
+If the distributions of similarity scores have significant differences, then the anonymization has failed.
+Alternatives to the 1 to 10110 Likert-scale include, for example, AB (are these samples the same person?) and ABX (which of AB is equal to X?) tests.
+For a comprehensive exposition of subjective evaluation methods, see [178].
+VI-C Objective performance in downstream tasks
 The utility of a privacy-performing system can be quantified by measuring the performance of the downstream, trusted task.
@@
 In applications where a computer analyzes the speech signal, we quantify utility by measuring the accuracy of analysis.
-For example, in automatic speech recognition, where the computer extracts the text content of a speech signal, a commonly used measure is word error rate (WER), which is the proportion of correctly identified words [rohit2024asr].
+For example, in automatic speech recognition, where the computer extracts the text content of a speech signal, a commonly used measure is word error rate (WER), which is the proportion of correctly identified words [179].
 By measuring the word error rate with and without privacy-preserving processing, we can determine how much privacy-protections impede utility.
-V-D Subjective performance in downstream tasks
-If a human user is to listen to the output signal, then that user is the ultimate measure of performance [zacharov2018sensory].
-For example, if we implement speech enhancement to reduce the perceived noise in a signal, then a user can subjectively evaluate how close the enhanced signal is to the noiseless original [vincent2018audio].
-Such perceptual evaluation is best implemented using subjective listening test standards such as P.800 or MUSHRA [P800, MUSHRA].
+VI-D Subjective performance in downstream tasks
+If a human user is to listen to the output signal, then that user is the ultimate measure of performance [178].
+For example, if we implement speech enhancement to reduce the perceived noise in a signal, then a user can subjectively evaluate how close the enhanced signal is to the noiseless original [180].
+Such perceptual evaluation is best implemented using subjective listening test standards such as P.800 or MUSHRA [181, 182].
 Similarly, if privacy protections degrade a speech signal, we can readily quantify such degradations by listening tests.
@@
 They estimate the outcome of subjective listening tests by computational means.
-For example, a classic objective measure known as PESQ estimates the outcome of a P.800 test [rix2001perceptual].
-Later improvements include POLQA and ViSQOL [beerends2013perceptual, chinen2020visqol].
+For example, a classic objective measure known as PESQ estimates the outcome of a P.800 test [183].
+Later improvements include POLQA and ViSQOL [184, 185].
 Such objective measures can be used to automate quality testing at a lower price and faster speed so that repeated estimates with the same input give the same result.
 Observe that researchers should carefully review the specifications of such algorithms to determine whether they apply to the problem at hand to ensure reliable results.
-V-E Privacy-preserving systems
-Even if individual modules of a system can assure privacy, their interaction in a system can expose privacy.
-Therefore, the privacy-preservation performance of processes involving speech or other biometric signals must be evaluated on a systems level.
-Such evaluation uses a concept of biometric reference, which is a data record that stores an attribute of the user, such as the speakers’ identity.
-New samples are compared to such biometric references to find matching attributes.
-The three key concepts in such evaluation are [ISOIEC224745_2022, nautsch2019preserving]
-  • Unlinkability – comparing two or more biometric references will not reveal whether they are from the same or different persons.
-  • Irreversibility – information about the biometric sample cannot be derived from the biometric reference.
-  • Renewability – multiple, unlikable biometric references can be created from the same biometric samples and used to identify the person without revealing information about the biometric samples.
-Together, unlinkability and irreversibility thus minimize the amount of information leaked if biometric references are revealed.
-At the same time, renewability ensures that leaked references can be revoked and new ones created to replace them.
-VI Perception and Psychology of Privacy
+VII Perception and psychology of privacy
+Much of the work of the speech community is focused on technical aspects of privacy-preserving algorithms.
+Yet speech devices are used by and designed for people, and they use people’s voice data.
+It is therefore essential to know what people think of speech devices, how they use and want to use them, and their preferences about privacy.
+Much of this work falls under the fields of psychology, user-interface design, and social sciences.
+This section is an introduction to these topics.
+VII-A Psychological privacy and ownership
 People tend to have a deep sense of ownership of some things, both material and immaterial.
-In particular, people have a feeling of ownership toward information about themselves [cichy2014extending, dawkins2017psychological].
+In particular, people have a feeling of ownership toward information about themselves [186, 187].
 Such feelings are, to some extent, detached from the material consequences of breaches in privacy.
@@
 To qualitatively or quantitatively measure such effects, we also need subjective user tests.
-Concurrently, people have well-established social rules regarding human-to-human privacy [petronio2002boundaries].
-Moreover, users also tend to anthropomorphize technology, that is, treat devices and services as if they were human [nass1993anthropomorphism, cornelius2021acceptance], such that they will likely assume that devices apply human-like social rules to privacy.
+Concurrently, people have well-established social rules regarding human-to-human privacy [25].
+Moreover, users also tend to anthropomorphize technology, that is, treat devices and services as if they were human [188, 189], such that they will likely assume that devices apply human-like social rules to privacy.
 Observe that such social rules are related to human-like behavior and performance.
 That does not reveal whether they are applicable to the super-human performance that computers can possess, such as the ability to integrate information over massive databases and to retain accurate records of events long past.
-How machines should then behave, especially with respect to super-human capabilities, is then not only a question of user-interface design but of moral psychology [thompson2021machine]: We need discussions on a societal level of how automated services should behave and how they are allowed to behave (see also section VIII).
-VI-A Perception and Experience of Privacy
-Irrespective of whether a system actually preserves privacy or not, some users can perceive the system as threatening, and others can be oblivious to the threats it poses [pepino2020distrust].
+How machines should then behave, especially with respect to super-human capabilities, is then not only a question of user-interface design but of moral psychology [190]: We need discussions on a societal level of how automated services should behave and how they are allowed to behave (see also section VIII).
+VII-B Perception and experience of privacy
+Irrespective of whether a system actually preserves privacy or not, some users can perceive the system as threatening, and others can be oblivious to the threats it poses [191].
 Clearly, users do not like using services they perceive as threatening.
@@
 Perception and experience of privacy can be approached in two alternative ways.
-We can make user studies where people interact with either machines or each other [leschanowsky2024review].
+We can make user studies where people interact with either machines or each other [192].
 The distinction is important in the sense that human-to-human interaction relies on social rules that are well-established and developed over a long time.
@@
 Moreover, by using studies of human-to-human interaction, we can learn effects related to human-like performance but can probably not rely on observing effects related to super-human performance.
-Studies of human-to-human interaction are, however, always a proxy if the actual target is to design human-to-computer interfaces [leschanowsky2024review].
+Studies of human-to-human interaction are, however, always a proxy if the actual target is to design human-to-computer interfaces [192].
 Studies of human-computer interaction thus characterize the desired phenomenon directly.
 The compromise is, however, that people’s understanding of privacy with devices might not reflect the true level of privacy, and conclusions made based on the users’ opinions might then not reflect their true preferences and those preferences are moving targets.
-As people gain more experience with and as they learn more about technology, their attitudes, and perceptions of it change [leschanowsky2024review].
+As people gain more experience with and as they learn more about technology, their attitudes, and perceptions of it change [192].
 Despite these shortcomings, both types of experiments are essential for improving our understanding of privacy and for improving technology.
-For example, human-to-human studies have revealed that people experience privacy differently in different acoustic environments; a noisy cafeteria can be better at masking sounds and defending against eavesdropping than a reverberant hallway [zarazaga19_interspeech, das2021thesis, zarazaga2022thesis, backstrom2021intuitive, Leschanowsky2020] and multiple-bed patient rooms in hospitals have privacy-concerns [reinten2017speech].
-Similarly, human-machine studies have revealed that when a chatbot actively communicates choices related to privacy it improves users’ experience of privacy [bruggemeier2022perceptions], voice interfaces with unknown features cause fear in users, and reduces retention of services [yeasmin2020privacy, Lau2018, lutz2021privacy] and breaches in privacy are highly detrimental to the trust in services and reduce users’ willingness to use the services, but such trust can be rebuilt [Kurz2021].
-VI-B User Interface Design
-For usable privacy, it is important that information about privacy is readily provided, changes in the extent of privacy are promptly notified, and that users have control of the level of privacy [bruggemeier2022perceptions, so2021voice, lewis1998designing, Lau2018, Kurz2021, dhiya2023privacy, iachello2007end, sivaraman2018smart].
+For example, human-to-human studies have revealed that people experience privacy differently in different acoustic environments; a noisy cafeteria can be better at masking sounds and defending against eavesdropping than a reverberant hallway [193, 17, 18, 194, 195] and multiple-bed patient rooms in hospitals have privacy-concerns [196].
+Similarly, human-machine studies have revealed that when a chatbot actively communicates choices related to privacy it improves users’ experience of privacy [197], voice interfaces with unknown features cause fear in users, and reduces retention of services [198, 199, 200] and breaches in privacy are highly detrimental to the trust in services and reduce users’ willingness to use the services, but such trust can be rebuilt [201].
+VII-C User interface design
+For usable privacy, it is important that information about privacy is readily provided, changes in the extent of privacy are promptly notified, and that users have control of the level of privacy [197, 202, 203, 199, 201, 204, 205, 206].
 As a parallel to speech interfaces, visual interfaces can use lights or icons for monitoring and tactile interfaces for controlling the level of privacy.
-While sound can be used to monitor system status with sonification [hildebrandt2016continuous, iber2021auditory], this is not in widespread use within signaling of privacy.
-It is a compromise between filling the acoustic space with information and the tendency of the human auditory system to block out monotonous sounds [pulkki2015communication].
+While sound can be used to monitor system status with sonification [207, 208], this is not in widespread use within signaling of privacy.
+It is a compromise between filling the acoustic space with information and the tendency of the human auditory system to block out monotonous sounds [209].
 To be effective, the sound should be perceivable when the user consciously wants to check the privacy level and changes in privacy level should evoke correspondingly and appropriately large changes in the soundscape that users consciously register such changes.
-Furthermore, user interfaces should give correct information about and allow controlling the privacy level [Lau2018, backstrom2020privacy].
+Furthermore, user interfaces should give correct information about and allow controlling the privacy level [199, 15].
 In any case, it is imperative that services are designed such that they reflect the true extent of privacy.
@@
 In fact, service providers may have short-term incentives to follow such approaches to design as long as they improve overall user satisfaction.
-However, such approaches to design are known as dark patterns or deceptive design patterns, and they are considered unethical [willis2020deception].
+However, such approaches to design are known as dark patterns or deceptive design patterns, and they are considered unethical [210].
 Through deception, such dark patterns lull users into believing they are safe when, in fact, they are abused.
 Therefore, good design of privacy should actively communicate and enable control of the true extent of privacy.
-Such design practice is not only ethical but also rewards service providers in the long term by improving retention of those services [Kurz2021].
-VI-C Privacy and Trust
-The level of privacy provided by a speech service is intimately linked to the level of trust users experience in that service [leschanowsky2024review].
+Such design practice is not only ethical but also rewards service providers in the long term by improving retention of those services [201].
+VII-D Privacy and trust
+The level of privacy provided by a speech service is intimately linked to the level of trust users experience in that service [192].
 In particular, if users experience that a service is detrimental to their privacy or, more generally, to their well-being, they will obviously not trust it.
-We thus discuss trust mainly from a psychological perspective, unlike in security research and cryptography, where trust has a specific and unambiguous mathematical definition [balsa2022cryptography].
-While the meaning of trust may seem intuitively clear, it is however difficult to define exactly. A classic definition is [deutch1973resolution]
+We thus discuss trust mainly from a psychological perspective, unlike in security research and cryptography, where trust has a specific and unambiguous mathematical definition [211].
+While the meaning of trust may seem intuitively clear, it is however difficult to define exactly. A classic definition is [212]
 Trust: confidence that [one] finds what is desired rather than feared.
-This definition highlights the psychological nature of trust [simpson2007psychological].
+This definition highlights the psychological nature of trust [213].
 From the current engineering perspective, this translates to the statistical confidence that the outcome of a process is useful.
@@
 Implicitly, this furthermore requires that we predict the outcome of that process in the form of a statistical distribution of possible outcomes and that the expectation over the possible outcomes is that the process is useful.
-Another way of expressing this is to say that trust requires dimensions of competence, integrity, and benevolence [Chen2003, Xie2009].
+Another way of expressing this is to say that trust requires dimensions of competence, integrity, and benevolence [214, 215].
 Observe the semantic difference between a system that is “benevolent” and one that is “useful”.
@@
 It is useful and ethical to design speech services that are trustworthy.
-This means that systems should be such that it would be logical for users to trust them [hardin1996trustworthiness].
+This means that systems should be such that it would be logical for users to trust them [216].
 Speech technology should thus be designed to have useful outcomes with statistically significant confidence, especially with respect to privacy.
-A practical approach for reducing the likelihood of negative outcomes is to build multiple layers of checks and verifications, similar to multi-factor authentication [ometov2018multi].
+A practical approach for reducing the likelihood of negative outcomes is to build multiple layers of checks and verifications, similar to multi-factor authentication [217].
 In fig.
-4, this can be implemented by applying layers of privacy protection in each successive connection and device as well as in their internal modules and components, as described in section IV-E.
+4, this can be implemented by applying layers of privacy protection in each successive connection and device as well as in their internal modules and components, as described in section V-E.
 Activation of an undesired functionality of the system or leaking private information would thus require that every subsequent module fails simultaneously, which is unlikely in most scenarios.
@@
 Thus, it seems obvious that users will tend to trust more those devices they have in their possession or homes, rather than remote cloud services.
-VII Content Categories and Applications
-To give a complete picture of the implications of privacy in speech technology, this section provides a brief discussion both on categories of private information, complementing table I as well as the application of such information.
-Observe that this is not a complete list (nor is table I) of the content nor of application categories.
-The purpose is merely to provide a characterization of work done and challenges in the research field, with accompanying references.
-First, note that while all presented categories are (potentially) private information, all sustained information (in the sense that it does not change over a “long” time) can potentially be used to identify a speaker.
-For example, while the current emotional state cannot alone identify a speaker, the tendency to display emotional states can aid in identification.
-However, where we want to verify that a speaker is who they claim to be, we can use only information that cannot be willfully changed.
-That is, speech style is a particular example of a property that a good voice actor can freely choose, making it “easy” to change also for fraudulent purposes.
-Using variable attributes, or their distributions, to identify or verify an identity is thus likely to be complicated.
-Recognizing the speaker’s identity then becomes the natural starting point for studies in privacy.
-We can attempt to recognize who is speaking (speaker recognition), verify whether a speaker has the claimed identity (speaker verification), clustering audio to segments with a single speaker (speaker diarization) and we can develop methods for deceiving identity (spoofing), e.g.
-[chouchane21_interspeech, granqvist20_interspeech, mawalim20_interspeech, mawalim20_interspeech, nautsch19b_interspeech, nelus19_interspeech, nelus19b_interspeech, noe21_interspeech, perero2022x, zewoudie2021federated, zewoudie2021use, teixeira22_interspeech, treiber2019privacy, nelus19_interspeech, nelus19b_interspeech, BAI202165, han2020voice, KINNUNEN201012, hansen2015speaker, parthasarathi2013wordless, pathak11_interspeech, parthasarathi11_interspeech, pathak2012ppsa, jimenez2017two, portelo2014bprivacy].
-Similarly, by voice conversion we can anonymize a speaker identity by replacing it with a random identity (anonymization) or a specific one (pseudonymization), e.g.
-[noe20_interspeech, pierre22_interspeech, prajapati21_interspeech, shamsabadi2023diff, srivastava20_interspeech, Srivastava2022taslp, mawalim20_interspeech, maouche20_interspeech, Fang2019].
-Speaker characterization is the natural complement of speaker idenfication [nautsch2019preserving, nguyen2010automatic].
-Such methods related to speaker identity can be applied for example to recognize the user of a device, verify a customer at a bank, or as a voice avatar in online gaming.
-Similarly, anonymization and pseudonymization can be used to hide the identity from the public media and gaming.
-Speech recognition, as in speech-to-text, is probably the largest sub-area of speech research.
-With respect to privacy, it contains two obvious challenges for privacy.
-We can try to limit side-information, such as eliminating all non-text information from the data stream, e.g.
-[shoemate2022sottovoce, nguyen2023federated, aloufi21_interspeech, garg20b_interspeech, shah21_interspeech, srivastava19_interspeech, tomashenko2022privacy], or we can use natural language processing to anonymize the text content, e.g.
-[liu22_interspeech, maouche2022enhancing, adelani20_interspeech, novotney21_interspeech, mahendran2021privacy, habernal2021differential].
-Privacy is even more important in always-on applications like wake-word detection, i.e.
-when the interface is triggered by a specific keyword like “Computer” in “Computer, lights-off”.
-This application is more sensitive exactly because it is always on as users cannot choose when data is processed [koppelmann21_interspeech, koppelmann22_interspeech, stoidis22_interspeech, yang21_interspeech].
-The always-on characteristic is also prominent in assisted-living applications, which can ”monitor people’s daily exercises, consumption of calories and sleeping patterns, and to provide coaching interventions to foster positive behavior” [haider19_interspeech].
-Such ambient voice interfaces are often implemented through acoustic sensor networks which pose their own challenges [koppelmann22_interspeech, oconnor20_interspeech, malekzadeh2018replacement, zewoudie2021use, zarazaga2020acoustic, nelus2021].
-Speech enhancement refers to removing background noises and distortions from the desired speech signal (private message) [benesty2008springer].
-This task can have an impact on privacy in two ways.
-First, the recording environment can reveal private information about the speaker, and second, background noises, like a competing talker, can contain private information [zarazaga21_interspeech, rech2022multi].
-Attenuating background noises and competing speakers, as well as removing reverberation, can thus improve privacy.
-Again, while using a sensor network to capture speech can improve utility, it introduces novel threats as well [nelus2021, zarazaga2022thesis, hendriks2013ppdistributed].
-Speech is typically considered to be biometric information in its entirety, including all information categories in table I as such categories can be used to identify a person [adler2006towards, nautsch20_interspeech, nautsch2019preserving, ISObiometric, edpb2021guidelines].
-Some, however, limit the term biometric to refer only to physical and behavioral characteristics.
-Such properties include health [natarajan2021prioris], emotions [feng22b_interspeech] and gender identity [nelus2018itg, stoidis21_interspeech, stoidis22_interspeech, nelus19_interspeech], each warranting their own treatment.
 VIII Legal and societal landscape
@@
 While we have not (yet) seen breaches related to speech technology with consequences of comparable magnitude, these parallels highlight the potential effect of breaches.
-Scandals directly related to speech technology include eavesdropping on private persons by employees and contractors of service providers [Lynskey2019GuardianPrivacyScandal, hern2019, day2019].
-Section VI-B also makes the point that service and technology providers have clear short-term incentives that conflict with users’ preferences for privacy.
+Scandals directly related to speech technology include eavesdropping on private persons by employees and contractors of service providers [5, 6, 221].
+Section VII-C also makes the point that service and technology providers have clear short-term incentives that conflict with users’ preferences for privacy.
 We can easily find examples where users do not have real choices in protecting their privacy.
-For example, suppose all friends and family of a user use a particular platform for social and voice interaction even if its privacy configuration is inadequate.
+For example, suppose all friends and family of a user use a particular platform for social and voice interaction even if its privacy configuration is inadequate [222].
 The individual user faces then the choice of either disconnecting from their social network or compromising privacy.
@@
 On the individual level, the examples of potential exploits in table II cause clear damage to individuals, including psychological harm in particular (e.g.
-[natarajan2021prioris]).
-Additional dangers include stalkers [lu2019abusers, hulsen2022stalkers, avast2021stalkerware].
-While such individual damage is “small” on the societal level, their prevalence makes their joint impact significant [levy2020intimate].
+[90]).
+Additional dangers include stalkers [223, 224, 225] and oppression [222].
+While such individual damage is “small” on the societal level, their prevalence makes their joint impact significant [8].
 These examples demonstrate the inherent need for society-level regulation of speech technology with respect to privacy.
-Governments have already responded to this need, with the European Union spearheading the process with the General Data Protection Regulation (GDPR) [eu2016gdpr, nautsch19c_interspeech], with the State of California following soon thereafter with the California Consumer Privacy Act (CCPA) of 2018 [california2018ccpa].
+Governments have already responded to this need, with the European Union spearheading the process with the General Data Protection Regulation (GDPR) [226, 227], with the State of California following soon thereafter with the California Consumer Privacy Act (CCPA) of 2018 [228].
 While these laws cover only a small percentage of the global population, as cloud services typically operate globally, they need the capability to follow local laws.
@@
 Service providers have, therefore, widely adopted the requirements of GDPR and CCPA, and that has likely had a large impact also on users outside the scope of these regulations.
-With respect to regulation, an important consequence of the objective measures of privacy in section V is that our tools and measurements will give as an output only statistical characterizations of the accuracy of estimated attributes, but they can never give absolute confidence.
-This is in stark contrast with the concept of unique identifiability used in legal documents, such as the General Data Protection Regulation (GDPR) by the European Union [eu2016gdpr], which does not explicitly leave room for statistical uncertainty.
-This is reflected for example in the Guidelines for virtual voice assistants by the European Data Protection Board, which states that: [edpb2021guidelines, page 13, §31]
+More recently, other countries have followed suit, for example through the “Personal Information Protection Law of the People’s Republic of China” [229] and “Lei Geral de Proteção de Dados” (Brazilian General Data Protection Law, LGPD) [230].
+With respect to regulation, an important consequence of the objective measures of privacy in section VI is that our tools and measurements will give as an output only statistical characterizations of the accuracy of estimated attributes, but they can never give absolute confidence.
+This is in stark contrast with the concept of unique identifiability used in legal documents, such as the General Data Protection Regulation (GDPR) by the European Union [226], which does not explicitly leave room for statistical uncertainty.
+This is reflected for example in the Guidelines for virtual voice assistants by the European Data Protection Board, which states that: [89, page 13, §31]
 … voice data is inherently biometric personal data. As a result, when such data is processed for the purpose of uniquely identifying a natural person … the processing must have a valid legal basis …
@@
 Both interpretations lead to absurdity, which suggests that the truth must lie somewhere in the middle.
-In fact, the GDPR in practice requires (see [edpb2021guidelines, page 4]) that the design process of voice assistants includes a data protection impact assessment, where the risks and consequences are evaluated such that the designer can take appropriate precautions to preserve privacy.
+In fact, the GDPR in practice requires (see [89, page 4]) that the design process of voice assistants includes a data protection impact assessment, where the risks and consequences are evaluated such that the designer can take appropriate precautions to preserve privacy.
 Authors of the GDPR are thus clearly aware that it is impossible to give absolute guarantees of privacy but that the impact assessment (i.e., objective measures of privacy) must necessarily be based on statistical measures, even if such measures have not been defined.
+Fortunately, first mathematical characterizations of the legal criteria have started to appear [169, 170].
 However, even in the absence of legal requirements, the current author would encourage all researchers, developers, data analysts, project managers, and policymakers to perform impact assessments in data projects, data management, and data policies, anytime when there is or could be the slightest concern that the dataset imposes risks to privacy.
-A practical tool to aid impact assessment is the “Data Ethics Decision Aid (DEDA)” that formalizes and enumerates all the questions that a representative group of stakeholders should review, discuss, and consider [franzke2021data].
+A practical tool to aid impact assessment is the “Data Ethics Decision Aid (DEDA)” that formalizes and enumerates all the questions that a representative group of stakeholders should review, discuss, and consider [231].
 While governments are in the process of regulating privacy, corporations and non-governmental organizations have also realized that proper privacy is an opportunity.
-For example, the Open Voice Network seeks to develop and standardize open technical standards and ethical guidelines for voice assistance [ovn2023] and the MyData Global seeks to help people and organizations to benefit from personal data in a human-centric way [mydata2023, poikola2020mydata].
-Within the research community, the author of this paper has been involved in establishing a special interest group within the International Speech Communication Association (ISCA) devoted to “Security and Privacy in Speech Communication” [spsc2023].
+For example, the Open Voice Network seeks to develop and standardize open technical standards and ethical guidelines for voice assistance [232] and the MyData Global seeks to help people and organizations to benefit from personal data in a human-centric way [233, 234].
+Within the research community, the author of this paper has been involved in establishing a special interest group within the International Speech Communication Association (ISCA) devoted to “Security and Privacy in Speech Communication” [235].
 It is as far as we know, the world-wide largest community focused on this topic.
@@
 Its most notable novelty is an exhaustive categorization of threats (see figs.
-1 and III).
-Protections against those threats are further categorized according to whether they relate to the private message or side information.
-The pertinent difference is that transmitting a private message is the whole purpose of communication, and there is not much we can do to protect it other than encryption.
+1 and IV).
+Protections against those threats are further categorized according to whether they relate to the target information or side information.
+The pertinent difference is that transmitting target information is the whole purpose of communication, and there is not much we can do to protect it other than encryption.
 With side-information, that is, all the other information that is bundled into a speech, like health status and gender identity, we have a much larger arsenal of protections.
 The primary approach is, however, to remove as much of the side information as possible as early as possible.
-As the private message is all that communication we need, all side-information should be removed to the extent possible.
-Such removal rapidly, however, demonstrates that paralinguistic information like speech style is often very useful in conveying the intended message.
-It is thus not always clear what constitutes a legitimate private message.
+As the target information is all the communication we need, all side-information should be removed to the extent possible.
+Such removal, however, rapidly demonstrates that paralinguistic information like speech style is often very useful in conveying the intended message.
+It is thus not always clear what constitutes the legitimate, target information.
 The first conclusion from this paper is that the range of possible threats to privacy is vast.
@@
 That means that we need a society-wide ethical discussion about what is allowed with respect to user privacy.
-Such discussions are needed to prevent a Cambridge Analytica-style scandal for speech technology [heikkila2023cambridge_ai].
+Such discussions are needed to prevent a Cambridge Analytica-style scandal for speech technology [219].
 A third implication of this paper is that, while research in this field has started to accelerate only recently, there is already a substantial body of research available.
@@
 Consent
-While management of acquiring informed consent has established traditions and best practices for most interface types [friedman2000informed], speech, audio, and ambient systems are notably unique.
+While management of acquiring informed consent has established traditions and best practices for most interface types [236], speech, audio, and ambient systems are notably unique.
 Namely, acoustic information is a time-varying stream.
@@
 Consent should thus be acquired on an actual need basis.
-In addition to being more usable, it would also make choices better connected to the actual needs since consent is acquired only once it is actually needed.
+In addition to being more usable, it would also make choices better connected to the actual needs since consent is acquired only once it is needed.
 Metrics for streaming
@@
 Metrics for out-of-category information
-The metrics discussed in section V are all related to specific categories of private information and in particular, we can provide protection only to identified threats.
+The metrics discussed in section VI are all related to specific categories of private information and in particular, we can provide protection only to identified threats.
 For example, we can measure the threat to privacy related to health information, but that does not say anything about the threat related to information about ethnic background.
-We thus need methods for evaluating privacy jointly with respect to all categories of private information except the private message.
+We thus need methods for evaluating privacy jointly with respect to all categories of private information except the target information.
 Future-proof metrics
@@
 The metrics are thus subject to change when those models are improved in the future, and it will likely expose new threats.
-Though it is likely difficult, it would be extremely useful if we could characterize, for example, privacy threats as a function of computational complexity, similarly as cryptography provides assurances over security based on the difficulty of solving certain difficult computational problems.
-Where cryptographic security is often seen as a threshold of computational complexity, however, for privacy in speech applications, it would be beneficial to also assess the leak magnitude (e.g.
-in bits) as a function of computational complexity.
-Over time, improvements in efficiency reduce algorithmic complexity, and as hardware with better capabilities becomes available, this would allow characterizing how protections will survive the test of time.
+Though it is likely difficult, it would be extremely useful if we could characterize, for example, privacy threats as a function of computational complexity, similarly to how cryptography provides assurances over security based on the difficulty of solving certain difficult computational problems.
+Where cryptographic security is often seen as a threshold of computational complexity, however, for privacy in speech applications, it would be beneficial to also assess the leak magnitude (e.g., in bits) as a function of computational complexity.
+Over time, efficiency improvements reduce algorithmic complexity, and as hardware with better capabilities becomes available, this would allow us to characterize how protections will survive the test of time.
 Attack and attacker characterization
-The biometrics community has a long tradition of characterizing attacks as, for example, opportunistic and motivated attacks, as well as in terms of the potential rewards of such attacks [roberts2007biometric].
+The biometrics community has a long tradition of characterizing attacks as, for example, opportunistic and motivated attacks, as well as in terms of the potential rewards of such attacks [237].
 Characterization of speech-specific attacks has received much less attention, though figs.
-1 and III in this paper are attempts in that direction.
+1 and IV in this paper are attempts in that direction.
 Further work is needed to gain privacy-protections on a systems level.
@@
 If user A records a discussion with user B, then both clearly have some level of ownership and privacy requirements on that recording.
-Another case is smart technology with multiple users, like smart TVs; even if one user has consented to data collection, that does not mean that others would agree.
+Another case is smart technology with multiple users, like smart TVs; even if one user has consented to data collection, that does not mean that others would agree [222].
 We do not yet have any widely accepted standard approaches for handling privacy, ownership, and consent in such multi-user scenarios.
@@
 Without exact definitions of those categories, we have no hope of developing methods for them.
-An alternative approach is to use representation learning methods to create unsupervised clustering of information categories, thus circumventing the need to explicitly define categories.
+An alternative approach is to use representation learning methods to create unsupervised clustering of information categories, thus circumventing the need to define categories explicitly.
 The compromise is that we cannot guarantee that the learned representations correspond to heuristically meaningful categories.
@@
 Perception, experience, and design of privacy
-Most of the speech-specific research on privacy has focused on privacy-preserving processing and systems structures.
+Most of the speech-specific research on privacy has focused on privacy-preserving processing and system structures.
 This is useful because it is the mandatory prerequisite for privacy-preserving technology.
-However, as discussed in section VI-B, users’ experience of services is to some extent independent of the objective level of privacy.
+However, as discussed in section VII-C, users’ experience of services is to some extent independent of the objective level of privacy.
 We need many more user studies on how, for example, voice characteristics and word choices influence trust, how the privacy level can be monitored during interactions and how changes are notified, how the environment and content of interaction influence user experiences, etc.
@@
 In conclusion, threats and breaches of privacy have significant negative consequences on individual, societal, ethical, and economic levels.
-While further improvements in smart technology are expected to improve the utility of the technology, it likely also introduce new threats.
-The protection of privacy in speech technology has thus been important already for a long time and the importance is increasing.
-Fortunately, research in the area has picked up speed and this tutorial presents the most important concepts, approaches, and methodology.
+While further improvements in smart technology are expected to improve the utility of the technology, it will likely also introduce new threats.
+The protection of privacy in speech technology has thus been important already for a long time, and its importance is increasing.
+Fortunately, research in the area has picked up speed, and this tutorial presents the most important concepts, approaches, and methodology.
 It is, however, likely that fundamental results and new technologies will be introduced in the near future.
@@
 A majority of the content in this paper has risen directly from discussions with a long list of colleagues, who were far too many to include as co-authors.
-These colleagues include, but are not limited to, in alphabetical order, Chris Brzuska, Mads Græsbøll Christensen, Sneha Das, Nicholas Evans, Johannes Fischer, Francisco Florez, Florin Ghido, Meeri Haataja, Ivan Habernal, Aki Härmä, Dorothea Kolossa, Michael Laakasuo, Martha Larson, Anna Leschanowsky, Rainer Martin, Joachim Meyer, Sebastian Möller, Andreas Nautsch, Birgit Popp, Nitin Sawhney, Ingo Siegert, Stephan Sigg, Rech Silas, Isabel Trancoso, Sanna Toropainen, Emmanuel Vincent, Jennifer Williams, and Pablo Pérez Zarazaga.
+These colleagues include, but are not limited to, in alphabetical order, Chris Brzuska, Mads Græsbøll Christensen, Sneha Das, Nicholas Evans, Johannes Fischer, Francisco Florez, Florin Ghido, Meeri Haataja, Ivan Habernal, Aki Härmä, Dorothea Kolossa, Michael Laakasuo, Martha Larson, Anna Leschanowsky, Rainer Martin, Joachim Meyer, Sebastian Möller, Andreas Nautsch, Francesco Nespoli, Birgit Popp, Silas Rech, Nitin Sawhney, Ingo Siegert, Stephan Sigg, Francisco Teixeira, Isabel Trancoso, Sanna Toropainen, Ville Vestman, Emmanuel Vincent, Jennifer Williams, and Pablo Pérez Zarazaga.
 I will remain forever grateful.
-[…]
-[Uncaptioned image] Tom Bäckström D.Sc.
-(tech.) is an associate professor at Aalto University, Finland (2019-).
-He obtained his Master’s and Doctoral degrees at Helsinki University of Technology (the predecessor of Aalto) in 2001 and 2004, respectively.
-During his time at International Audio Laboratories Erlangen, Germany (2008-2016), he made contributions to several international speech and audio coding standards such as MPEG USAC and 3GPP EVS, and became a professor (W2) at Friedrich-Alexander University Erlangen-Nürnberg (FAU) (2012-2016).
-Before his current position, he was a professor practice at Aalto University (2016-2019).
-He was the initiator and chair of ISCA SIG “Security and Privacy in Speech Communication” (2019-2022).
-His current research interests include privacy, coding, enhancement, and transmission of speech as well as machine learning.
+A part of this work was supported by “Designing Inclusive & Trustworthy Digital Public Services for Migrants in Finland (Trust-M)” (Grant Number: 353529) funded by the Strategic Research Council of Finland.

### v4

@@
 Such technology, however, also presents proven threats to privacy.
-Speech is a tool for communication and it will thus inherently contain private information.
-Importantly, it however also contains a wealth of side information, such as information related to health, emotions, affiliations, and relationships, all of which are private.
+Speech is a tool for communication, and it will thus inherently contain private information.
+Importantly, it also contains a wealth of side information, including details about health, emotions, affiliations, and relationships, all of which are private.
 Exposing such private information can lead to serious threats such as price gouging, harassment, extortion, and stalking.
-This paper is a tutorial on privacy issues related to speech technology, modeling their threats, approaches for protecting users’ privacy, measuring the performance of privacy-protecting methods, perception of privacy as well as societal and legal consequences.
+This paper is a tutorial on privacy issues related to speech technology, modeling their threats, approaches for protecting users’ privacy, measuring the performance of privacy-protecting methods, perception of privacy, as well as societal and legal consequences.
 In addition to a tutorial overview, it also presents lines for further development where improvements are most urgently needed.
@@
 Speech is a mode of communication and thus inherently contains a wealth of information (see table I).
-In addition to the desired information of the communication, speech thus also contains a wide range of side information like the state of health and emotions, as well as physical, psychological, and social identity, most of which is private information.
+In addition, to the desired information of the communication, speech thus also contains a wide range of side information like the state of health and emotions, as well as physical, psychological, and social identity, most of which is private information.
 Moreover, speaking is a dynamic interaction between two or more speakers.
@@
 From an information content perspective, we can thus expect that privacy is a multifaceted issue in all areas of speech technology.
-TABLE I: Private information potentially identifiable from speech signals, and the extent to which they are sustained over time and can be willfully controlled [1, 2, 3, 4].
+TABLE I: Private information potentially identifiable from speech signals, and the extent to which they are sustained over time and can be willfully controlled [singh2019profiling, kroger2019privacy, kroger2022personal, singh2025humanvoiceunique].
 Category Examples Permanence Control
@@
 dignity
-The subjective feeling of being continuously monitored may cause psychological damage and stifle political expression damaging democratic societies.
+The subjective feeling of being continuously monitored may cause psychological damage and stifle political expression, damaging democratic societies.
 Users may feel they lose control of personal aspects of their lives.
@@
 The list is incomplete, and we can expect new threats to be discovered.
-The generic solution template is, however, typically always the same: Minimize transmission and storage of as well as access to sensitive information that is irrelevant to the service that the users want.
+The generic solution template is, however, typically always the same: Minimize transmission, storage, and access to sensitive information that is irrelevant to the service that the users want.
 It is then ”merely” a question of how such minimization is achieved, what information is relevant, and how to determine what the user wants (provide information and enable control of services and threats).
-With respect to threats, media attention is often focused on breaches which have large economic consequences [5, 6, 7].
+With respect to threats, media attention is often focused on breaches which have large economic consequences [Lynskey2019GuardianPrivacyScandal, hern2019, brewster2021].
 While such breaches are important by themselves, the media attention introduces unfortunate biases in two ways.
@@
 Importantly, users can avoid systems that are perceived threatening, even when there is no actual threat.
-To maintain users’ trust, it is therefore important to both uphold the users’ actual privacy but also design systems to actively and frequently communicate the level of privacy, including known threats, their current status, and measures taken to protect against them.
-Second, while a single breach in a large service can have large economic, psychological, societal, and legal consequences, small crimes are so common that their joint effect is comparable or larger in size [8].
+To maintain users’ trust, it is therefore essential to both uphold the users’ actual privacy but also design systems that actively and frequently communicate the level of privacy, including known threats, their current status, and measures taken to protect against them.
+Second, while a single breach in a large service can have large economic, psychological, societal, and legal consequences, small crimes are so common that their joint effect is comparable or larger in size [levy2020intimate].
 With speech technology, such ”small” crimes include stalking, extortion, harassment, humiliation, and inappropriate advances (see table II).
-While the economic damage of a single such incident can be small, when combined, their joint psychological and societal effect is potentially large and their prevalence makes them a considerable threat also economically.
+While the economic damage of a single such incident can be small, when combined, their joint psychological and societal effect is potentially large, and their prevalence makes them a considerable threat also economically.
 This paper presents a tutorial overview of privacy in speech technology that covers a wide range of threats, methodologies, and algorithms.
 The range of applications and content types where privacy-preserving processing methods are needed is vast, and section II presents a brief overview thereof, and section III defines the vocabulary.
-Analysis of threats demonstrates that while attack surfaces have great variety, the threat models are similar (see section IV).
-Protections against those threats have distinct categories of approaches, isolating target information by removing private information, limiting access to private the message through cryptography, reducing centralized threats through non-centralized processing, limiting access to reproduced audio, improving overall system performance, and design principles from theoretical frameworks (see section V).
+Analysis of threats demonstrates that while attack surfaces have a great variety, the threat models are similar (see section IV).
+Protections against those threats have distinct categories of approaches, isolating target information by removing private information, limiting access to private information through cryptography, reducing centralized threats through non-centralized processing, limiting access to reproduced audio, improving overall system performance, and design principles from theoretical frameworks (see section V).
 To quantify the extent of protections, we further need methods for evaluation of privacy and utility (see section VI).
@@
 For the best user experience, therefore, we need to quantify and understand users’ perceptions, experiences, and preferences related to privacy, as well as design systems accordingly (see section VII).
-Finally, threats to and breaches in privacy have considerable consequences on both societal, economic, and individual levels, which has motivated governments to increasingly regulate the use of technology (see section VIII).
+Finally, threats to and breaches in privacy have considerable consequences on both societal, economic, and individual levels, which have motivated governments to increasingly regulate the use of technology (see section VIII).
 This paper focuses primarily on speech technology in human-computer interfaces (see fig.
 1), but will also address telecommunication where appropriate.
-Focus is here, however, limited to the acoustic speech signal, also known as the voice since natural language processing is a distinct and largely independent field that warrants treating it separately (e.g.
-[9]).
-Concerning telecommunication, approaches to ethical, legal, and technological questions related to privacy in telecommunication over landlines are well-established, and open discussions are primarily related to how the existing regulation and oversight should be extended to cover also mobile telecommunication and voice-over-Internet protocols (VoIP) [10, 11, 12].
+Focus is here, however, limited to the acoustic speech signal, also known as the voice, since natural language processing is a distinct and largely independent field that warrants treating it separately (e.g.
+[mahendran2021privacy]).
+Concerning telecommunication, approaches to ethical, legal, and technological questions related to privacy in telecommunication over landlines are well-established, and open discussions are primarily related to how the existing regulation and oversight should be extended also to cover mobile telecommunication and voice-over-Internet protocols (VoIP) [penders2004privacy, wong2009privacy, arapinis2017analysis].
 To the author’s knowledge, this is the widest tutorial overview of privacy in speech technology.
-Recent works in the field, however, include a technical overview [13], a survey on privacy issues with voice assistants [14], as well as a popular-science review [15] of this area, and several doctoral theses have their respective summaries, e.g.
-[16, 17, 18, 19, 20, 21, 22, 23].
-Privacy has been extensively discussed in other areas of science, such as the neighboring field of natural language processing [9], in statistical theory of privacy [24], in social sciences [25] and psychology [26].
-There is even an excellent and thorough meta-study of all areas of science that discuss privacy [27].
+Recent works in the field, however, include a technical overview [nautsch2019preserving], a survey on privacy issues with voice assistants [acosta2022survey], as well as a popular-science review [backstrom2020privacy] of this area, and several doctoral theses have their respective summaries, e.g.
+[Nelus2022, das2021thesis, zarazaga2022thesis, pathak2012privacy, williams2022learning, portelo2015privacy, vestman2020phd, nespoli2025phd].
+Privacy has been extensively discussed in other areas of science, such as the neighboring field of natural language processing [mahendran2021privacy], in statistical theory of privacy [cummings2023challenges], in social sciences [petronio2002boundaries] and psychology [Derlega1977PrivacyAS].
+There is even an excellent and thorough meta-study of all areas of science that discuss privacy [van2018handbook].
 Central contributions of this work include 1.
@@
 II Content categories and applications
-To give a complete picture of the implications of privacy in speech technology, this section provides a brief discussion both on categories of private information, complementing table I as well as the application of such information.
+To give a complete picture of the implications of privacy in speech technology, this section provides a brief discussion on both categories of private information, complementing table I as well as the application of such information.
 Observe that this is not a complete list (nor is table I) of the content nor of application categories.
@@
 We can attempt to recognize who is speaking (speaker recognition), verify whether a speaker has the claimed identity (speaker verification), clustering audio to segments with a single speaker (speaker diarization) and we can develop methods for deceiving identity (spoofing), e.g.
-[28, 37, 44, 44, 45, 46, 47, 48, 49, 29, 30, 31, 32, 46, 47, 33, 34, 35, 36, 38, 39, 40, 41, 42, 43].
+[chouchane21_interspeech, granqvist20_interspeech, mawalim20_interspeech, mawalim20_interspeech, nautsch19b_interspeech, nelus19_interspeech, nelus19b_interspeech, noe21_interspeech, perero2022x, zewoudie2021federated, zewoudie2021use, teixeira22_interspeech, treiber2019privacy, nelus19_interspeech, nelus19b_interspeech, BAI202165, han2020voice, KINNUNEN201012, hansen2015speaker, parthasarathi2013wordless, pathak11_interspeech, parthasarathi11_interspeech, pathak2012ppsa, jimenez2017two, portelo2014bprivacy].
 Similarly, by voice conversion we can anonymize a speaker identity by replacing it with a random identity (anonymization) or a specific one (pseudonymization), e.g.
-[50, 51, 52, 53, 54, 55, 44, 56, 57].
-Speaker characterization is the natural complement of speaker idenfication [13, 58].
-Such methods related to speaker identity can be applied for example to recognize the user of a device, verify a customer at a bank, or as a voice avatar in online gaming.
+[noe20_interspeech, pierre22_interspeech, prajapati21_interspeech, shamsabadi2023diff, srivastava20_interspeech, Srivastava2022taslp, mawalim20_interspeech, maouche20_interspeech, Fang2019].
+Speaker characterization is the natural complement of speaker idenfication [nautsch2019preserving, nguyen2010automatic].
+Such methods related to speaker identity can be applied, for example, to recognize the user of a device, verify a customer at a bank, or as a voice avatar in online gaming.
 Similarly, anonymization and pseudonymization can be used to hide the identity from the public media and gaming.
 Speech recognition, as in speech-to-text, is probably the largest sub-area of speech research.
-With respect to privacy, it contains two obvious challenges for privacy.
+With respect to privacy, it presents two obvious challenges.
 We can try to limit side-information, such as eliminating all non-text information from the data stream, e.g.
-[59, 60, 61, 62, 63, 64, 65], or we can use natural language processing to anonymize the text content, e.g.
-[66, 67, 68, 69, 9, 70].
-Privacy is even more important in always-on applications like wake-word detection, i.e.
-when the interface is triggered by a specific keyword like “Computer” in “Computer, lights-off”.
-This application is more sensitive exactly because it is always on as users cannot choose when data is processed [71, 72, 73, 74].
-The always-on characteristic is also prominent in assisted-living applications, which can “monitor people’s daily exercises, consumption of calories and sleeping patterns, and to provide coaching interventions to foster positive behavior” [75, 76].
-Such ambient voice interfaces are often implemented through acoustic sensor networks which pose their own challenges [72, 77, 78, 30, 79, 80, 81].
-Speech enhancement refers to removing background noises and distortions from the desired speech signal (target information) [82].
+[shoemate2022sottovoce, nguyen2023federated, aloufi21_interspeech, garg20b_interspeech, shah21_interspeech, srivastava19_interspeech, tomashenko2022privacy], or we can use natural language processing to anonymize the text content, e.g.
+[liu22_interspeech, maouche2022enhancing, adelani20_interspeech, novotney21_interspeech, mahendran2021privacy, habernal2021differential].
+Privacy is even more critical in always-on applications like wake-word detection, i.e., when the interface is triggered by a specific keyword like “Computer” in “Computer, lights-off”.
+This application is more sensitive exactly because it is always on, as users cannot choose when data is processed [koppelmann21_interspeech, koppelmann22_interspeech, stoidis22_interspeech, yang21_interspeech].
+The always-on characteristic is also prominent in assisted-living applications, which can “monitor people’s daily exercises, consumption of calories and sleeping patterns, and to provide coaching interventions to foster positive behavior” [haider19_interspeech, salah2025privacy].
+Such ambient voice interfaces are often implemented through acoustic sensor networks, which pose their own challenges [koppelmann22_interspeech, oconnor20_interspeech, malekzadeh2018replacement, zewoudie2021use, zarazaga2020acoustic, nelus2021, backstrom2025privacy].
+Speech enhancement refers to removing background noises and distortions from the desired speech signal (target information) [benesty2008springer].
 This task can have an impact on privacy in two primary ways.
-First, the recording environment can reveal private information about the speaker, and second, background noises, like a competing talker, can contain private information [83, 84].
+First, the recording environment can reveal private information about the speaker, and second, background noises, like a competing talker, can contain private information [zarazaga21_interspeech, rech2022multi].
 Attenuating or removing background noises and competing speakers, as well as removing reverberation, can thus improve privacy.
-In addition, while using a sensor network to capture speech can improve utility, it introduces novel threats as well [80, 18, 85].
-Speech is typically considered to be biometric information in its entirety, including all information categories in table I as such categories can be used to identify a person [86, 87, 13, 88, 89].
+In addition, while using a sensor network to capture speech can improve utility, it introduces novel threats as well [nelus2021, zarazaga2022thesis, hendriks2013ppdistributed].
+Speech is typically considered to be biometric information in its entirety, including all information categories in table I as such categories can be used to identify a person [adler2006towards, nautsch20_interspeech, nautsch2019preserving, ISObiometric, edpb2021guidelines].
 Some, however, limit the term biometric to refer only to physical and behavioral characteristics.
-Such properties include health [90], emotions [91] and gender identity [92, 93, 73, 46], each warranting their own treatment.
+Such properties include health [natarajan2021prioris], emotions [feng22b_interspeech] and gender identity [nelus2018itg, stoidis21_interspeech, stoidis22_interspeech, nelus19_interspeech], each warranting their own treatment.
 Bob Acoustic pathway User
@@
 XX5765432X1
-Figure 1: Threats to the privacy of a user Alice when speaking with another person Bob or a local device, where information is leaked or shared through the acoustic pathways, the local device, the network or the cloud service, to another person Eve, device or service, contrary to the preferences of Alice (red arrows).
+Figure 1: Threats to the privacy of a user Alice when speaking with another person Bob or a local device, where information is leaked or shared through the acoustic pathways, the local device, the network, or the cloud service, to another person Eve, device or service, contrary to the preferences of Alice (red arrows).
 Each threat is marked with the number of the corresponding part of section IV-D.
-Dotted red arrows marked with “X” indicate threats outside the current scope and dashed black arrows indicate the potential retransmission of leaked data beyond the primary attack surface.
+Dotted red arrows marked with “X” indicate threats outside the current scope, and dashed black arrows indicate the potential retransmission of leaked data beyond the primary attack surface.
 III Definitions
@@
 Here, information security refers to enforcing rightful access to information in scenarios where the access rights are binary (granted or not granted).
-In contrast, information privacy refers to the application of rules to information access and use, in scenarios where access permissions are partial and rules context-dependent (adapted from [94]).
+In contrast, information privacy refers to the application of rules to information access and use, in scenarios where access permissions are partial and rules context-dependent (adapted from [nissenbaum2004privacy]).
 The rules may distinguish between the ability to access and de facto accessing the data, different ways and purposes of using the data, and sharing the data with third parties.
-The rules may further be explicit (e.g., verbalized or documented) or implicit (as based on, e.g., social norms and tacit expectations) and can include requirements on monitoring, reporting, and control.
+The rules may be explicit (e.g., verbalized or documented) or implicit (based on social norms and tacit expectations), and can include requirements for monitoring, reporting, and control.
 For example, a voice interface can be used to control home automation, but if the service provider shares that voice data with advertisers against the users’ preferences, then it violates privacy (see fig.
@@
 III-B Information categories
-Target information is the content of a communication that is required to fulfill the purpose of said communication.
-Side-information (also known as auxiliary information, external knowledge, or background knowledge [95]) is any and all information in a communication other than the target information.
+Target information is the content of a message that is required to fulfill the purpose of that communication.
+Side-information (also known as auxiliary information, external knowledge, or background knowledge [ganta2008composition]) is any and all information in a message other than the target information.
 See fig.
 2.
-Attributes (or characteristics [4]) of speech are properties that can be distinguished from a speech signal and describe the speaker(s) or their state, the speech content, style or environment, the recording configuration, or any other related features.
+Attributes (or characteristics [singh2025humanvoiceunique]) of speech are properties that can be distinguished from a speech signal and describe the speaker(s) or their state, the speech content, style, or environment, the recording configuration, or any other related features.
 Table I lists examples of speech attributes with private information.
@@
 For example, in medical analysis, the target information is the health state of the speaker.
-In that case, the lexical information may be private side-information.
+In that case, the lexical information may be private side information.
 Sender Channel RecipientRefer to caption
@@
 !! Target information Side-information
-Figure 2: The high-level threat model of speech interaction, where a target information is sent through a channel to the legitimate recipient, but consequential side-information is bundled to that message.
+Figure 2: The high-level threat model of speech interaction, where target information is sent through a channel to the legitimate recipient, but consequential side-information is bundled to that message.
 It is a threat to privacy when an undesired recipient gains access to that target information or side information (marked by red arrows and exclamation marks).
@@
 In contrast, we use threat tasks to model the activity of attackers trying to extract information or perform other nefarious tasks.
-[47, 92].
+[nelus19b_interspeech, nelus2018itg].
 Voice conversion refers to any speech processing method that replaces the speaker’s identity or other attributes in a speech signal with those of another.
@@
 The anonymized speech signal then corresponds to a speech signal of an artificial person, who is known as a pseudospeaker.
-Attribute or property inference refers to the identification of private attributes of speakers of known identity or otherwise [96].
-Re-identification refers to identifying individuals from apparently anonymized information [88, 97].
+Attribute or property inference refers to the identification of private attributes of speakers of known identity or otherwise [gong2018attribute].
+Re-identification refers to identifying individuals from apparently anonymized information [ISObiometric, ISOIEC224745_2022].
 In anonymization tasks, enrollment data is reference data, typically labeled with, for example, speakers’ identities or other attribute(s), to which observations are compared with the purpose of identifying said identity or attribute(s).
@@
 Conversely, speech samples that are compared to a database of samples are here referred to as observations, and in literature, they are sometimes known as trial samples, recognition samples, or public data.
-[98]
+[tomashenko2022voiceprivacy]
 III-D Attack and attacker properties
 We use the terms attacker and unintended recipient loosely to describe any entity or agent that gains access to or uses private information in a way that conflicts with the user’s preferences.
-A malicious attacker is intentionally seeking to gain private information, a targeted attacker attempts to gain such access to a specific person, an opportunistic attacker to any person, a motivated attacker is seeking ways to attack, and a honest-but-curious attacker seizes emerging opportunities [99, 100].
-The attack surface or vector is the specific point in the system where private information leaves the authorized path to make its way to the attacker [101].
+A malicious attacker is intentionally seeking to gain private information, a targeted attacker attempts to gain such access to a specific person, an opportunistic attacker to any person, a motivated attacker is seeking ways to attack, and a honest-but-curious attacker seizes emerging opportunities [CASEY20038, le2023privacy].
+The attack surface or vector is the specific point in the system where private information leaves the authorized path to make its way to the attacker [THEISEN201894].
 IV Threat and attack models
@@
 Speech data can be exposed in two ways (see fig.
-2): First, when a target information is transmitted, it is a threat to privacy when contrary to the preferences of the user, it is used in an unexpected way or transmitted to a third party.
+2): First, when target information is transmitted, it is a threat to privacy when, contrary to the preferences of the user, it is used in an unexpected way or transmitted to a third party.
 Second, since speech contains a wide range of private information (see table I) which is bundled in the acoustic signal in complicated ways, it is challenging to extract only the desired message.
@@
 It is a threat to privacy when such private side-information is transmitted to an undesired recipient alongside the target information.
-Per definition, we here assume that the legitimate recipient can be trusted with the side information and that any undesirable use of it is labeled as an undesirable service.
+By definition, we here assume that the legitimate recipient can be trusted with the side information and that any undesirable use of it is labeled as an undesirable service.
 The difference between the two forms of information makes an important difference in approaches to mitigation.
-Transmitting the target information to the recipient is the desired outcome of the interaction and we can thus not complete the desired task without sending the message.
-In some cases, though, we can reduce the amount of information content in the target information, and more generally, we should always secure the transmission for example through cryptography (see sections V-A, V-B and V-F).
+Transmitting the target information to the recipient is the desired outcome of the interaction, and we can thus not complete the desired task without sending the message.
+In some cases, though, we can reduce the amount of information content in the target information, and more generally, we should always secure the transmission, for example, through cryptography (see sections V-A, V-B and V-F).
 Some cryptographic methods, such as homomorphic encryption, even keep the message secret from the receiver but allow the processing of a desired task.
@@
 IV-B Attack types
-There are three main types of attacks; motivated, opportunistic, and incidental.
+There are three main types of attacks: motivated, opportunistic, and incidental.
 In motivated attacks, the attacker attempts to extract information about a specific person.
@@
 For example, if a database contains speech samples from the members of a particular association, then the re-identification of a target person reveals that person’s affiliation with that association.
-Opportunistic attacks are usually known as birthday attacks, referring to the birthday paradox [102], where the attacker uses a known weakness on a large pool of potential victims.
-It is like spam email; if the attacker sends a large enough number of such emails, then randomly, someone triggers the trap by mistake.
+In opportunistic attacks, the attacker uses a known weakness on a large pool of potential victims.
+It is similar to spam email; if the attacker sends a large enough number of such emails, then, randomly, someone triggers the trap by mistake as in a birthday attack [menezes2018handbook].
 In this form of attack, anyone could be the victim.
 For example, a speech service could analyze private speaker attributes (for, e.g., health problems) or affiliations (with, e.g., minority groups), for purposes of targeted advertisement or surveillance.
-Where motivated attacks typically target high-value victims such as famous persons, where both protections and rewards are high, birthday attacks typically target ordinary people, where the likelihood of success is larger but rewards are smaller.
+Where motivated attacks typically target high-value victims such as famous persons, where both protections and rewards are high, opportunistic attacks typically target ordinary people, where the likelihood of success is larger but rewards are smaller.
 The third category of “attacks” is incidental, where an agent gains excessive access to information unintentionally or by mistake.
@@
 In other words, by putting themselves in the attacker’s place, safeguards designers can better understand pertinent aspects of weaknesses and quantify risks.
-Figure 3 illustrates an attack model for measuring the extent of privacy [56].
+Figure 3 illustrates an attack model for measuring the extent of privacy [maouche20_interspeech].
 We assume that there is some private data, which is anonymized such that it can be used in a trusted task.
@@
 The attacker has also access to some other data about speakers, found from some other source (found data), which helps in extracting private information.
-Again, the term “found data” is a loose term and denotes any data other than found data that is available to the attacker.
-To complement an Attack(er) model, it can also be useful to define a Protector model for the use-case of the system.
-Jointly, the two can be documented using the Scenario-of-Use scheme, as illustrated in table III [103].
-The accurate definition of the protector model and use-cases is essential, since they define both the target information needed as well as how performance should be evaluated [104].
-TABLE III: The Scenario-of-Use scheme [103].
+Again, the term “found data” is a loose term and denotes any data other than public data that is available to the attacker.
+To complement an Attack(er) model, it can also be helpful to define a Protector model for the use case of the system.
+Jointly, the two can be documented using the Scenario-of-Use scheme, as illustrated in table III [rahman24_spsc].
+The accurate definition of the protector model and use-cases is essential, since they define both the target information needed as well as how performance should be evaluated [meyer2025usecasesvoiceanonymization].
+TABLE III: The Scenario-of-Use scheme [rahman24_spsc].
 Attacker Model (Attack on Privacy) Protector Model (Protection of Privacy)
@@
 Attacks can be classified according to the amount of information available for the attacker, such as information about the speakers and about the models used.
-Access to speech samples from the target user (known as enrollment data) are particularly important for the attacker.
-For example, in the anonymization task of the VoicePrivacy 2022 challenge, the objective was to replace (pseudonymize) speaker identity but retain all other speech characteristics such as linguistic information.
+Access to speech samples from the target user (known as enrollment data) is particularly important for the attacker.
+For example, in the anonymization task of the VoicePrivacy 2022 challenge, the objective was to replace (pseudonymize) speaker identity but retain all other speech characteristics, such as linguistic information.
 The anonymized sentences are known as trial utterances.
-The attack scenarios were then classified as [98]:
+The attack scenarios were then classified as [tomashenko2022voiceprivacy]:
   1. 1.
@@
   3. 3.
-    Lazy-informed: Both trial and enrollment data are available to attackers, but anonymized with different pseudo-speaker.
+    Lazy-informed: Both trial and enrollment data are available to attackers, but anonymized with a different pseudo-speaker.
   4. 4.
-    Semi-informed: As in lazy-informed, attackers can also train their model with the same anonymization system but different pseudo-speakers.
-With increasing information, obviously, the attacker’s task becomes easier and the accuracy of the attacker’s models improves.
+    Semi-informed: Similar to lazy-informed, attackers can also train their model using the same anonymization system but with different pseudo-speakers.
+With increasing information, obviously, the attacker’s task becomes easier, and the accuracy of the attacker’s models improves.
 Further, such attack scenarios can then be devised according to the specific use case.
-For example, if the task is to anonymize emotional or health status, we can define different attack scenarios depending on the extent to which the attacker has information about categories of emotions and health status used in anonymization.
+For example, suppose the task is to anonymize emotional or health status.
+In that case, we can define different attack scenarios depending on the extent to which the attacker has information about categories of emotions and health status used in anonymization.
 Private data Anonymization Public data Downstream application ! Private
@@
 An attacker uses any available (found) data and anonymized public data to infer private information contrary to the users’ preferences.
-Anonymized data flow is indicated by dashed lines and the attack by red lines and an exclamation mark.
+Anonymized data flow is indicated by dashed lines, and the attack is represented by red lines with an exclamation mark.
 IV-D Attack Surfaces
@@
 The best quality data correspond closely to scenarios where the services are used.
-Recording users’ interactions with their devices is then attractive for training improved models since it corresponds exactly to the use case.
-This presents a considerable threat to privacy because such unrestricted recording could capture a wide range of private information, including all interactions with the device but potentially also any and all speech in the vicinity of the device.
+Recording users’ interactions with their devices is then attractive for training improved models since it corresponds precisely to the use case.
+This presents a considerable threat to privacy because such unrestricted recording could capture a wide range of private information, including all interactions with the device, but potentially also any and all speech in the vicinity of the device.
 Observe that though figs.
@@
 5).
-A popular approach is to use a specific utterance, known as a wake word, to start all interactions with the interface [105].
-The wake word is then like a rudimentary password, which prevents the interface from activating when speech is not directed to the device.
+A popular approach is to use a specific utterance, known as a wake word, to start all interactions with the interface [wang2021wake].
+The wake word serves as a rudimentary password, preventing the interface from activating when speech is not directed to the device.
 Unfortunately, designing wake word detectors is non-trivial, and they will occasionally make mistakes.
@@
 While false negatives are annoying for the user when the service does not activate, false positives potentially present serious threats to privacy.
-In some famous cases, speech interfaces have activated from sounds on the television to buy unwanted items at the users’ behest, and users’ private conversations have been leaked to third parties [106, 107].
+In some famous cases, speech interfaces have been activated from sounds on the television to buy unwanted items at the users’ behest, and users’ private conversations have been leaked to third parties [liptak2017alexa, moscaritolo2018alexa].
 User Local device AliceRefer to caption Refer to caption EveRefer to caption Refer to caption Remote service Cloud storage Refer to caption Refer to caption !!
@@
 This attack surface is similar to the Cloud leak scenario, with the main difference being the recipient, which is here a person, whereas, in the Cloud leak scenario, it is an automated agent.
-An example of this threat is medical records, which can be collected from patients to a centralized database.
-A researcher with authorization to access the database can then potentially extract private information beyond the expected [108].
+An example of this threat is medical records, which can be collected from patients and stored in a centralized database.
+A researcher with authorization to access the database can then potentially extract private information beyond the expected [Teixeira2018].
 User Devices in same network Remote service Refer to caption Refer to captionRefer to captionRefer to caption Refer to caption Refer to caption Refer to caption ! Outside the room Room
@@
 IV-D4 WASN authentication
-The audio quality of speech pickup (in terms of, for example, signal-to-noise ratio or perceptual quality), as well as the usability of voice interfaces, can be improved by using all available connected devices with microphones in the same room or acoustic space [109, 79, 110].
+The audio quality of speech pickup (in terms of, for example, signal-to-noise ratio or perceptual quality), as well as the usability of voice interfaces, can be improved by using all available connected devices with microphones in the same room or acoustic space [das2021enhancement, zarazaga2020acoustic, sigg2020provable].
 Such collaboration can be realized with wireless acoustic sensor networks (WASNs), where several independent devices simultaneously pick up speech, and those channels are combined to obtain a high-quality signal (see fig.
-7[16, 18].
-A sensitive question is, however, authorization: Which devices are allowed to share information with each other?
-One approach is to assume that devices that are in the same room or acoustic space can hear the same signal [18, 72].
+7[Nelus2022, zarazaga2022thesis].
+A sensitive question, however, is authorization: Which devices are allowed to share information?
+One approach is to assume that devices that are in the same room or acoustic space can hear the same signal [zarazaga2022thesis, koppelmann22_interspeech].
 Presence in the same space is thus already an implicit authorization to participate in a joint signal pickup.
@@
 In the case where Eve overhears Alice’s speech, we assume that people have a learned awareness of which other people are present in the same room and adjust their speech accordingly.
-Then either Alice does not mind that Eve overhears her speech (inconsequential information), Alice can change her speech style to a whisper, or change content such that Eve does not hear anything private (reduced information transmission), or she can go to a different room to continue the interaction in private (modified acoustic channel).
+Then, either Alice does not mind that Eve overhears her speech (inconsequential information), Alice can change her speech style to a whisper, or change the content such that Eve does not hear anything private (reduced information transmission), or she can go to a different room to continue the interaction in private (modified acoustic channel).
 Whether users are similarly aware of devices present in the room, through anthropomorphization of devices or otherwise, and adjust their speech accordingly is an open question.
 Local dialogue Local device Remote service Alice Bob Refer to caption Refer to caption Refer to caption Refer to caption !!
-Figure 9: Threat scenario “Discussion leak”, where a dialogue between two users, Alice and Bob, is picked up by a local device and potentially shared to a remote service contrary to preferences (red arrow and exclamation mark).
+Figure 9: Threat scenario “Discussion leak”, where a dialogue between two users, Alice and Bob, is picked up by a local device and potentially shared with a remote service contrary to preferences (red arrow and exclamation mark).
 IV-D6 Discussion leak
@@
 The main difference to a Speech interface leak (sections IV-D5 and 8) is that the type of speech is different.
-Interaction with a speech interface at the current level of technology is, roughly speaking, limited to speech commands, question-answer pairs, and one-sided monologues, whereas spoken human-to-human interaction has potentially a much more dynamic character.
-This is potentially beneficial in the sense that the speaking style of human-to-human interaction could be recognized, preventing unintended triggering of functionalities.
+Interaction with a speech interface at the current level of technology is, roughly speaking, limited to speech commands, question-answer pairs, and one-sided monologues.
+In contrast, spoken human-to-human interaction has potentially a much more dynamic character.
+This is potentially beneficial in that the speaking style of human-to-human interaction can be recognized, thereby preventing unintended triggering of functionalities.
 UserAliceRefer to caption EveRefer to caption Local device Storage Refer to caption Refer to caption !
@@
 Many practical scenarios include multiple users sharing one or several speech-operated devices.
-For example, a family can share a smart speaker or television, an office meeting room can have smart devices and customer service points can have a phone shared across duty officers.
+For example, a family can share a smart speaker or television, an office meeting room can have smart devices, and customer service points can have a phone shared across duty officers.
 Each of these devices can collect private information about the user(s) over time and can potentially share it with other users (see fig.
 10).
-Notably, this scenario highlights that leak happens over a distance in time, whereas threats occurring over a spatial distance often receive most of the attention.
+Notably, this scenario highlights that a leak happens over a distance in time, whereas threats occurring over a spatial distance often receive most of the attention.
 This threat model clearly demonstrates that devices and services used by multiple users need to employ access control and authorization management if they store any private information.
@@
 By isolation of information, we refer to the design of a channel of information where the target information is isolated from other (undesired) private information.
-This applies at least and in particular to the attack vectors Cloud leak (section IV-D1), False activation (section IV-D2) and Cloud access (section IV-D3), where a reduction in flow of private information would reduce threats to privacy.
+This applies at least and in particular to the attack vectors Cloud leak (section IV-D1), False activation (section IV-D2), and Cloud access (section IV-D3), where a reduction in flow of private information would reduce threats to privacy.
 Central concepts in information isolation include 1.
-information bottlenecks(or privacy funnels) , where information is compressed to a low-capacity channel such that only target information is passed through (see section V-A1) and where the neural network implementation usually takes the structure of an autoencoder [111], 2.
+information bottlenecks(or privacy funnels) , where information is compressed to a low-capacity channel such that only target information is passed through (see section V-A1) and where the neural network implementation usually takes the structure of an autoencoder [li2023comprehensive], 2.
 adversarial training, where an information channel (such as an information bottleneck) is trained to contain target information to fulfill the trusted task and simultaneously trained such that a competing, attacking adversarial model fails to extract private information (threat task), (see section V-A2) and 3.
@@
 The overall approach is generic and can be applied, for example, in cases where 1.
-the decoder resynthesizes the signal with some attributes anonymized [112] or 2.
-information extraction scenarios such as classification [113], where the decoder can often be trivial.
-A concept related to information bottlenecks is the privacy funnel [114].
-Where the original definition of the information bottleneck [113] aims to maximize the utility of a channel with a fixed rate, the privacy funnel aims to minimize the information rate with a fixed utility.
+the decoder resynthesizes the signal with some attributes anonymized [wang2019vector] or 2.
+information extraction scenarios such as classification [tishby2000information], where the decoder can often be trivial.
+A concept related to information bottlenecks is the privacy funnel [makhdoumi2014information].
+Where the original definition of the information bottleneck [tishby2000information] aims to maximize the utility of a channel with a fixed rate, the privacy funnel seeks to minimize the information rate with a fixed utility.
 The information bottleneck is, however, much more widely known in the community and often informally used as an umbrella term to cover both cases.
-To distill only the target information and discard any side-information, speech information can be passed through a bottleneck so tight, that only the legitimate message at full accuracy can pass through [115, 114] (see fig.
+To distill only the target information and discard any side-information, speech information can be passed through a bottleneck so tight that only the legitimate message at full accuracy can pass through [Tishby2015, makhdoumi2014information] (see fig.
 11).
-This approach thus provides protection against an attacker who has access to the output of the bottleneck.
-The challenges are to design a training methodology and model structure with a sufficiently tight bottleneck so that the legitimate message is accurately retained while private side information is discarded.
-Observe that a generic implementation of an information bottleneck does not guarantee privacy as side-information can sneak through the bottleneck alongside the target information [116].
+This approach thus protects against an attacker who has access to the output of the bottleneck.
+The challenges lie in designing a training methodology and model structure with a sufficiently tight bottleneck, so that the legitimate message is accurately retained while private side information is discarded.
+Observe that a generic implementation of an information bottleneck does not guarantee privacy as side-information can sneak through the bottleneck alongside the target information [raj2019probing].
 Instead, to be privacy preserving, a channel must be designed or optimized for privacy, warranting the term privacy funnel.
-A frequently used approach to implementing a bottleneck is an autoencoder structure [111], consisting of an encoder, bottleneck, and a decoder, where the objective is to reconstruct the input signal from the bottleneck output [78, 49, 117] (see fig.
+A frequently used approach to implementing a bottleneck is an autoencoder structure [li2023comprehensive], consisting of an encoder, bottleneck, and a decoder, where the objective is to reconstruct the input signal from the bottleneck output [malekzadeh2018replacement, perero2022x, lu2013speech] (see fig.
 12).
@@
 Designing a loss function is then challenging when the bottleneck is so tight that the decoded signal is very far from the original.
-One solution akin to the machine learning method known as representation learning is to feed the output of the autoencoder again to the encoder and compare the bottleneck features [118, 119].
+One solution akin to the machine learning method known as representation learning is to feed the output of the autoencoder again to the encoder and compare the bottleneck features [neumann2019improving, chorowski2019unsupervised].
 Information content at the bottleneck can be reduced by either reducing the signal rank (viz.
-dimensionality reduction, e.g., [49]) or by quantizing and coding the signal at a low bitrate (e.g., [112, 120]).
-The trade-off between the accuracy of trusted task and bottleneck entropy then determines the extent of privacy (see section VI).
-Observe that dimensionality reduction is, in general, not alone sufficient because if the channel’s floating- (or fixed)-point representation has $K$ dimensions each with $B$ bits, then a sufficiently advanced encoder can pack $KB$ bits of information to that channel.
+dimensionality reduction, e.g., [perero2022x]) or by quantizing and coding the signal at a low bitrate (e.g., [wang2019vector, williams2021revisiting]).
+The trade-off between the accuracy of the trusted task and bottleneck entropy then determines the extent of privacy (see section VI).
+Observe that dimensionality reduction is, in general, not alone sufficient because if the channel’s floating- (or fixed)-point representation has $K$ dimensions each with $B$ bits, then a sufficiently advanced encoder can pack $KB$ bits of information into that channel.
 Since typical representations have $B=32$ or $B=64$, which can encode up to $2^{B}$ unique speaker identities, we see that the bottleneck is not particularly tight.
@@
 This applies to single vectors of target information passed through a bottleneck with full accuracy.
-Streaming data through the bottleneck, however, over time, leaks a flow of private information that increases in size without bound [121, 122].
-This approach also addresses only the average disclosure of private information but does not give guarantees of worst-case disclosure [87, 123].
+Streaming data through the bottleneck, however, over time, leaks a flow of private information that increases in size without bound [kairouz2015composition, backstrom2025privacydisclosure].
+This approach addresses only the average disclosure of private information, without guaranteeing worst-case disclosure [nautsch20_interspeech, vali2024porcupine].
 Information bottlenecks, or more accurately, privacy funnels, are thus theoretically well-suited for privacy-preserving applications.
-They have also been used in other areas, such as privacy-aware transmission of images [124] and face recognition [125].
+They have also been used in other areas, such as privacy-aware transmission of images [sun2024privacyaware] and face recognition [razeghi2024deep].
 However, the above-mentioned applications in speech either do not specifically apply it for privacy or use information bottlenecks optimized for privacy with additional components.
@@
 V-A2 Adversarial training
-Privacy-preserving neural networks can be trained using adversarial methods, such that side-information in the bottleneck is minimized [48, 64, 92].
+Privacy-preserving neural networks can be trained using adversarial methods, such that side-information in the bottleneck is minimized [noe21_interspeech, srivastava19_interspeech, nelus2018itg].
 Figure 13 illustrates the model structure, where the trusted and threat tasks correspond respectively to utility and privacy, which in turn respectively correspond to the extraction of the target information and side-information.
@@
 Adversarial training thus combines maximizing and minimizing the accuracy of private side information.
-This forces the encoder to minimize private side information passed through the bottleneck since the threat task tries to maximize the extraction of private side information.
-Similar effects can be achieved by a variety of other formulations, for example, by Siamese training, where the analysis model is specifically optimized to minimize and maximize feature distances between speakers within and across target classes [47].
+This forces the encoder to minimize private side information passed through the bottleneck, as the threat task aims to maximize the extraction of such information.
+Similar effects can be achieved by a variety of other formulations, for example, by Siamese training, where the analysis model is specifically optimized to minimize and maximize feature distances between speakers within and across target classes [nelus19b_interspeech].
 This approach should cluster extracted information into groups corresponding to the trusted task while minimizing private information.
@@
 The principal issue with adversarial training is that theoretical frameworks for certifying performance (e.g.
-[126]) are not yet widely accepted and there are thus no formal assurances of privacy (cf.
+[NEURIPS2023_e8b0c97b]) are not yet widely accepted and there are thus no formal assurances of privacy (cf.
 section VI-A).
@@
 User EncoderBottleneckDecoderRefer to caption OutputAnonymization!!
-Figure 14: Privacy-preserving speech processing through disentanglement, where speech is decomposed into independent streams of information, each representing a distinct category of information and where the level of anonymization can be individually chosen for each category.
+Figure 14: Privacy-preserving speech processing through disentanglement, where speech is decomposed into independent streams of information, each representing a distinct category of information, and where the level of anonymization can be individually chosen for each category.
 Potential points of attack are indicated by red, dashed arrows and exclamation marks.
@@
 With disentanglement, we refer to methods that split a speech signal into channels corresponding to distinct categories of information.
-For example, selecting from table I we can choose to decompose speech into linguistic information, the speaker’s (voice) identity, and emotions [127].
+For example, selecting from table I, we can choose to decompose speech into linguistic information, the speaker’s (voice) identity, and emotions [zhou2021vaw].
 Such disentanglement can be achieved by either or both through adversarial training or by generalizing information bottlenecks to encompass multiple bottleneck channels, where each bottleneck represents a disentangled representation (see fig.
 14).
-Note that provable privacy requires that such channels are statistically independent of each other.
-That is, each channel represents a distinct and statistically independent attribute of the signal such that the extent of anonymization can be cherry-picked per channel as per use-case e.g.
-[128, 93, 127, 20].
+Note that provable privacy requires that such channels be statistically independent of each other.
+That is, each channel represents a distinct and statistically independent attribute of the signal such that the extent of anonymization can be cherry-picked per channel as per use-case e.g., [aloufi2020privacy, stoidis21_interspeech, zhou2021vaw, williams2022learning].
 The privacy-preserving performance of a disentangling system can be readily formulated, provided that channels $k$ are uncorrelated.
-If the bitrates for private information of respective channels are $B_{k}$, then using the composition theorem, the joint disclosure is $\sum_{k}B_{k}$ [121].
+If the bitrates for private information of respective channels are $B_{k}$, then using the composition theorem, the joint disclosure is $\sum_{k}B_{k}$ [kairouz2015composition].
 Per-channel disclosures are thus additive such that the design and performance analysis of the privacy-preserving systems is straightforward.
@@
 For example, a model can be trained to match specific channels with labels in the dataset (e.g.
-[93, 127]).
-Alternatively, using representation learning, we can constrain channels based on objective criteria such as time-scope, i.e.
-the length of time over which data is integrated (e.g.
-[129, 118, 130]).
-Specifically, attributes such as the speaker’s identity are constant over an utterance, while phonemes, intensity and pitch change over a time scale of tens of milliseconds.
+[stoidis21_interspeech, zhou2021vaw]).
+Alternatively, using representation learning, we can constrain channels based on objective criteria such as time-scope, i.e., the length of time over which data is integrated (e.g.
+[Niekerk2020, neumann2019improving, lotfidereshgi2022cognitive]).
+Specifically, attributes such as the speaker’s identity are constant over an utterance, while phonemes, intensity, and pitch change over a time scale of tens of milliseconds.
 By modeling the signal in terms of such time scales, we can disentangle attributes that appear on the corresponding time scales.
@@
 Figure 15: Protecting privacy by encrypting communication and processing.
-The gray area indicates the domain that is encrypted and red, double arrows marked with exclamation marks indicate the protected stream of data.
+The gray area indicates the domain that is encrypted, and the red, double arrows marked with exclamation marks indicate the protected stream of data.
 V-B Secure processing
 If the transmitted information is encrypted, then it cannot be used for malicious purposes without access to the key.
-Surprisingly, however, it is possible to process encrypted information when using secure multi-party computation (SMPC), such that the encrypted result of the computation can be returned and opened [131, 132, 133, 108, 134] (see fig.
+Surprisingly, however, it is possible to process encrypted information when using secure multi-party computation (SMPC), such that the encrypted result of the computation can be returned and opened [yao1986generate, portelo2015logsum, portelo2014privacy, Teixeira2018, cramer2015mpc] (see fig.
 15).
-This applies largely to the same attack vectors as Information isolation, including Cloud leak (section IV-D1), False activation (section IV-D2) and Cloud access (section IV-D3), but also WASN authentication (section IV-D4) and Shared device (section IV-D7).
+This applies largely to the same attack vectors as Information isolation, including Cloud leak (section IV-D1), False activation (section IV-D2), and Cloud access (section IV-D3), but also WASN authentication (section IV-D4) and Shared device (section IV-D7).
 The SMPC umbrella contains several categories of methods such as homomorphic encryption, oblivious transfer, secret sharing, and garbled circuits.
@@
 Furthermore, when Alice receives Bob’s response, she cannot deduce the operations that Bob has performed, such that the desired task is completed with mutual privacy preserved.
-Another category of approaches is distance-preserving hashing techniques, where data is mapped into a space where distances between points are approximately maintained, allowing some operations like clustering, without revealing the actual data [21, 135].
-Such secure techniques can be applied, for example, in the extraction of spectral features of speech or speech recognition in the cloud, without plain text access to the speech signal [136, 137].
-It can also be applied in privacy-preserving speaker recognition, authentication or verification [31, 39, 41, 43], speech search [138], or for secure sound classification [139].
-Finally, we can compare the similarity of audio signals without revealing them [140, 141, 18].
-A central problem with such cryptographic protocols is that the floating- or fixed-point operations that can be performed in the encrypted domain are limited, typically to only additions and multiplications [142, 143, 144, 145].
-The limitation does not apply to the boolean operations offered by fast fully homomorphic encryption over the torus scheme, garbled circuits, and boolean secret sharing [146].
-In contrast, typical speech-processing algorithms also require other floating- or fixed-point operations, such as non-linearities, conditionals, and recursions [147], making the algorithms impossible to implement.
-While many operations have standard workarounds [148], such as approximation with additions and multiplications, such approximations are often computationally prohibitively complex, see e.g.
-[136].
+Another category of approaches is distance-preserving hashing techniques, where data is mapped into a space where distances between points are approximately maintained, allowing some operations like clustering, without revealing the actual data [portelo2015privacy, jimenez2015secure].
+Such secure techniques can be applied, for example, in the extraction of spectral features of speech or speech recognition in the cloud, without plain text access to the speech signal [thaine19_interspeech, Zhang2019].
+It can also be applied in privacy-preserving speaker recognition, authentication or verification [teixeira22_interspeech, pathak11_interspeech, pathak2012ppsa, portelo2014bprivacy], speech search [portelo2015bprivacy], or for secure sound classification [shashanka2006secure].
+Finally, we can compare the similarity of audio signals without revealing them [rane2013privacy, perez2020, zarazaga2022thesis].
+A central problem with such cryptographic protocols is that the floating-point or fixed-point operations that can be performed in the encrypted domain are limited, typically to only additions and multiplications [smaragdis2007framework, pathak2012ppsv, pathak2011privacy, pathak2013privacy].
+The limitation does not apply to the boolean operations offered by fast fully homomorphic encryption over the torus scheme, garbled circuits, and boolean secret sharing [chillotti2020tfhe].
+In contrast, typical speech-processing algorithms also require other floating- or fixed-point operations, such as non-linearities, conditionals, and recursions [chialva2018conditionals], making the algorithms impossible to implement.
+While many operations have standard workarounds [teixeira2019privacy], such as approximation with additions and multiplications, such approximations are often computationally prohibitively complex, see e.g.
+[thaine19_interspeech].
 In addition, the communication overhead, in terms of the amount of data to send in comparison to the size of the secret message, is very large.
-For a broader overview of cryptographic methods in speech, see [13].
+For a broader overview of cryptographic methods in speech, see [nautsch2019preserving].
 User Local device Remote service Secondary service (a)Refer to caption Refer to caption Refer to caption Refer to caption (b)Refer to caption Refer to caption Reduced information ! Local processing
@@
 V-C Privacy-preserving architectures
-Perhaps the most obvious solution to protecting privacy, for example in the Cloud leak–scenario, is to limit information flow (to some extent) from the local edge device to the cloud server (see fig.
+Perhaps the most obvious solution to protecting privacy, for example, in the Cloud leak–scenario, is to limit information flow (to some extent) from the local edge device to the cloud server (see fig.
 16).
@@
 At the extreme, the edge device can be entirely disconnected from the cloud or network, and process information only locally, as in fig.
-16(b) [149].
+16(b) [shi2016edge].
 In doing so, we then need to assume that the local device has sufficient capacity and software to complete the requested tasks, that it is not compromised, and that the restriction of information flow is sufficient to ensure that private information is not leaked.
@@
 1, local processing is the best protection as nothing is transmitted to a cloud service.
-Distributed learning is an approach to training models without the need for centralized data collection [150].
+Distributed learning is an approach to training models without the need for centralized data collection [liu2021machine].
 Models are trained on local edge devices and only model updates are shared between the nodes and/or with a central server (see fig.
@@
 Federated learning is one of the flavors of distributed learning, where a central, cloud server collects, merges, and redistributes model updates.
-It has been used for example in speaker and emotion recognition, language modeling, as well as for unsupervised estimation of microphone clusters in sensor networks [91, 37, 29, 80, 59, 151, 60, 152].
+It has been used, for example, in speaker and emotion recognition, language modeling, as well as for unsupervised estimation of microphone clusters in sensor networks [feng22b_interspeech, granqvist20_interspeech, zewoudie2021federated, nelus2021, shoemate2022sottovoce, ro21_interspeech, nguyen2023federated, feng2021attribute].
 Overall, distributed learning is a promising approach, even if it has several challenges.
-First, constructing, training, and testing systems architectures is much more complicated than regular machine learning.
+First, constructing, training, and testing system architectures is much more complicated than regular machine learning.
 Second, model updates are model-specific and cannot easily be reused if the model structure is updated.
 The learning accumulated during training is then effectively lost every time the model is updated, jeopardizing the fair comparison of competing approaches.
-Third, even if distributed learning does improve privacy, it is not a guarantee for privacy since model updates can also contain private information [153, 150].
+Third, even if distributed learning does improve privacy, it is not a guarantee for privacy since model updates can also contain private information [Orekondy2018gradientleaks, liu2021machine].
 To be specific, except for federated learning, distributed learning protects efficiently against threats 1 to 3 in fig.
@@
 Figure 19: Protection against a reproduction leak by using sound zones, where constructive interference between loudspeakers is used to retain intelligibility for user Alice, and destructive interference distorts it for user Eve (red dashed lines and exclamation mark).
-When a user interacts with a speech interface, then the spoken answer of the speech interface can contain private information.
-This private information can be overheard by other users in the same acoustic space, and that presents a threat to privacy (see figs.
+When a user interacts with a speech interface, the spoken answer of the speech interface can contain private information.
+Other users can overhear this private information in the same acoustic space, and that presents a threat to privacy (see figs.
 19 and IV-D5), as in attack vectors False activation (section IV-D2), WASN authentication (section IV-D4), Speech interface leak (section IV-D5), Discussion leak (section IV-D6), and Shared device (section IV-D7).
@@
 Such systems have, however, not yet been widely published.
-Figure 19 presents another solution, which uses acoustic interventions to create physical spaces where the private information is, respectively, intelligible or inaudible [154, 155].
+Figure 19 presents another solution, which uses acoustic interventions to create physical spaces where the private information is, respectively, intelligible or inaudible [gardner1998_3daudio, pulkki2018parametric].
 This acoustic effect can be achieved either as a sound zone through constructive and destructive interference between loudspeakers (see e.g.
-[156, 157, 158]) or as an audio spotlight through non-linear interference effects between ultrasound sources [159].
+[donley2016, nielsen2018, WALLACE2022101285]) or as an audio spotlight through non-linear interference effects between ultrasound sources [yoneyama1983audio].
 By choosing the spatial location where speech is intelligible and assuming we know where the target user is located, we can thus limit access to private information only to the target speaker.
@@
 The treatment thus addresses the cause rather than the symptoms thereof.
-Preserving privacy thus becomes an issue of systems design and implies using the privacy-by-design approach [160, 161].
+Preserving privacy thus becomes an issue of systems design and implies using the privacy-by-design approach [Kalloniatis2009, van2011privacy].
 This approach has, however, two principal challenges.
 First, improving utility often requires an increase in computational power and other resource consumption.
-This is financially costly and has an environmental penalty [162].
+This is financially costly and has an environmental penalty [patterson2021carbon].
 In particular, balancing the computational load is easier on a cloud server, and thus, existing resources can be more efficiently used.
 Second, speech interfaces will always have occasional errors, even with improved utility.
-That is, the design of privacy-preserving speech technology must include multiple layers of protection, especially when requested operations have large consequences [163, 164].
-Say, when the wake word detector is activated, then before any actions which would potentially breach privacy, the system could require that the speaker is in the same room, or the speaker identity is verified, or an extra confirmation step “Are you sure?” or similar.
-The intrusiveness of such additional protections should then reflect the severity of the potential breach such that the protections are not perceived as overly obtrusive and decrease the utility of the service.
+That is, the design of privacy-preserving speech technology must include multiple layers of protection, especially when requested operations have large consequences [hong2004privacy, backstrom2021edpbcomments].
+Say, when the wake word detector is activated, then before any actions which would potentially breach privacy, the system could require that the speaker is in the same room, or the speaker’s identity is verified, or an extra confirmation step “Are you sure?” or similar.
+The intrusiveness of such additional protections should then reflect the severity of the potential breach, such that the protections are not perceived as overly obtrusive and decrease the utility of the service.
 Improving performance can also mean that the system requires additional functionalities.
-For example, in the case a second local device overhears Alice’s interaction with Bob (Discussion leak) or the primary device (Speech interface leak), the required privacy protection is that the secondary device either 1.
+For example, in the case of a second local device overhearing Alice’s interaction with Bob (Discussion leak) or the primary device (Speech interface leak), the required privacy protection is that the secondary device either 1.
 is aware that it is not the intended recipient of speech (speech analysis), or 2.
@@
 V-F Privacy-by-design and provable protections
-Privacy-by-design is a philosophy or approach to design, where privacy is taken into account from the start of the design process and not as an afterthought [165].
+Privacy-by-design is a philosophy or approach to design, where privacy is taken into account from the start of the design process and not as an afterthought [schaar2010privacy].
 This increases the chances that the system does not violate privacy unnecessarily, since the privacy-preserving performance is considered at every step of the design process.
-The privacy-preserving performance of a system can be designed to satisfy design requirements provably, using theoretical frameworks, and differential privacy in particular [166, 167, 13].
+The privacy-preserving performance of a system can be designed to satisfy design requirements provably, using theoretical frameworks, and differential privacy in particular [dwork2006differential, dwork2008differential, nautsch2019preserving].
 Privacy is generally enforced by reducing the accuracy of private information sufficiently, for example, through the addition of noise or by quantization (see fig.
@@
 To determine how much accuracy must be reduced, differential privacy quantifies how the amount of information available about a speaker’s affiliation to a particular group changes due to an observation.
-Specifically, suppose an algorithm $M()$ is applied on a set $D$ to give an answer $M(D)=S$.
+Specifically, suppose an algorithm $M()$ is applied on a set $D$ to answer $M(D)=S$.
 The limit $\epsilon$ for leakage of information (in nits) is then the amount of new information gained through a slightly different dataset $D^{\prime}$ as
@@
 It is arguably more informative for users to state the absolute group size within which they remain anonymous.
-If, through the processing of an observation, an attacker can reduce the group size within which the user remains anonymous to $k$, then we say that the process is $k$-anonymous [168].
+If, through the processing of an observation, an attacker can reduce the group size within which the user remains anonymous to $k$, then we say that the process is $k$-anonymous [sweeney2002k].
 Even if individual modules of a system can assure privacy, their interaction in a system can expose privacy.
 Therefore, the privacy-preservation performance of processes involving speech or other biometric signals must be evaluated on a systems level.
-Such evaluation uses a concept of biometric reference, which is a data record that stores an attribute of the user, such as the speakers’ identity.
+Such evaluation uses a concept of biometric reference, which is a data record that stores an attribute of the user, such as the speaker’s identity.
 New samples are compared to such biometric references to find matching attributes.
-The three key concepts in such evaluation are [97, 13]
+The three key concepts in such evaluation are [ISOIEC224745_2022, nautsch2019preserving]
   • Unlinkability – comparing two or more biometric references will not reveal whether they are from the same or different persons.
@@
 At the same time, renewability ensures that leaked references can be revoked and new ones created to replace them.
-Such concepts are central to the legal definitions of privacy and first proposals for their mathematical characterization have emerged [169, 170].
+Such concepts are central to the legal definitions of privacy, and first proposals for their mathematical characterization have emerged [vauquier25_interspeech, cohen2020towards].
 These theoretical design tools and concepts can be used in the design of, say, information bottlenecks (section V-A), to quantify the maximal leak of private information.
 However, care must be taken to ensure that prior assumptions, such as the equal likelihood of all labels, are met (or enforced) in practical scenarios (cf.
-[123]).
-Theoretical methodology for privacy has been applied in, among others, differentially private speech recognition [59], emotion recognition [91], speaker anonymization [53], and federated learning in acoustic sensor network [171].
+[vali2024porcupine]).
+Theoretical methodology for privacy has been applied in, among others, differentially private speech recognition [shoemate2022sottovoce], emotion recognition [feng22b_interspeech], speaker anonymization [shamsabadi2023diff], and federated learning in an acoustic sensor network [el2022differential].
 As theoretical proofs of privacy have obvious advantages, the adoption of this methodology will likely increase.
@@
 With utility, we refer here to the usefulness of the service in its intended task.
-While the objective performance in both utility and privacy are prerequisites for any good speech application, to the user, it is central that their experience with the service or application is good.
+While the objective performance in both utility and privacy is a prerequisite for any good speech application, to the user, it is central that their experience with the service or application is good.
 Correspondingly, performance can always be measured on two dimensions, privacy and utility, as well as objective and subjective (see table IV).
-Performance with respect to objective and subjective privacy and utility, are presented in the following subsections (sections VI-A, VI-B, VI-C and VI-D), but the subjective experience of privacy is a sufficiently large and independent topic that it deserves its own section (section VII).
+Performance with respect to objective and subjective privacy and utility is presented in the following subsections (sections VI-A, VI-B, VI-C and VI-D), but the subjective experience of privacy is a sufficiently large and independent topic that it deserves its own section (section VII).
 We further need to define how the objective measures are applied in systems design to ensure privacy (see section V-F).
-The evaluation protocol of the VoicePrivacy challenge is a state-of-the-art, practical example of evaluation of both privacy and utility [98].
+The evaluation protocol of the VoicePrivacy challenge is a state-of-the-art, practical example of evaluation of both privacy and utility [tomashenko2022voiceprivacy].
 It is based on direct measurement of identifiability using signal processing and machine learning methods.
@@
 The leak is thus equal to the mutual information of $X$ and $Y$.
-A central issue with these metrics is that they are based on mathematical models of probability distributions, whereas speech is a natural signal that does not submit to exact formulas [87].
+A central issue with these metrics is that they are based on mathematical models of probability distributions, whereas speech is a natural signal that does not submit to exact formulas [nautsch20_interspeech].
 To use these formulas, we must, therefore, first choose models to describe the distributions of observations and then estimate their parameters.
 Both steps lead to inaccuracies that dilute the evidentiary power of the proofs of privacy.
-Measures commonly used to analyze privacy in speaker recognition, -verification, and biometrics include [172]:
+Measures commonly used to analyze privacy in speaker recognition, verification, and biometrics include [natgunanathan2016protection]:
   1. 1.
@@
   3. 3.
-    The equal error rate (EER) is the proportion of misclassified samples when the threshold of acceptance is adjusted to give equal FJR and FAR [56, 88].
+    The equal error rate (EER) is the proportion of misclassified samples when the threshold of acceptance is adjusted to give equal FJR and FAR [maouche20_interspeech, ISObiometric].
 An increase in EER means that the attacker has made more errors, and privacy has improved.
   4. 4.
-    The log-likelihood ratio (LLR), defined as the logarithm of the ratio of likelihoods $p(Y|H_{k})$ of the observation $Y$ conditioned with the hypotheses that either the sample is $H_{1}$, or is not $H_{0}$, the target speaker [173]
+    The log-likelihood ratio (LLR), defined as the logarithm of the ratio of likelihoods $p(Y|H_{k})$ of the observation $Y$ conditioned with the hypotheses that either the sample is $H_{1}$, or is not $H_{0}$, the target speaker [nist2021SRE]
     $LLR(s)=\log\left(\frac{p\left(Y|H_{0}\right)}{p\left(Y|H_{1}\right)}\right)=I(Y|H_{0})-I(Y|H_{1}).$ (4)
@@
   5. 5.
-    The application-independent log-likelihood-ratio cost function $C^{\min}_{llr}$ generalizes the EER by considering optimal thresholds over all possible prior probabilities and all possible error cost functions [56, 174].
+    The application-independent log-likelihood-ratio cost function $C^{\min}_{llr}$ generalizes the EER by considering optimal thresholds over all possible prior probabilities and all possible error cost functions [maouche20_interspeech, brummer2006application].
   6. 6.
-    The expected privacy disclosure from the Privacy ZEBRA framework [87] quantifies the empirical cross-entropy (the leak magnitude) integrated over all possible prior likelihoods.
+    The expected privacy disclosure from the Privacy ZEBRA framework [nautsch20_interspeech] quantifies the empirical cross-entropy (the leak magnitude) integrated over all possible prior likelihoods.
   7. 7.
-    The worst-case disclosure also from the Privacy ZEBRA framework [87] is the largest disclosure observed in a dataset.
+    The worst-case disclosure also from the Privacy ZEBRA framework [nautsch20_interspeech] is the largest disclosure observed in a dataset.
 Private information can be potentially extracted (and privacy-preserving performance evaluated) from either 1.
@@
 section V-F).
-As mentioned above, empirical evaluation of privacy using EER has been applied, for example, in the VoicePrivacy challenge [98].
-Both EER and $C^{\min}_{llr}$ are widely used in any speaker recognition [35].
-The worst-case disclosure in a quantized bottleneck in speaker recognition has been studied in [123].
-More comprehensive treatments of objective privacy metrics are provided in [13, 175].
+As mentioned above, empirical evaluation of privacy using EER has been applied, for example, in the VoicePrivacy challenge [tomashenko2022voiceprivacy].
+Both EER and $C^{\min}_{llr}$ are widely used in any speaker recognition [KINNUNEN201012].
+The worst-case disclosure in a quantized bottleneck in speaker recognition has been studied in [vali2024porcupine].
+More comprehensive treatments of objective privacy metrics are provided in [nautsch2019preserving, noe2022towards].
 VI-B Subjective evaluation of privacy
-In speaker recognition, computers are now on par or better than humans [36] and are not fooled by voice mimicry (i.e., voice actors) even when humans struggle to do so [176, 177].
+In speaker recognition, computers are now on par or better than humans [hansen2015speaker] and are not fooled by voice mimicry (i.e., voice actors) even when humans struggle to do so [vestman2020methods, vestman2020voice].
 We can thus expect that computers are better at evaluating privacy in well-defined tasks.
@@
 Similarly, if the threat is speakers’ attribute inference, the subjects should evaluate the extent to which they can identify such attributes.
-As an example, in the VoicePrivacy 2020 challenge, subjective speaker verifiability was measured by asking subjects to rate, on a scale of 1 to 10110, the similarity between an anonymized target and a reference speaker, where the reference was either the same or a different speaker [98].
+As an example, in the VoicePrivacy 2020 challenge, subjective speaker verifiability was measured by asking subjects to rate, on a scale of 1 to 10110, the similarity between an anonymized target and a reference speaker, where the reference was either the same or a different speaker [tomashenko2022voiceprivacy].
 If the distributions of similarity scores have significant differences, then the anonymization has failed.
 Alternatives to the 1 to 10110 Likert-scale include, for example, AB (are these samples the same person?) and ABX (which of AB is equal to X?) tests.
-For a comprehensive exposition of subjective evaluation methods, see [178].
+For a comprehensive exposition of subjective evaluation methods, see [zacharov2018sensory].
 VI-C Objective performance in downstream tasks
@@
 In applications where a computer analyzes the speech signal, we quantify utility by measuring the accuracy of analysis.
-For example, in automatic speech recognition, where the computer extracts the text content of a speech signal, a commonly used measure is word error rate (WER), which is the proportion of correctly identified words [179].
-By measuring the word error rate with and without privacy-preserving processing, we can determine how much privacy-protections impede utility.
+For example, in automatic speech recognition, where the computer extracts the text content of a speech signal, a commonly used measure is word error rate (WER), which is the proportion of correctly identified words [rohit2024asr].
+By measuring the word error rate with and without privacy-preserving processing, we can determine how much privacy protections impede utility.
 VI-D Subjective performance in downstream tasks
-If a human user is to listen to the output signal, then that user is the ultimate measure of performance [178].
-For example, if we implement speech enhancement to reduce the perceived noise in a signal, then a user can subjectively evaluate how close the enhanced signal is to the noiseless original [180].
-Such perceptual evaluation is best implemented using subjective listening test standards such as P.800 or MUSHRA [181, 182].
+If a human user is to listen to the output signal, then that user is the ultimate measure of performance [zacharov2018sensory].
+For example, if we implement speech enhancement to reduce the perceived noise in a signal, then a user can subjectively evaluate how close the enhanced signal is to the noiseless original [vincent2018audio].
+Such perceptual evaluation is best implemented using subjective listening test standards such as P.800 or MUSHRA [P800, MUSHRA].
 Similarly, if privacy protections degrade a speech signal, we can readily quantify such degradations by listening tests.
@@
 Individual subjects, especially those without training (known as naïve subjects), are typically unable to repeat a sequence of tests such that results would match.
-Subjects are also slow since they need to listen to samples one at a time, at real-time speeds, require time to answer between each sample, a multitude of subjects need to do the same test to gain statistical confidence in results, and tests need oversight by trained personnel.
+Subjects are also slow because they must listen to samples one at a time, at real-time speeds, and require time to answer between each sample.
+Additionally, a multitude of subjects need to complete the same test to gain statistical confidence in the results, and tests require oversight by trained personnel.
 Automated tests, known as objective measures of perceptual quality, are an alternative to such subjective tests.
 They estimate the outcome of subjective listening tests by computational means.
-For example, a classic objective measure known as PESQ estimates the outcome of a P.800 test [183].
-Later improvements include POLQA and ViSQOL [184, 185].
-Such objective measures can be used to automate quality testing at a lower price and faster speed so that repeated estimates with the same input give the same result.
+For example, a classic objective measure known as PESQ estimates the outcome of a P.800 test [rix2001perceptual].
+Later improvements include POLQA and ViSQOL [beerends2013perceptual, chinen2020visqol].
+Such objective measures can be used to automate quality testing at a lower price and faster speed, so that repeated estimates with the same input give the same result.
 Observe that researchers should carefully review the specifications of such algorithms to determine whether they apply to the problem at hand to ensure reliable results.
@@
 Yet speech devices are used by and designed for people, and they use people’s voice data.
-It is therefore essential to know what people think of speech devices, how they use and want to use them, and their preferences about privacy.
+It is therefore essential to understand people’s perceptions of speech devices, including how they use and want to use them, as well as their preferences regarding privacy.
 Much of this work falls under the fields of psychology, user-interface design, and social sciences.
@@
 People tend to have a deep sense of ownership of some things, both material and immaterial.
-In particular, people have a feeling of ownership toward information about themselves [186, 187].
+In particular, people have a feeling of ownership toward information about themselves [cichy2014extending, dawkins2017psychological].
 Such feelings are, to some extent, detached from the material consequences of breaches in privacy.
@@
 To qualitatively or quantitatively measure such effects, we also need subjective user tests.
-Concurrently, people have well-established social rules regarding human-to-human privacy [25].
-Moreover, users also tend to anthropomorphize technology, that is, treat devices and services as if they were human [188, 189], such that they will likely assume that devices apply human-like social rules to privacy.
+Concurrently, people have well-established social rules regarding human-to-human privacy [petronio2002boundaries].
+Moreover, users also tend to anthropomorphize technology, that is, treat devices and services as if they were human [nass1993anthropomorphism, cornelius2021acceptance], such that they will likely assume that devices apply human-like social rules to privacy.
 Observe that such social rules are related to human-like behavior and performance.
-That does not reveal whether they are applicable to the super-human performance that computers can possess, such as the ability to integrate information over massive databases and to retain accurate records of events long past.
-How machines should then behave, especially with respect to super-human capabilities, is then not only a question of user-interface design but of moral psychology [190]: We need discussions on a societal level of how automated services should behave and how they are allowed to behave (see also section VIII).
+That does not reveal whether they apply to the super-human performance that computers can possess, such as the ability to integrate information over massive databases and to retain accurate records of events long past.
+How machines should then behave, especially with respect to super-human capabilities, is then not only a question of user-interface design but of moral psychology [thompson2021machine]: We need discussions on a societal level of how automated services should behave and how they are allowed to behave (see also section VIII).
 VII-B Perception and experience of privacy
-Irrespective of whether a system actually preserves privacy or not, some users can perceive the system as threatening, and others can be oblivious to the threats it poses [191].
+Irrespective of whether a system actually preserves privacy or not, some users can perceive the system as threatening, and others can be oblivious to the threats it poses [pepino2020distrust].
 Clearly, users do not like using services they perceive as threatening.
@@
 Perception and experience of privacy can be approached in two alternative ways.
-We can make user studies where people interact with either machines or each other [192].
+We can make user studies where people interact with either machines or each other [leschanowsky2024review].
 The distinction is important in the sense that human-to-human interaction relies on social rules that are well-established and developed over a long time.
 We can thus expect them to be stable over time, whereas human-computer interaction is continuously evolving as people learn more.
-Moreover, by using studies of human-to-human interaction, we can learn effects related to human-like performance but can probably not rely on observing effects related to super-human performance.
-Studies of human-to-human interaction are, however, always a proxy if the actual target is to design human-to-computer interfaces [192].
+Moreover, by using studies of human-to-human interaction, we can learn effects related to human-like performance, but can probably not rely on observing effects related to super-human performance.
+Studies of human-to-human interaction are, however, always a proxy if the actual target is to design human-to-computer interfaces [leschanowsky2024review].
 Studies of human-computer interaction thus characterize the desired phenomenon directly.
 The compromise is, however, that people’s understanding of privacy with devices might not reflect the true level of privacy, and conclusions made based on the users’ opinions might then not reflect their true preferences and those preferences are moving targets.
-As people gain more experience with and as they learn more about technology, their attitudes, and perceptions of it change [192].
+As people gain more experience with and as they learn more about technology, their attitudes and perceptions of it change [leschanowsky2024review].
 Despite these shortcomings, both types of experiments are essential for improving our understanding of privacy and for improving technology.
-For example, human-to-human studies have revealed that people experience privacy differently in different acoustic environments; a noisy cafeteria can be better at masking sounds and defending against eavesdropping than a reverberant hallway [193, 17, 18, 194, 195] and multiple-bed patient rooms in hospitals have privacy-concerns [196].
-Similarly, human-machine studies have revealed that when a chatbot actively communicates choices related to privacy it improves users’ experience of privacy [197], voice interfaces with unknown features cause fear in users, and reduces retention of services [198, 199, 200] and breaches in privacy are highly detrimental to the trust in services and reduce users’ willingness to use the services, but such trust can be rebuilt [201].
+For example, human-to-human studies have revealed that people experience privacy differently in different acoustic environments; a noisy cafeteria can be better at masking sounds and defending against eavesdropping than a reverberant hallway [zarazaga19_interspeech, das2021thesis, zarazaga2022thesis, backstrom2021intuitive, Leschanowsky2020], and multiple-bed patient rooms in hospitals have privacy-concerns [reinten2017speech].
+Similarly, human-machine studies have revealed that when a chatbot actively communicates choices related to privacy it improves users’ experience of privacy [bruggemeier2022perceptions], voice interfaces with unknown features cause fear in users, and reduces retention of services [yeasmin2020privacy, Lau2018, lutz2021privacy] and breaches in privacy are highly detrimental to the trust in services and reduce users’ willingness to use the services, but such trust can be rebuilt [Kurz2021].
 VII-C User interface design
-For usable privacy, it is important that information about privacy is readily provided, changes in the extent of privacy are promptly notified, and that users have control of the level of privacy [197, 202, 203, 199, 201, 204, 205, 206].
+For usable privacy, information about privacy must be readily provided, changes in the extent of privacy are promptly notified, and that users have control of the level of privacy [bruggemeier2022perceptions, so2021voice, lewis1998designing, Lau2018, Kurz2021, dhiya2023privacy, iachello2007end, sivaraman2018smart].
 As a parallel to speech interfaces, visual interfaces can use lights or icons for monitoring and tactile interfaces for controlling the level of privacy.
-While sound can be used to monitor system status with sonification [207, 208], this is not in widespread use within signaling of privacy.
-It is a compromise between filling the acoustic space with information and the tendency of the human auditory system to block out monotonous sounds [209].
-To be effective, the sound should be perceivable when the user consciously wants to check the privacy level and changes in privacy level should evoke correspondingly and appropriately large changes in the soundscape that users consciously register such changes.
-Furthermore, user interfaces should give correct information about and allow controlling the privacy level [199, 15].
+While sound can be used to monitor system status with sonification [hildebrandt2016continuous, iber2021auditory], this is not in widespread use within signaling of privacy.
+It is a compromise between filling the acoustic space with information and the tendency of the human auditory system to block out monotonous sounds [pulkki2015communication].
+To be effective, the sound should be perceivable when the user consciously wants to check the privacy level, and changes in privacy level should evoke correspondingly and appropriately large changes in the soundscape so that users can consciously register the changes.
+Furthermore, user interfaces should provide accurate information about and enable control over the privacy level [Lau2018, backstrom2020privacy].
 In any case, it is imperative that services are designed such that they reflect the true extent of privacy.
@@
 In fact, service providers may have short-term incentives to follow such approaches to design as long as they improve overall user satisfaction.
-However, such approaches to design are known as dark patterns or deceptive design patterns, and they are considered unethical [210].
-Through deception, such dark patterns lull users into believing they are safe when, in fact, they are abused.
+However, such approaches to design are known as dark patterns or deceptive design patterns, and they are considered unethical [willis2020deception].
+Through deception, such dark patterns lull users into believing they are safe when, in fact, they are being abused.
 Therefore, good design of privacy should actively communicate and enable control of the true extent of privacy.
-Such design practice is not only ethical but also rewards service providers in the long term by improving retention of those services [201].
+Such design practice is not only ethical but also rewards service providers in the long term by improving retention of those services [Kurz2021].
 VII-D Privacy and trust
-The level of privacy provided by a speech service is intimately linked to the level of trust users experience in that service [192].
+The level of privacy provided by a speech service is intimately linked to the level of trust users experience in that service [leschanowsky2024review].
 In particular, if users experience that a service is detrimental to their privacy or, more generally, to their well-being, they will obviously not trust it.
-We thus discuss trust mainly from a psychological perspective, unlike in security research and cryptography, where trust has a specific and unambiguous mathematical definition [211].
-While the meaning of trust may seem intuitively clear, it is however difficult to define exactly. A classic definition is [212]
+We thus discuss trust mainly from a psychological perspective, unlike in security research and cryptography, where trust has a specific and unambiguous mathematical definition [balsa2022cryptography].
+While the meaning of trust may seem intuitively clear, it is, however, difficult to define exactly. A classic definition is [deutch1973resolution]
 Trust: confidence that [one] finds what is desired rather than feared.
-This definition highlights the psychological nature of trust [213].
+This definition highlights the psychological nature of trust [simpson2007psychological].
 From the current engineering perspective, this translates to the statistical confidence that the outcome of a process is useful.
-Here, usefulness refers to what the user, subjectively perceives as useful.
+Here, usefulness refers to what the user subjectively perceives as useful.
 Implicitly, this furthermore requires that we predict the outcome of that process in the form of a statistical distribution of possible outcomes and that the expectation over the possible outcomes is that the process is useful.
-Another way of expressing this is to say that trust requires dimensions of competence, integrity, and benevolence [214, 215].
+Another way of expressing this is to say that trust requires dimensions of competence, integrity, and benevolence [Chen2003, Xie2009].
 Observe the semantic difference between a system that is “benevolent” and one that is “useful”.
 The casual use of the concept of trust seems to require that the actor is indeed benevolent.
-However, it is possible to imagine a system with malicious objectives concerning Alice but produces side effects that consistently benefit Bob.
-Bob thus could potentially trust the system to act beneficially even if the system is not benevolent towards him.
+However, it is possible to imagine a system with malicious objectives concerning Alice, but that produces side effects that consistently benefit Bob.
+Bob thus could potentially trust the system to act beneficially even if it is not benevolent towards him.
 It is furthermore questionable whether we can assign an “intent” to a computational system due to its close association with psychology.
@@
 It is useful and ethical to design speech services that are trustworthy.
-This means that systems should be such that it would be logical for users to trust them [216].
+This means that systems should be such that it would be logical for users to trust them [hardin1996trustworthiness].
 Speech technology should thus be designed to have useful outcomes with statistically significant confidence, especially with respect to privacy.
-A practical approach for reducing the likelihood of negative outcomes is to build multiple layers of checks and verifications, similar to multi-factor authentication [217].
+A practical approach for reducing the likelihood of adverse outcomes is to build multiple layers of checks and verifications, similar to multi-factor authentication [ometov2018multi].
 In fig.
-4, this can be implemented by applying layers of privacy protection in each successive connection and device as well as in their internal modules and components, as described in section V-E.
+4, this can be implemented by applying layers of privacy protection in each successive connection and device, as well as in their internal modules and components, as described in section V-E.
 Activation of an undesired functionality of the system or leaking private information would thus require that every subsequent module fails simultaneously, which is unlikely in most scenarios.
@@
 While we have not (yet) seen breaches related to speech technology with consequences of comparable magnitude, these parallels highlight the potential effect of breaches.
-Scandals directly related to speech technology include eavesdropping on private persons by employees and contractors of service providers [5, 6, 221].
+Scandals directly related to speech technology include eavesdropping on private persons by employees and contractors of service providers [Lynskey2019GuardianPrivacyScandal, hern2019, day2019].
 Section VII-C also makes the point that service and technology providers have clear short-term incentives that conflict with users’ preferences for privacy.
 We can easily find examples where users do not have real choices in protecting their privacy.
-For example, suppose all friends and family of a user use a particular platform for social and voice interaction even if its privacy configuration is inadequate [222].
-The individual user faces then the choice of either disconnecting from their social network or compromising privacy.
+For example, suppose all friends and family of a user use a particular platform for social and voice interaction, even if its privacy configuration is inadequate [backstrom2025beyonduser].
+The individual user then faces the choice of either disconnecting from their social network or compromising privacy.
 This applies to all users with a sufficiently large portion of their communities participating in that platform.
@@
 On the individual level, the examples of potential exploits in table II cause clear damage to individuals, including psychological harm in particular (e.g.
-[90]).
-Additional dangers include stalkers [223, 224, 225] and oppression [222].
-While such individual damage is “small” on the societal level, their prevalence makes their joint impact significant [8].
+[natarajan2021prioris]).
+Additional dangers include stalkers [lu2019abusers, hulsen2022stalkers, avast2021stalkerware] and oppression [backstrom2025beyonduser].
+While such individual damage is “small” on the societal level, their prevalence makes their joint impact significant [levy2020intimate].
 These examples demonstrate the inherent need for society-level regulation of speech technology with respect to privacy.
-Governments have already responded to this need, with the European Union spearheading the process with the General Data Protection Regulation (GDPR) [226, 227], with the State of California following soon thereafter with the California Consumer Privacy Act (CCPA) of 2018 [228].
+Governments have already responded to this need, with the European Union spearheading the process with the General Data Protection Regulation (GDPR) [eu2016gdpr, nautsch19c_interspeech], with the State of California following soon thereafter with the California Consumer Privacy Act (CCPA) of 2018 [california2018ccpa].
 While these laws cover only a small percentage of the global population, as cloud services typically operate globally, they need the capability to follow local laws.
 In many cases, it can be easiest to apply the strictest laws on all users, such that the most strict laws benefit the privacy of all users.
-Service providers have, therefore, widely adopted the requirements of GDPR and CCPA, and that has likely had a large impact also on users outside the scope of these regulations.
-More recently, other countries have followed suit, for example through the “Personal Information Protection Law of the People’s Republic of China” [229] and “Lei Geral de Proteção de Dados” (Brazilian General Data Protection Law, LGPD) [230].
+Service providers have, therefore, widely adopted the requirements of GDPR and CCPA, and that has likely had a significant impact also on users outside the scope of these regulations.
+More recently, other countries have followed suit, for example, through the “Personal Information Protection Law of the People’s Republic of China” [china2021pip] and “Lei Geral de Proteção de Dados” (Brazilian General Data Protection Law, LGPD) [brazil2020lgpd].
 With respect to regulation, an important consequence of the objective measures of privacy in section VI is that our tools and measurements will give as an output only statistical characterizations of the accuracy of estimated attributes, but they can never give absolute confidence.
-This is in stark contrast with the concept of unique identifiability used in legal documents, such as the General Data Protection Regulation (GDPR) by the European Union [226], which does not explicitly leave room for statistical uncertainty.
-This is reflected for example in the Guidelines for virtual voice assistants by the European Data Protection Board, which states that: [89, page 13, §31]
+This is in stark contrast with the concept of unique identifiability used in legal documents, such as the General Data Protection Regulation (GDPR) by the European Union [eu2016gdpr], which does not explicitly leave room for statistical uncertainty.
+This is reflected, for example, in the Guidelines for virtual voice assistants by the European Data Protection Board, which states that: [edpb2021guidelines, page 13, §31]
 … voice data is inherently biometric personal data. As a result, when such data is processed for the purpose of uniquely identifying a natural person … the processing must have a valid legal basis …
@@
 Both interpretations lead to absurdity, which suggests that the truth must lie somewhere in the middle.
-In fact, the GDPR in practice requires (see [89, page 4]) that the design process of voice assistants includes a data protection impact assessment, where the risks and consequences are evaluated such that the designer can take appropriate precautions to preserve privacy.
+In fact, the GDPR in practice requires (see [edpb2021guidelines, page 4]) that the design process of voice assistants includes a data protection impact assessment, where the risks and consequences are evaluated such that the designer can take appropriate precautions to preserve privacy.
 Authors of the GDPR are thus clearly aware that it is impossible to give absolute guarantees of privacy but that the impact assessment (i.e., objective measures of privacy) must necessarily be based on statistical measures, even if such measures have not been defined.
-Fortunately, first mathematical characterizations of the legal criteria have started to appear [169, 170].
+Fortunately, the first mathematical characterizations of the legal criteria have started to appear [vauquier25_interspeech, cohen2020towards].
 However, even in the absence of legal requirements, the current author would encourage all researchers, developers, data analysts, project managers, and policymakers to perform impact assessments in data projects, data management, and data policies, anytime when there is or could be the slightest concern that the dataset imposes risks to privacy.
-A practical tool to aid impact assessment is the “Data Ethics Decision Aid (DEDA)” that formalizes and enumerates all the questions that a representative group of stakeholders should review, discuss, and consider [231].
+A practical tool to aid impact assessment is the “Data Ethics Decision Aid (DEDA)” that formalizes and enumerates all the questions that a representative group of stakeholders should review, discuss, and consider [franzke2021data].
 While governments are in the process of regulating privacy, corporations and non-governmental organizations have also realized that proper privacy is an opportunity.
-For example, the Open Voice Network seeks to develop and standardize open technical standards and ethical guidelines for voice assistance [232] and the MyData Global seeks to help people and organizations to benefit from personal data in a human-centric way [233, 234].
-Within the research community, the author of this paper has been involved in establishing a special interest group within the International Speech Communication Association (ISCA) devoted to “Security and Privacy in Speech Communication” [235].
-It is as far as we know, the world-wide largest community focused on this topic.
+For example, the Open Voice Network seeks to develop and standardize open technical standards and ethical guidelines for voice assistance [ovn2023] and the MyData Global seeks to help people and organizations to benefit from personal data in a human-centric way [mydata2023, poikola2020mydata].
+Within the research community, the author of this paper has been involved in establishing a special interest group within the International Speech Communication Association (ISCA) devoted to “Security and Privacy in Speech Communication” [spsc2023].
+It is, as far as we know, the world’s largest community focused on this topic.
 IX Discussion and conclusions
@@
 The pertinent difference is that transmitting target information is the whole purpose of communication, and there is not much we can do to protect it other than encryption.
-With side-information, that is, all the other information that is bundled into a speech, like health status and gender identity, we have a much larger arsenal of protections.
+With side information, that is, all the other information that is bundled into a speech, like health status and gender identity, we have a much larger arsenal of protections.
 The primary approach is, however, to remove as much of the side information as possible as early as possible.
-As the target information is all the communication we need, all side-information should be removed to the extent possible.
-Such removal, however, rapidly demonstrates that paralinguistic information like speech style is often very useful in conveying the intended message.
+As the target information is all the communication we need, all side information should be removed to the extent possible.
+Such removal, however, rapidly demonstrates that paralinguistic information like speech style is often instrumental in conveying the intended message.
 It is thus not always clear what constitutes the legitimate, target information.
-The first conclusion from this paper is that the range of possible threats to privacy is vast.
-Each agent – be it human or device – participating in an interaction as well as the acoustic pathways and network connection through which they are connected, is a potential attack surface.
+The first conclusion from this paper is that the range of possible privacy threats is vast.
+Each agent – be it human or device – participating in an interaction, as well as the acoustic pathways and network connection through which they are connected, is a potential attack surface.
 Any actor that can interact with the other agent or listen to the connection is a potential eavesdropper.
@@
 That means that we need a society-wide ethical discussion about what is allowed with respect to user privacy.
-Such discussions are needed to prevent a Cambridge Analytica-style scandal for speech technology [219].
+Such discussions are necessary to prevent a Cambridge Analytica-style scandal for speech technology [heikkila2023cambridge_ai].
 A third implication of this paper is that, while research in this field has started to accelerate only recently, there is already a substantial body of research available.
 The research is not, however, mature but in a phase of rapid development, and there are important sub-areas that have not yet seen much work.
-This makes it a fruitful area for research as we can expect important understanding to be discovered in the coming years.
+This makes it a fruitful area for research, as we can expect important understanding to be discovered in the coming years.
 Particular research questions where the author sees an urgent need for and expects to see new results include:
 Consent
-While management of acquiring informed consent has established traditions and best practices for most interface types [236], speech, audio, and ambient systems are notably unique.
+While management of acquiring informed consent has established traditions and best practices for most interface types [friedman2000informed], speech, audio, and ambient systems are notably unique.
 Namely, acoustic information is a time-varying stream.
@@
 Consent should thus be acquired on an actual need basis.
-In addition to being more usable, it would also make choices better connected to the actual needs since consent is acquired only once it is needed.
+In addition to being more usable, it would also make choices better connected to the actual needs, since consent is acquired only once it is needed.
 Metrics for streaming
@@
 We would thus need methodologies for characterizing the effect of the length of observation on privacy.
+The current author has recently made first attempts in this direction [backstrom2025privacydisclosure].
 Metrics for out-of-category information
-The metrics discussed in section VI are all related to specific categories of private information and in particular, we can provide protection only to identified threats.
+The metrics discussed in section VI are all related to specific categories of private information, and in particular, we can provide protection only to identified threats.
 For example, we can measure the threat to privacy related to health information, but that does not say anything about the threat related to information about ethnic background.
@@
 Attack and attacker characterization
-The biometrics community has a long tradition of characterizing attacks as, for example, opportunistic and motivated attacks, as well as in terms of the potential rewards of such attacks [237].
+The biometrics community has a long tradition of characterizing attacks as, for example, opportunistic and motivated attacks, as well as in terms of the potential rewards of such attacks [roberts2007biometric].
 Characterization of speech-specific attacks has received much less attention, though figs.
 1 and IV in this paper are attempts in that direction.
-Further work is needed to gain privacy-protections on a systems level.
+Further work is needed to gain privacy protections on a system-level.
 Multi-user interaction
@@
 If user A records a discussion with user B, then both clearly have some level of ownership and privacy requirements on that recording.
-Another case is smart technology with multiple users, like smart TVs; even if one user has consented to data collection, that does not mean that others would agree [222].
+Another case is smart technology with multiple users, like smart TVs; even if one user has consented to data collection, that does not mean that others would agree [backstrom2025beyonduser].
 We do not yet have any widely accepted standard approaches for handling privacy, ownership, and consent in such multi-user scenarios.
@@
 The issue is that we do not yet have sufficiently sophisticated methods for such disentanglement.
-The difficulty in developing disentanglement algorithms is that information categories in table I are vaguely and heuristically defined and there is significant overlap between them.
+The difficulty in developing disentanglement algorithms is that information categories in table I are vaguely and heuristically defined, and there is significant overlap between them.
 We cannot even demonstrate that this would be a complete list of information categories.
@@
 The protection of privacy in speech technology has thus been important already for a long time, and its importance is increasing.
-Fortunately, research in the area has picked up speed, and this tutorial presents the most important concepts, approaches, and methodology.
+Fortunately, research in the area has picked up speed, and this tutorial presents the most important concepts, approaches, and methodologies.
 It is, however, likely that fundamental results and new technologies will be introduced in the near future.
@@
 I will remain forever grateful.
+Feedback from the anonymous reviewers was outstanding in both quality and extent. It pains me that I can never adequately acknowledge their exceptional work.
 A part of this work was supported by “Designing Inclusive & Trustworthy Digital Public Services for Migrants in Finland (Trust-M)” (Grant Number: 353529) funded by the Strategic Research Council of Finland.
+[Uncaptioned image] Tom Bäckström D.Sc.
+(tech.) is an associate professor at Aalto University, Finland (2019-).
+He obtained his Master’s and Doctoral degrees at Helsinki University of Technology (the predecessor of Aalto) in 2001 and 2004, respectively.
+During his time at International Audio Laboratories Erlangen, Germany (2008-2016), he made contributions to several international speech and audio coding standards such as MPEG USAC and 3GPP EVS, and became a professor (W2) at Friedrich-Alexander University Erlangen-Nürnberg (FAU) (2012-2016).
+Before his current position, he was a professor practice at Aalto University (2016-2019).
+He is the President (2025–2027) and Board Member (2023–2027) of the International Speech Communication Association (ISCA).
+He was the initiator and chair of ISCA SIG “Security and Privacy in Speech Communication” (2019-2022).
+His current research interests include privacy, trust, coding, enhancement, and transmission of speech as well as machine learning.